A Complete Bibliography of Publications in Foundations of Physics Letters

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

28 May 2018 Version 1.02

Title word cross-reference

(1 + 1) [453]. (2 + 1) [734]. (4 + 1) [686]. 1 + 1 [937]. 1/2 [695]. $19.95 [17]. 2 [724]. $22.00 [678]. $24.95 [9]. $25.00 [18]. 4 [724, 572]. $42.95 [289]. $45.00 [8]. $49.50 [19]. $56.00 [180]. 6 [703]. ^{(3)} [708, 807, 312, 363, 455, 508, 560]. ^{1} [914, 938]. ^{5+} [713]. _{2} [724, 723, 720, 713]. 3 [723, 720, 716, 713]. _{5} [720, 713]. _{7} [720]. B [500, 648, 568]. B^{(3)} [357, 359, 345, 396, 365, 404, 381, 402, 383, 410, 374, 373]. B^{(3)} = B^{(0)}k [502]. B^{0} − B̄^{0} [167]. B^{3} [592]. Bα [572]. c [784]. D [752, 766, 550]. υ̇ = f(t) + c [506]. e^{+}e^{−} → 2γ [29]. E^{α} [572]. F [892]. G [754, 958]. γ [29]. B̂^{(3)} [324]. J [814]. K [390]. K^{0} − K̄^{0} [167]. K^{0}K^{o} [681]. L = R + ΩR_{μν}R^{μν}/R + τR_{αβγδ}R^{αβγδ}/R [238]. N [306, 443, 627, 667, 726, 725]. N > 4 [163]. O(3) [669, 735, 791]. ϕ [133]. π [711]. π^{∗} [711]. Q [754]. R^{3} [236]. S^{3} [742]. SO(2n) [169]. × [592, 601, 630, 626, 742]. U(1) [609]. V [911, 955]. Z [144].

-antenna [144]. -Body [443, 627, 667, 306]. -butyldithiocarbamato [725]. -Constant [754]. -Cyclic [648, 500]. -D [703]. -dimensional [453, 766]. -Factor [754]. -Gauge [955, 911]. -imino- [724]. -Matrix [814]. -meson [390]. -Momentum [572]. -pentanononate [724]. -propyldithiocarbamate [726]. -Transitions [711]. -Vectors [572]. -Wave [550].

0 [678]. 0-691-00699-7 [678].

1 [630, 860]. 1-trifluoro- [724]. 1/ [671, 793]. 10-phenanthroline [726, 725]. 11th [553, 570]. 12-14 [434]. 12th [49, 57]. 1984 [18]. 1996 [414]. 1999 [569].

2-bipyridyl [726]. 21-26 [435]. 2nd [26].

4 [713]. 4’-bipy [713].

6= [437]. 6=-9 [437].

7 [678].

= [713].

A. [374]. abelian [381, 172, 583]. Aberration [430, 100, 138, 217]. Abraham [9, 819]. absence [340]. absolute [100, 352, 366]. absorber [507]. Accelerated [650, 597]. Accelerating [951]. Acceleration [662, 42, 275]. accepting [721]. According [800]. account [183]. acetylacetonato [720]. acting [76]. Action [587, 852, 201, 37, 39, 259]. Addendum [155, 632, 648, 43]. Adding [202]. ADM [638, 311]. Adomian [656]. Advanced [588, 338]. Æther [789]. affected [405]. Affine [877]. after [573]. Again [677]. Aging [937]. Aharonov [907, 765, 948, 348, 839, 308]. Aharonov-bohm [765]. ain’t [263]. al [568, 855]. al. [955, 961, 620]. Albert [391]. Algebra [954, 523, 135, 169, 833, 214, 109, 364]. Algebraic [545, 527, 124, 189, 121]. Algebras [696, 379, 398, 380, 98, 134, 209]. algorithm [378, 727]. alia [508]. Allen [221]. Allman [620]. allow [479, 223]. almost [416]. Altering [794]. Alternative [808, 31, 44]. Ambiguity [840, 326]. Ampère [396, 315]. Amplification [12, 41]. Amplified [41]. Amplitude [606]. analog [746]. Analogue [520, 541]. Analogy [62, 446]. Analysis [913, 830, 806, 827, 819, 239, 198, 343, 353, 197]. Analytic [351, 856, 858, 372, 54]. Analytical [397, 862, 867]. Analyzed [787]. Analyzers [750]. Analyzing [752]. Angular [48]. Anharmonic [752, 798]. anholonomic [429]. Anisotropy [855]. Anistropy [847]. Announcement [389, 376, 436, 435, 434, 437, 422, 485, 494, 518, 553, 552, 526, 219, 27, 20].

Announcements [368, 444, 561]. Anomalous [827, 83, 97]. Ansatz [752, 798]. answer [296]. antenna [144]. Anti [710, 295, 296]. Anti-Gravity [710]. anti-realism [295, 296]. Antimatter [611]. Antippa [75]. any [629]. Apparatus [252]. Apparent [304, 216, 243]. Appeal [770]. Applications [789, 588, 778, 784, 729]. Approach [605, 748, 545, 533, 412, 523, 73, 394, 171, 314, 292, 55, 146, 181, 426]. Approaches [619]. approximation [428]. Arbitrary [783, 679, 867, 705]. Argument [544, 672]. Arguments [527]. arises [453]. arising [98, 272]. Arithmetic [557]. Aron [153]. Arrow [587, 643, 928, 492, 327]. Arthur [220]. Aspects [640, 675, 48, 222, 81, 442]. Associated [948]. Associative [920, 209]. Assumptions [830, 362]. Astrophysical [910]. asymmetry [117]. Asymptotic [152]. Athens [18]. atom [714]. Atomic [639, 822]. attempt [63]. attempts [254]. Attraction [550]. August [389]. Autocorrelation [858]. autonoma [569]. auxiliary [362]. average [405]. Avoid [419]. Axiom [892]. Axioma [682]. Axiomatic [892, 55].

B [708, 807, 312, 363, 455, 433, 508, 620, 903, 183, 560]. Background [754]. ball [236, 446]. Ballentine [180]. Barrier [635]. barriers [467]. Barut [512]. baryon [471]. baryons [741]. Based [619, 919, 646, 255, 153]. Basic [853]. Basis [840, 731, 765, 293, 310, 371]. Battery [882]. Bayesian [389]. Be [546, 344, 624, 138, 405]. Beam [889]. Beams [612]. beamsplitter [12]. Beauregard [295, 200]. been [277, 337]. before [112]. Behavior [786, 266, 269, 486]. behaviour [477]. behind [464, 12]. Bell [193, 323, 182, 190, 657, 753, 347, 242, 267, 167, 378, 689, 129, 104, 681, 690, 377, 527, 339, 147, 757, 511, 163, 462, 271, 253]. Bell-like [242, 267]. Bell-type [167, 378, 377, 163]. Bergmann [957]. Bergmann-Thomson [957]. Bering [315]. Berry [172, 82, 531]. Between [865, 959, 793, 884, 851, 928, 923, 770, 365, 455, 323, 104, 298, 88, 62, 339, 216, 204, 649]. Beyond [30, 19]. Biot [396]. bipy [713]. bipyridyl [726]. Biquaternions [668]. Birkhäuser [19]. bis [725]. Bitsakis [18]. Bivector [892]. Black [589, 659, 895, 915, 957, 192, 320]. Black-Hole [915]. blackbody [62]. Bodies [939, 420]. Body [443, 865, 607, 627, 705, 667, 306, 743]. bohm [765, 907, 837, 948, 651, 795, 261, 605, 348, 839, 748, 861, 919, 654, 308, 216, 579]. Bohmian [842, 872, 899, 497]. Bohr [922]. bond [728]. Book [493, 36, 35, 48, 8, 47, 9, 180, 385, 87, 65, 678, 289, 18, 332, 19, 17, 386]. boolean [379, 398]. Boost [859]. Boosts [783]. Born [587, 950]. Bose [134, 364]. Boson [652]. Bosonic [842, 938]. bosons [565]. Boston [19]. Bound [754, 9]. Boundary [331, 536, 656, 92]. Bounds [323]. bpa [713]. Bradley [100, 138, 217]. Bradyon [498]. Branching [908]. Breakdown [666]. Breaking [790, 460, 16]. Breakthrough [771]. bridge [88]. Brief [568, 949, 760]. brightness [83]. Brillet [827]. Brillouin [118]. British [437]. Broglie [837, 605, 748, 216, 10, 151, 505, 404, 548, 464, 246]. broken [321]. Brownian [266]. Brukner [909]. Buckingham [431]. Bundle [537, 660, 203, 228, 235, 213, 276, 287, 330]. Buniy [954]. Butterfield [188]. butyldithiocarbamato [725].

C [720]. C. [175, 232]. cadmium [725]. Calculation [817, 714]. Calculations [716, 546, 547, 718, 184]. calculus [447]. Call [388, 466]. Cambridge [289]. Can [788, 344, 624, 889, 653, 482, 405, 2]. cannot [51, 138]. Canonical [878, 823, 205, 206, 506]. Canonically [554]. Carmeli [961, 951]. Carrying [882]. Cartan [820]. Cartesian [888]. Case [926, 863, 806]. Casimir [61, 879, 247, 438]. Cat [534, 439]. Categorical [674, 564]. Causal [660, 783, 873, 615, 928, 645, 483, 179, 355, 273, 442]. Causality [755, 170]. Causation [673]. Caused [621, 620, 304]. Censorship [737]. center [719]. Central [798]. Centre [515]. Century [573]. CH [716]. Chang [44]. Channel [624]. channeling [211]. Channels [843]. chaos [358]. Chaotic [698, 446]. Characterization [890]. Charge [628, 560, 597, 912, 764, 272, 395]. Charged [703, 957, 481, 343]. Charges [882, 787, 461, 428]. chemical [718, 727]. chemistry [48]. Cherenkov [927]. Chiral [601]. chloro [724]. chromium [719]. Chubykalo [955]. circularly [465]. Circulating [889]. Clarendon [8]. Class [593, 307, 245]. Classical [698, 817, 911, 520, 513, 841, 955, 573, 777, 906, 934, 786, 797, 818, 953, 522, 286, 942, 186, 704, 162, 215, 229, 379, 398, 418, 566, 313, 470, 453, 571, 425, 98, 124, 123, 327, 450]. Clausius [428]. Clifford [523, 668, 169, 214]. Clifford-algebra [523]. Cliffs [180]. Clifton [188]. Clock [862, 914, 940, 917, 207]. Close [652, 882]. Close-Packed [652]. Closed [788, 593, 584, 815, 768, 100]. Closed-form [768]. Cohen [87]. Coherence [826]. Coherent [612, 157]. Coincidence [612, 949]. Cole [817, 818]. Coleman [328]. Collapse [910, 897, 782, 659, 829, 832, 240, 230, 756, 304]. College [389]. collisions [417]. Colored [868]. Comay [465, 502, 490, 560]. Comment [621, 668, 955, 77, 198, 961, 99, 818, 620, 855, 94, 197, 465, 568, 362, 135, 193, 109, 33]. Comments [817, 74, 72, 580, 663, 655, 75, 44, 39, 4, 738]. Common [544, 880]. Communication [829, 51]. commutativity [760]. Comoving [943]. Compacted [599]. Company [18]. Comparison [619, 104, 44]. compatibility [85]. compatible [399]. Complementarity [824, 796, 871]. completeness [265, 165]. Complex [276, 555, 934]. complexes [729, 726, 713]. Complexified [776]. Complexity [801]. component [474]. Composite [652]. composition [244, 406]. compound [478]. compounds [723, 715]. compressible [397]. Compton [212]. Computation [736]. Computations [818]. computer [364]. Computers [788]. Concavity [280, 142]. concept [14, 318]. Concepts [873, 106]. concerning [743, 225]. Concomitant [708]. condition [375, 194, 204]. Conditional [684, 59, 158, 175]. conditionals [232]. Conditions [536, 517, 92]. Conductivity [732]. Cones [865]. Conference [182, 190, 49, 57, 290]. Configuration [657]. configurations [127]. Confinement [621, 620]. Conflict [928, 216, 204]. conformai [321]. Conformal [640]. Conformally [599]. Conformational [717]. Congress [553, 570]. Congruences [696].

Conjectured [801]. Conjugate [554, 205, 206, 344, 711]. Conjugation [560]. Connection [919, 649, 365]. Connections [923]. Conscious [634]. Consciousness [101, 692, 829]. Consequence [558]. Consequences [943]. Conservation [631, 501, 761, 324, 23]. Considerations [420]. Consistency [676, 319]. Consistent [407, 278]. constancy [305]. Constant [754, 852, 784, 516, 61, 314]. constants [719]. constrained [73]. Constraints [651, 407]. Construction [946, 712]. constructions [249]. contain [416]. containing [719]. context [445]. Continuity [543]. continuous [763, 64]. continuum [344]. Contracted [954]. contraction [426]. Contradictory [533]. contribution [746]. Contributions [903]. controlled [832]. Controversy [819]. Conventionality [812, 896, 733, 227, 254]. conventions [177]. Cooper [550]. Coordinates [650]. coordination [719]. Copenhagen [301, 654]. Coriolis [746, 758]. correct [484]. Correlated [894, 488]. correlation [76]. Correlations [824, 825, 581, 622, 695, 793, 763, 757, 248]. Correspondence [706, 450]. Corrigenda [870, 933, 641]. Corrigendum [181]. Cosmic [930, 943, 961, 949, 831]. Cosmological [789, 943, 350, 784]. Cosmology [951, 216]. Costa [200, 295]. Could [546]. Coulomb [195]. Counterfactual [935, 499, 131, 130]. Coupled [804]. Coupling [580, 169, 314, 288]. Covariance [648, 848, 500, 106, 473, 232]. Covariant [510, 823, 799, 811, 844, 240, 572, 354, 277]. cover [8, 19]. CP [520, 541, 68, 390, 903, 349]. CP-violation [68]. CPCB [652]. CPT [552, 569, 745]. Created [517]. Creation [800, 29]. Critical [619, 909, 34, 722]. criticisms [410]. Critique [129, 193, 942]. crossings [13]. crown [729]. crucial [483]. Cryptography [932]. Crystal [724, 723, 726, 725]. crystals [728]. cultural [19]. Curing [855]. Curious [657]. Current [628, 664, 882, 787, 746, 764, 214, 143]. Curvature [873, 256, 213]. Curved [807, 610, 641, 81]. Curves [788]. Curving [893]. cutoff [247]. Cyclic [648, 500, 433]. Cyclotron [843]. cysteine [719]. cysteine-containing [719].

D [17, 703]. Dark [951]. Data [630, 951, 763]. Dating [546]. David [87, 88, 261]. Debates [917]. Debye [212]. decay [297, 390]. Decelerating [951]. Decoherence [602, 939, 681, 582]. Decomposition [825, 656, 136]. Deconstructing [743]. defined [353]. definiteness [131, 130]. definition [502, 313]. Deformation [814]. Deformed [783]. Delta [734, 530, 334]. demon [476, 772]. Density [608, 708, 515, 943, 743, 82]. Dependent [585, 691, 795, 894, 633, 596, 351]. dephasing [487]. Depth [539]. derivable [15]. Derivation [628, 708, 735, 791, 532, 357, 404, 828, 852, 844, 859, 539, 841, 806, 245, 129, 496, 193, 125, 571, 314, 50]. Derived [618, 448, 470]. Deriving [938, 647, 407]. Descartes [436]. Describe [887]. Described [825]. describes [457]. Describing [673, 76]. Description [691, 634, 425]. design [727]. Desloge [249]. detect [10, 345, 335, 126]. detecting [362]. Detection [824, 792, 636, 151, 246, 292, 274]. detector [85, 230]. determination [113]. determine [328]. Determinism [70, 953, 379, 72, 71, 18]. Deterministic [946, 399, 477]. Development [709, 849]. Device [754, 352]. Devices [810]. di- [726, 725]. diatomic [714]. Difference [884, 309, 323, 298]. Differential [820, 846, 937, 307]. difficulties [188]. Diffraction [353]. diffusion [215]. dilation [509, 426]. Dimension [591, 894]. Dimensional [907, 515, 585, 559, 59, 496, 453, 334, 54, 766]. Dimensions [752, 937, 867]. dimer [730]. diphenylguanidine [718]. Dirac [135, 907, 926, 375, 789, 168, 683, 382, 795, 152, 120, 665, 734, 805, 86, 357, 431, 418, 828, 668, 453, 81, 109, 334]. Direction [783]. directly [277]. Disc [830]. Discontinuity [797, 855]. Discover [936]. Discovery [800, 521]. Discrete [309, 317, 946, 954, 205, 206, 344, 458]. Discretization [440]. Disk [549]. dispersion [730]. Dispersions [854]. disprove [119]. Dissipation [369, 202]. dissipative [449]. Distance [227, 37, 39]. distances [728]. Distant [753, 254]. Distinguish [959]. Distribution [532, 775, 869, 606]. Distributions [671, 565, 245, 140]. disturbance [122]. divergence [317]. divergence-free [317]. divinylacetylene [717]. division [570]. DNA [810]. Do [959, 119, 164, 328]. Does [479, 223, 782, 623, 416, 37, 39]. Domain [660, 322]. Dominated [593]. don’t [262]. Dordrecht [17]. double [308]. Douglas [580]. Dowe [296]. drive [315]. Dual [31, 461, 481, 44]. Duality [922, 645, 237, 204]. Dudley [17]. dumbbell [481, 343]. During [944]. dye [41]. dye-laser [41]. Dynamic [805]. Dynamical [164, 114, 897, 445, 134, 196, 426]. Dynamics [540, 869, 596, 688, 443, 91, 102, 407]. Dyson [145].

E. [465, 502]. early [329]. Echo [905]. Economical [779]. edited [18]. edition [17]. Edwards [940]. Effect [639, 785, 2, 839, 635, 879, 730, 557, 927, 746, 765, 61, 348, 758, 380, 308, 421, 247, 438, 315]. Effects [710, 796, 369, 69, 402, 211, 711, 355, 156, 143]. efficiencies [85]. Efficiently [788]. Eftichios [18]. Ehrenfest [212]. Eigenfunctions [803]. Eigenstates [795]. eigenvalues [344]. Einstein [735, 791, 212, 585, 651, 736, 807, 523, 830, 361, 548, 68, 544, 546, 21, 28, 30, 547, 784, 473, 649, 678]. Elastic [697]. Elasticity [535]. Electric [580, 882, 884, 746, 288]. Electrical [732]. electrodynamic [324]. Electrodynamics [583, 577, 609, 710, 708, 709, 669, 693, 735, 791, 145, 683, 911, 513, 955, 775, 808, 697, 428, 128, 381, 448, 740, 522, 571, 266, 269, 143, 144, 154, 484]. Electrogravitic [849]. Electromagnetic [669, 693, 666, 691, 161, 558, 864, 881, 893, 572, 519, 516, 765, 127, 310, 357, 340, 465, 277, 425]. Electromagnetism [588, 799, 892, 915, 474, 312, 455, 448, 135, 160, 109]. Electron [910, 573, 832, 357, 345, 393, 195, 711, 721, 97, 336, 124, 148, 713, 54, 275]. electron-accepting [721]. electron-nuclear [832]. Electronic [912]. electrons [211, 311]. Electroweak [592, 601, 630]. Elementary [877, 307, 95]. Eli [19]. Embedded [807]. Embedding [816]. Emden [372]. Emergent [596]. emission [41]. Empirically [484, 328]. Empty [362, 156, 479, 371, 335]. Empty-wave [156]. Endogenous [692]. Energy [577, 639, 708, 661, 585, 591, 691, 943, 875, 787, 775, 951, 779, 856, 929, 511, 13, 831, 957, 649, 765, 501, 322, 160]. Energy-Dependent [585, 691]. energy-momentum [160]. Engine [772]. England [8, 9]. Englewood [180]. Enk [433]. Ensemble [878, 162, 245]. ensure [164]. Entails [589, 900]. entangled [744, 486]. Entanglement [679, 759, 663, 851]. Entrance [843]. Entropy [389, 625, 801, 286]. Environment [602, 787, 172]. epistemology [18]. EPR [800, 52, 76, 861, 672, 836, 527, 153, 183, 446, 248]. EPR-Experiments [800]. EPRB [750]. Equation [849, 926, 665, 600, 752, 734, 799, 811, 828, 852, 859, 539, 586, 841, 892, 934, 607, 627, 613, 883, 618, 646, 372, 367, 449, 837, 408, 382, 152, 120, 309, 86, 357, 404, 431, 418, 769, 76, 567, 125, 453, 469, 571, 768, 202, 50, 766, 238]. Equations [628, 857, 585, 558, 593, 638, 820, 873, 548, 586, 629, 668, 936, 540, 637, 578, 307, 351, 412, 145, 310, 317, 523, 433, 448, 454, 401]. Equilibrium [775]. Equivalence [683, 904, 782, 747, 597, 430, 84, 456, 311]. Equivalences [649]. equivalent [106, 472]. Era [517]. ergodic [96]. Errata [333]. Erratum [56, 78, 79, 89, 139, 622]. error [264, 391]. error-prone [264]. essence [142]. essential [326]. Established [770]. Estimate [516]. eternal [350]. eternity [223]. Ether [44, 729, 430, 99, 33, 207]. Euclidean [868, 590, 150]. Evans [849, 529, 359, 404, 418, 828, 852, 844, 859, 846, 839, 864, 881, 873, 893, 373]. evaporating [320]. Evaporation [589]. Event [623, 674, 675, 845]. events [445, 452, 763]. Everett [4, 475, 136, 75, 687, 840, 3]. evidence [45]. Evolution [910, 643, 873, 267, 833, 329, 739]. Exact [593, 600, 607, 627, 883, 705, 766]. Example [804, 566]. Examples [819, 894]. Excellence [428]. Exchange [550]. excitation [440]. excitations [174]. excited [440]. Existence [956, 860, 400, 163]. Expanding [551]. Experiment [905, 345, 750, 612, 892, 861, 780, 653, 774, 684, 631, 649, 151, 505, 338, 335, 308, 757, 183, 44, 413]. Experimental [666, 948, 374, 508, 866, 702, 827, 511, 718, 96, 146, 181]. experimentally [277, 337]. Experiments [651, 800, 690, 701, 692, 853, 400, 237, 52, 362, 21, 28, 122, 156, 126, 446]. Explanation [669, 693, 892, 390, 413]. explicit [469]. Extension [737, 837, 460, 653, 137, 760]. extensions [380]. Eye [887].

F. [88]. fact [416]. Factor [754]. Factorization [200]. factory [133]. family [459]. Faraday [785, 340, 393]. faster [51, 355]. faster-than-light [51, 355]. fate [482]. Faux [379, 398]. Faux-boolean [379, 398]. Fe [389]. feature [199]. Features [875]. Fermi [364]. fermion [525, 418]. Fermionic [872, 331]. fermions [565, 98]. Feynman [394, 946, 507]. Fibration [876]. Field [628, 708, 849, 529, 926, 558, 807, 592, 593, 560, 799, 811, 844, 820, 839, 864, 881, 873, 893, 548, 775, 892, 914, 906, 934, 519, 781, 559, 842, 872, 527, 127, 489, 95, 523, 51, 357, 359, 345, 312, 396, 404, 363, 402, 383, 410, 455, 418, 465, 508, 490, 356, 338, 373, 268, 336, 833, 504, 166, 150, 441, 238, 303, 432, 568]. Fields [874, 698, 537, 642, 610, 601, 641, 773, 893, 884, 615, 879, 838, 331, 835, 425, 214, 81, 427]. Fifteenth [389]. Fifth [591]. final [756]. Finding [952, 378]. fine [61, 220]. Finite [559, 613, 637, 128, 223, 259, 309]. Finite-Dimensional [559]. Finsler [228, 330]. Finsler-spacetime [228, 330]. Finslerian [737, 642]. First [785, 737, 666, 821, 745, 62, 134]. First-Order [737, 62]. Fits [951]. Five [585]. Five-Dimensional [585]. Fixing [409]. Fizeau [774]. Flow [841, 397]. fluctuation [833, 191, 243]. Fluctuations [636, 781, 114, 230, 171]. Fluid [58, 397]. Fluorescence [866]. Flux [868]. Force [873, 719, 923, 310, 758, 277, 764, 454, 160, 257]. Forces [948, 574, 168, 310, 298, 218]. Form [824, 894, 898, 740, 768]. Form-Preserving [898]. Formal [106, 731]. Formalism [823, 475, 205, 206, 523, 479, 279, 198, 123, 197]. Formation [679, 901, 566]. Forming [782]. Forms [359, 719]. Formula [539, 480]. Formulas [805]. Formulating [668]. Formulation [866, 572, 892, 559, 540, 86, 769, 25, 43, 268, 54, 273]. foundation [153]. Foundational [81, 353]. Foundations [588, 299, 616, 771, 182, 190, 250, 260, 26, 133, 319, 34, 66, 103, 110]. Four [86, 856, 858]. Four-space [86]. Fourier [803]. Fractional [173]. fragments [727, 712]. Frame [699, 184, 207]. Frames [751, 549, 165, 106, 472, 427]. Framework [816, 862, 777, 251]. Free [538, 823, 558, 913, 894, 947, 860, 883, 474, 317]. Free-Particle [894]. frequencies [719]. Frequency [519]. Friedman [593]. Frontiers [281, 290]. Fuchsian [614]. Fues [867]. Function [897, 869, 241, 457, 230, 304]. Functions [798, 530, 858, 317, 76, 230, 255]. Fundamental [921, 281, 290, 282, 291, 414, 435].

G [580]. Gamow [632, 603, 901]. Gas [515, 566, 714, 492, 567, 54]. Gauge [874, 628, 698, 683, 911, 955, 519, 877, 745, 214, 305]. Gauss [796, 618, 646]. Gaussian [856, 858, 894]. Gave [546, 547]. Gell [480]. Gell-Mann [480]. General [732, 583, 791, 785, 651, 638, 852, 873, 862, 806, 782, 894, 640, 49, 57, 448, 740, 470, 223, 521, 14, 473, 429, 143]. Generalization [400, 646]. Generalizations [906]. Generalized [610, 681, 700, 367, 491, 457, 135, 346, 109, 98, 123, 641]. Generalizing [668]. Generally [799, 811, 844]. Generator [669, 693, 169]. Geodesics [863]. Geometric [543, 805, 892, 165, 273, 214, 314, 98]. geometric-multispinor [214]. Geometrical [657, 683, 844, 447]. Geometro [46, 53, 81, 105]. Geometro-stochastic [46, 53, 81, 105]. Geometry [736, 574, 830, 820, 846, 411, 162, 105]. Geons [686]. Ghosal [855]. ghost [312, 373]. Ginzburg [837, 539]. Global [576]. go [407]. Goos [635]. Gordon [449, 567]. Grand [811, 820, 495]. Granularity [950]. graviation [146]. Gravitation [691, 799, 747, 257, 813, 890, 49, 57, 455, 448, 150, 154, 303, 181]. Gravitational [926, 754, 893, 863, 914, 595, 659, 838, 835, 756, 238].

Gravitationally [582]. Gravitoelectric [580, 288]. Gravitoelectric-Electric [580, 288]. graviton [107]. Gravity [710, 686, 649, 349, 321, 46, 53, 495, 456, 209, 441, 473, 115, 116]. Greece [18]. Greechie [594]. Green [317]. Greenberger [825]. Ground [817, 956, 818, 714]. Group [735, 552, 807, 665, 569, 74, 365, 7, 55]. Groups [809]. GRWP [460]. Gudder [47]. guided [140]. Gutenberg [18]. Gyro [802].

H [724, 720]. hadron [314]. Hall [180, 746, 652, 843, 827, 557]. halogen [714]. Hamilton [279, 404, 457, 934, 447, 618]. Hamilton-Schrödinger [934]. Hamiltonian [926, 342, 841, 123]. Hamiltonians [351, 898]. Hänchen [635]. Hard [788, 8, 19]. hard-cover [19]. hardcover [9, 180, 678, 289]. Hardy [622, 581, 695]. Hardy-Type [695]. Harmonic [627, 804, 705]. Harmony [770]. Healey [289, 406]. Heart [546, 547]. Heisenberg [77, 122, 29, 687, 45, 115]. Heisenberg-Picture [687]. Helmholtz [955, 911]. Helps [657]. Hertz [813]. Hestenes [805]. hexadecuplet [480]. Heywood [121]. Hidden [854, 617, 325, 475, 400, 322, 362, 459, 503, 163]. hidden-variable [459]. hierarchy [196]. High [831, 322]. higher [238]. higher-order [238]. highly [440]. Hilbert [87, 825, 171]. Hilbert-Spaces [171]. Hilgevoord [137]. historians [360, 387]. Histories [267, 676, 183, 278]. History [533, 231, 570, 521, 19]. HJW [916]. Hole [589, 895, 915, 957, 409]. Holes [659, 192, 320]. Holographic [792, 802]. Holography [876]. Home [194, 129]. Homodyne [246]. homogeneity [236]. Homogeneous [914, 469]. Honeycombs [749]. Hopf [876]. Horne [825]. Horne-Zeilinger [825]. Hsu [954]. hydrate [724]. Hydrodynamical [512]. Hydrogen [817, 795, 818]. Hyperboloid [688]. Hyperchaos [689]. hyperpolarizability [730]. Hypothesis [925, 524]. hypothetical [207].

ideal [283, 451]. Identification [475, 695]. Identity [616]. II [724, 568, 726, 414, 206, 601, 317, 550, 725, 28, 53, 294, 858, 269, 617, 144, 675]. III [719, 154, 845]. imino [724]. Implications [830]. importance [241]. Important [27, 903]. impossibility [121, 451]. inclusion [495]. incoherent [12]. Incompatible [533, 82, 419, 253]. Incomplete [530]. Incompleteness [8]. Inconsistencies [577, 609]. independence [252, 759]. independent [339]. Indeterminacy [797]. indeterminateness [284]. Indirect [903, 514]. Indistinguishable [503, 617]. individual [29]. Individuals [902]. Induced [906, 582, 525, 230, 266, 269]. induction [340, 393, 201, 144]. Inequalities [753, 347, 267, 378, 323, 40, 377, 339, 163, 462, 253]. Inequality [689, 681, 690, 242, 167, 129, 104, 193, 757, 511]. inert [714]. Inertia [535, 649, 302]. Inertial [857, 775, 781, 885, 152, 496, 14, 427]. Inference [11]. Infinite [551, 635, 400, 19]. Infinitesimal [538]. infinity [19]. inflationary [225]. Influence [711]. Information [91, 102, 736, 772, 619, 900, 909, 853, 325, 416]. Information-Based [619]. Information-Theoretic [900]. Informational [780, 165]. inherent [765]. inhomogeneity [566]. inhomogeneous [336, 425]. Initial [517]. Initializing [606]. Initio [716]. Instability [931]. Instantaneous [911, 955, 952]. Instead [676]. Integer [557]. integral [394]. integrals [64, 134]. interacting [408, 428]. Interaction [913, 851, 947, 161, 730, 116]. Interaction-Free [913, 947]. Interactions [886, 891, 501, 239]. interactive [289]. interatomic [728]. Interference [861, 780, 853, 338, 274]. Interferometry [609]. Intermediate [511]. Intermediate-Energy [511]. Internal [944]. International [57, 182, 190, 226, 233, 282, 291, 389, 414, 553, 570, 822, 219, 49, 290, 435, 26, 18, 552, 570, 569]. Interpretation [434, 301, 676, 690, 615, 934, 672, 255, 670, 899, 804, 687, 909, 273, 697, 853, 6, 96, 121, 97, 99, 308, 499, 406, 447, 33, 289]. Interpretational [675]. Interpretations [437, 654, 653, 262, 25, 43, 399]. Intersubjectivity [232]. Intertial [168]. Intramolecular [713]. Intrinsic [838]. Introducing [736]. introduction [87]. Introductory [1]. intuitive [446]. Invariance [666, 691, 354, 106, 640, 204, 305]. Invariant [809, 356]. Inverse [785, 600]. Inverse-Power [600]. investigate [133]. Investigation [909]. investigations [17]. involving [403]. Inward [9]. ion [714]. Irreducible [735, 807]. Irregular [613]. ISBN [678]. Isodualities [611]. Isometries [626]. Isotopic [495]. Isotropic [865]. Itamar [65]. Ithaca [670].

J [77, 655, 4, 175, 232]. J.-P [77]. Jacobi [279, 404, 457, 447]. Jammer [678]. Jammerian [239]. Jersey [180, 678]. John [389, 147]. joint [24, 253]. Jordan [209]. Joseph [261]. July [389, 414, 435]. June [434, 569].

Kähler [235]. Karl [38]. KeV [211]. Killing [415]. Kind [624]. Kinematic [749]. kinematics [212, 58]. Kinetic [875, 856]. Kinetics [777]. Kingdom [289]. Klein [567]. knowledge [17]. Kochen [657]. Kolmogorov [367]. Kolmogorovian [104]. Korté [328]. Kratzer [867]. Kraus [38]. ktf [724].

L [713]. label [205, 206]. Labels [902]. Laboratory [910, 100, 441]. Lagrangian [888, 342, 852, 769, 222]. Lakhtakia [374]. Lamb [128]. Lanczos [595]. Landau [174, 837, 539]. Landscapes [875]. Lane [372]. lanthanide [729]. Large [626, 524]. Large-Scale [626]. laser [41]. Lattice [900, 309, 317]. Lavenda [677]. Law [532, 844, 476, 821, 280, 396, 277, 23, 142]. Laws [923, 293, 310, 761, 324, 360, 387]. Least [618, 646]. Lecture [65]. Left [903]. Lehnert [628]. Lemma [846]. Length [921, 426]. Lengths [563]. LEP [630]. Leptons [703]. Leslie [180]. level [13]. levels [174]. Li [580]. Lie [134]. Light [556, 636, 847, 925, 822, 815, 889, 694, 576, 855, 784, 26, 237, 51, 12, 355, 417, 472, 305]. like [127, 242, 267, 501, 741, 320]. Limit [849, 512, 942, 84]. Limitation [878]. limiting [303]. Line [590]. Linear [898, 831, 351, 463, 471, 355]. Linearized [548]. Linearly [804]. link [455, 464]. Liouvillian [860]. liquid [722]. Lith [655]. Local [666, 691, 750, 663, 654, 576, 658, 85, 241, 104, 40, 119, 148, 462]. Locality [285, 925, 662, 687, 631, 252, 194, 149, 68, 153]. Localizability [179]. localization [369, 420, 64]. localized [712]. locally [472]. Loewer [391]. Logarithmic [956]. Logic [553, 570, 470, 442, 65, 87]. Logical [830]. Logics [594, 327]. London [437]. longitudinal [474, 489, 357, 340, 345, 312, 465, 490, 373]. Look [780, 112, 189]. Looking [666]. loophole [361]. Lorentz [552, 569, 99, 538, 354, 737, 666, 823, 691, 31, 32, 310, 152, 120, 665, 924, 859, 277, 884, 160, 519, 808, 33, 44, 923, 848, 7, 778]. Lorentz-covariant [354, 277]. Lorentz-Gauge [519]. Low [803, 85]. Low-Lying [803]. luxon [498]. Lying [803].

M [621, 175, 232]. Mach [181, 67, 682, 22, 784, 146, 275, 441, 649, 516]. Mach-Style [649]. Machian [313, 191, 243]. Machine [889, 319]. macrorealism [316]. Macroscopic [939, 582, 384, 420, 433]. Magnetic [907, 884, 528, 474, 489, 465, 490, 97, 336, 432]. Magnetization [345, 393]. Magneto [54, 402]. magneto-optics [402]. Magneto-oscillations [54]. major [763]. make [482]. Making [360, 387]. Malament [896]. Manifestly [510]. Mann [480]. Manuscript [546]. Many [171, 711, 231, 270]. many-electron [711]. many-successive [270]. many-worlds [231]. Maor [19]. Marchildon [75]. Maric [77]. Markov [739]. Marzke [249, 650]. Mass [463, 510, 471, 515, 943, 775, 703, 781, 649, 174, 114, 480, 365, 383, 313, 14, 191, 243, 275, 395]. Mass-Proper [510]. Masses [943]. Massive [874, 823, 959, 941, 107]. Massless [823, 584, 629, 107, 331]. Mathematical [120, 704, 310]. Mathematics [770]. matrices [82]. Matrix [814, 608, 515, 743]. Matter [593, 548, 951, 947, 408, 404, 88, 440]. Matter-Dominated [593]. Maurer [820]. Max [678]. MaxEnt [233, 226]. Maximal [42, 783, 275, 424]. Maximum [389]. Maxwell [412, 857, 558, 310, 593, 772, 586, 668, 454, 476, 936]. mean [833]. Meaning [608, 848, 271, 473]. Means [819]. Measure [596, 100, 217, 352]. Measure-Dependent [596]. measured [297]. Measurement [263, 772, 810, 284, 623, 804, 391, 947, 509, 300, 301, 262, 374, 487, 221, 169, 268, 171, 499, 459, 105, 220, 16, 253]. Measurements [913, 576, 191, 514, 164, 283, 69, 400, 419, 264, 346, 24, 113, 451]. Measures [685, 328]. Measuring [754, 336, 270]. Mechanical [661, 651, 901, 634, 514, 172, 52, 132, 477]. Mechanics [905, 434, 510, 920, 555, 605, 575, 614, 873, 816, 536, 680, 956, 700, 748, 777, 906, 919, 672, 797, 670, 780, 935, 899, 953, 533, 653, 687, 840, 682, 813, 942, 909, 831, 554, 755, 562, 26, 6, 316, 178, 96, 205, 206, 524, 162, 133, 215, 310, 85, 241, 262, 392, 319, 313, 470, 289, 452, 25, 43, 84, 179, 170, 308, 63, 286, 346, 406, 421, 92, 123, 13, 399, 497, 50, 154, 458, 273, 318, 180]. Mechanism [832, 661, 787, 901, 598, 5]. Media [927]. meeting [18]. meson [463, 480, 390]. Mesons [903]. Method [814, 752, 656, 728]. Methodology [553, 570]. Methods [389, 729]. Metric [529, 736]. Metrics [707, 585]. Mexico [389, 569]. Michael [8]. Michelson [609]. Micro [801].

Micro-Origin [801]. microcanonical [245]. Microcausality [642]. microstates [457]. microwave [345, 393]. Mikhailov [649]. mind [743, 88]. mind-body [743]. minimal [224]. Minkowski [660, 783, 819, 442]. Misrepresentation [544]. Missing [952, 464, 126]. mixed [759, 713]. mixed-valence [713]. Mixtures [15]. Mobility [185]. Modal [434, 301, 262]. Model [887, 602, 636, 371, 652, 292, 856, 858, 903, 93, 210, 460, 321, 355, 124, 148, 741, 216, 94, 270, 477]. Models [897, 750, 573, 164, 229, 398, 104, 58]. Modern [299, 182, 190, 250, 260, 66, 103, 110]. Modes [584]. modifications [723]. Modified [772, 748, 596, 32, 257]. Molecular [720, 729, 718, 715, 726, 727, 712, 730]. Molecules [810, 711]. moment [97]. Momentum [888, 577, 572, 701, 894, 898, 631, 205, 206, 48, 23, 122, 160, 45, 24]. Monopole [907, 195]. Monopoles [204]. month [521]. most [740]. Motion [865, 578, 152, 571, 401, 425, 150]. Motionless [669, 693]. motor [201]. Mott [211]. movements [509]. Moving [887, 604]. Multi [871]. Multiple [944]. multiply [297]. multispinor [214]. multivector [214]. multivector-current [214]. mutual [122].

N [742, 48, 713, 141]. Nanoparticle [866]. Nathan [423]. Nature [780, 912, 117, 22]. Navier [397]. near [722]. necessary [163]. needed [350]. neglected [149]. Nelsonian [562]. Netherlands [17, 434]. Network [887]. Neural [887]. Neutral [903, 390]. Neutrino [680, 826]. Neutrinos [552, 959, 891, 569]. Neutron [853]. Neutronization [910]. Neutrons [612]. Newton [821]. Newtonian [706, 298]. NH [713]. Nikos [18]. Ning [580]. nitrostilbenes [721]. NMR [568, 729, 431]. No [840, 407, 416]. no-go [407]. Noble [892]. Noether [363]. noise [173]. Non [283, 172, 583, 800, 920, 614, 381, 549, 806, 659, 902, 860, 207, 831, 442, 554, 443, 252, 96, 293, 310, 185, 149, 390, 153, 124, 760, 424, 320, 209]. Non-abelian [381, 172]. Non-Associative [920, 209]. non-commutativity [760]. non-electromagnetic [310]. non-ergodic [96]. non-evaporating [320]. Non-Existence [860]. Non-Fuchsian [614]. Non-ideal [283]. Non-Individuals [902]. Non-Linear [831]. non-locality [252, 149, 153]. non-maximal [424]. Non-Occurrence [659]. non-perturbative [443]. non-relativistic [293, 124]. non-separability [185]. Non-Spatiality [800]. Non-standard [207, 442, 390]. non-violation [390]. nonassociative [55]. nonautonomous [506]. noncommutativity [159]. noncontextuality [194]. Nonequivalence [99, 33]. nonet [463]. Noninertial [917]. Noninvariant [417, 472]. Nonlinear [162, 773, 134, 307, 351, 95, 196, 739]. Nonlocal [239, 201]. Nonlocality [551, 581, 622, 945, 199, 70, 93, 189, 72, 8, 188, 337, 94, 71, 131, 130]. nonperiodic [469]. Nonperturbative [773, 790]. Nonquantum [532]. Nonrelativistic [777]. Nonseparability [159, 60]. Nonstatistical [450]. Normal [613, 637]. normalization [241]. Normalized [145]. Note [924, 940, 890, 949, 941, 645, 656, 578, 500, 279, 760]. Notes [401, 65]. Novel [780]. nuclear [832, 835]. Nuclei [910, 832]. Nucleus [910]. Number [867, 234]. Numbers [668, 524].

O [724, 720]. O. [200]. obeying [458]. Object [845, 112]. Objective [589, 816, 670]. Objects [910]. observability [479, 402]. observable [344]. Observables [554, 82, 424, 253]. Observation [634, 508]. Observed [694, 138]. observer [136, 223, 405]. Observers [753, 650, 917]. Observing [534, 76]. Occam [731, 834]. Occur [623]. Occurrence [659, 452]. octet [471]. off [160]. off-shell [160]. Okubo [480]. once [411]. One [556, 515, 433, 586, 892, 521, 815, 894, 855, 2, 59, 566, 711, 166, 417, 472, 204, 873]. one- [711]. One-Dimensional [515, 59]. one-particle [566]. One-Photon [586, 166, 204]. One-Way [556, 815, 417, 472]. Oneway [352]. Only [753, 892]. opaque [467]. Operation [2]. Operational [564]. operations [69]. Operator [773, 76]. operators [205, 206, 309, 134]. Optical [541, 792, 431, 568, 550, 151, 348, 323, 119, 166, 44]. Optics [839, 347, 402, 323, 255]. optimal [727]. Orbital [715]. orbitals [712]. Order [737, 745, 62, 238]. organic [711]. Origin [852, 625, 775, 747, 801, 358, 302, 395]. orthodoxy [216]. Orthogonal [549, 806]. Orthogonality [536]. Orthomodularity [900]. oscillation [157]. Oscillations [959, 680, 826, 54]. Oscillator [907, 627, 804, 705, 766]. Oscillators [851]. Oscillatory [873]. outcomes [164, 400, 264]. Oviedo [414, 435]. Oxford [8, 9].

P [47, 77]. Packed [652]. Packet [643]. Packets [826, 856, 858, 894, 111]. Painlevé [196]. pair [29]. Pairs [550]. Pais [9]. Palpitations [546, 547]. paper [75, 18]. papers [388, 466]. Paradox [812, 904, 862, 914, 952, 801, 132, 361, 244, 176, 208, 662]. Paradoxes [928, 407]. Paradoxical [948]. paramagnetic [729]. parameter [356]. Parameters [555]. parametrization [7]. Parlett [431]. Part [266, 674, 154, 873, 269, 617, 860, 143, 144, 675, 845]. partial [307]. Particle [922, 944, 630, 803, 624, 796, 842, 872, 895, 894, 640, 860, 883, 869, 307, 127, 161, 237, 566, 5, 522, 292, 266, 269, 488, 156, 768, 417, 438, 58]. particle-like [127]. particle-wave [5]. Particles [823, 695, 629, 793, 877, 617, 310, 95, 29, 214, 503, 339, 458]. Parts [519]. Passage [589]. Past [794]. Path [824, 64, 394]. Pathology [626]. Paths [815, 946]. Pauli [795]. Peat [88]. peculiarities [149]. Pendulum [754]. Penetration [539]. Penrose [186]. pentanononate [724]. Peres [419]. Perfect [515, 581, 622]. performed [119]. periodic [464]. Permutation [705]. Perspectives [549]. Perturbation [497]. perturbative [443]. Peshkin [621]. Petrovskii [367]. Petrovskii-Piskunov [367]. Phase [621, 948, 599, 805, 844, 946, 620, 531, 869, 172, 474, 309, 317, 82, 62, 98, 327, 450]. phase-free [474]. phases [178, 165]. phenanthroline [726, 725]. Phenomena [666, 161, 767]. Phenomenological [585]. phenomenology [370]. phenomenon [464]. Phil [296]. Philosophy [437, 553, 570, 289, 17]. Phipps [342]. Phonon [550]. Phonon-Exchange [550]. Phonons [550]. photomagneton [340, 312, 324]. Photon [944, 945, 924, 605, 383, 586, 542, 702, 489, 365, 433, 490, 41, 338, 711, 274, 166, 353, 204, 432]. Photon-Solitons [702]. Photons [538, 695, 941, 371, 29, 486, 462]. Physical [437, 673, 120, 592, 601, 636, 839, 901, 731, 234, 293, 310, 121, 166, 199, 451, 407]. Physics [817, 619, 299, 886, 903, 949, 633, 770, 110, 771, 704, 182, 190, 250, 260, 281, 290, 282, 291, 414, 435, 322, 65, 18, 453, 11, 66, 103, 360, 387, 186, 48]. Picture [687]. pictures [488]. Pilot [598, 531]. Pinch [315]. Piskunov [367]. Pitowsky [65, 148]. Planck [874, 532, 852, 517, 492, 567, 767, 850, 516]. Planck-scale [767]. Plane [595]. plasma [345, 393]. Plates [882, 52]. Platform [556, 847, 855, 468]. Platforms [576, 563]. platinum [724]. Podolsky [651, 21, 28, 361, 544]. Podolsky-Rosen [361, 544]. Poincaré [365]. Point [775, 895, 722, 522, 266, 269, 428]. Polar [860]. Polarimeters [511]. polarizability [730]. Polarization [538]. polarized [465]. polyatomic [712]. polymorphous [723]. Popper [77, 653, 684, 631]. Position [894, 122, 45, 108, 24]. position-momentum [45, 24]. positive [255]. positrons [211]. Possibilities [702]. possibility [336, 425]. Possible [910, 666, 654, 808, 646, 10, 45, 413]. Postulate [558, 747]. postulates [738]. Potential [600, 734, 882, 595, 867, 528, 740, 308, 447, 202, 259, 366]. Potentials [907, 752, 798, 519]. Power [600]. pp [8, 9, 180, 678, 289, 18, 19, 17]. Pre [574, 770]. Pre-Established [770]. Pre-Relativistic [574]. Precession [751]. precognition [421]. predicting [728]. predictions [151, 52, 323, 44]. Preferential [875]. Preferred [699]. Preliminary [20]. Premature [658]. Prentice [180]. Prequantum [934]. Presence [907, 636]. Preserving [898]. Press [8, 9, 678, 289]. Pressure [943]. Princeton [678]. Principle [587, 610, 641, 792, 802, 904, 866, 900, 782, 258, 885, 597, 631, 784, 618, 516, 67, 491, 137, 84, 106, 22, 456, 146, 181, 275, 311, 441, 224, 618, 646]. Principles [785, 937, 354]. Probabilistic [651, 125]. Probabilities [651, 861, 670, 596, 711, 399]. Probability [403, 664, 671, 65, 786, 685, 647, 379, 59, 210, 21, 28, 274, 176, 158, 175, 47]. Problem [587, 911, 636, 955, 942, 634, 306, 262, 221, 391, 141, 220]. Problems [788, 656, 282, 291, 414, 435, 153]. Proca [668, 504]. Proceedings [26, 18]. Process [782, 832, 215, 499, 270, 739]. processes [242, 460, 355]. Produce [889]. Production [630]. products [377]. Program [728, 483, 445]. Prohibit [782]. projection [329]. prone [264]. Proof [418, 70, 93, 189, 400, 72, 188, 94, 71]. proofs [252]. Propagation [673, 694, 179]. Proper [510, 513, 688, 42]. Proper-Time [513, 688]. Properties [673, 445, 31, 120, 715, 721, 44, 329]. Property [406]. Proposal [126, 217, 270]. proposals [316]. Proposed [948, 237, 366, 100]. Propulsion [771]. propyldithiocarbamate [726]. Protective [300]. Protocol [932].

Protonization [910]. provide [51]. Proving [657]. Pseudoeffect [696]. pseudoharmonic [766]. PtC1 [724]. Published [546]. Publishing [18]. pulses [345, 393]. pump [315]. Pure [770, 487]. puzzle [836]. pym [713]. pyz [713].

QMSL [267]. Qualms [225]. quanitzation [73]. Quantal [786]. Quantaloids [673]. quantification [117]. quantised [427]. Quantization [698, 773, 519, 115, 116, 375, 81, 46, 53, 429]. Quantized [843]. Quantum [922, 621, 265, 434, 127, 887, 145, 718, 537, 642, 2, 301, 581, 651, 736, 215, 151, 610, 641, 792, 52, 85, 241, 798, 945, 920, 51, 555, 575, 614, 873, 487, 59, 644, 392, 539, 816, 810, 536, 586, 629, 680, 956, 652, 843, 841, 65, 210, 793, 700, 748, 932, 615, 777, 198, 901, 906, 851, 37, 565, 919, 672, 535, 797, 607, 624, 302, 747, 670, 780, 935, 781, 545, 63, 559, 842, 872, 899, 184, 804, 459, 533, 620, 616, 653, 527, 687, 840, 512, 860, 942, 909, 829, 918, 557, 685, 647]. Quantum [831, 197, 209, 606, 667, 596, 674, 675, 845, 739, 579, 634, 554, 130, 477, 514, 172, 300, 755, 80, 90, 282, 291, 414, 435, 174, 26, 6, 483, 347, 316, 178, 838, 96, 478, 205, 206, 524, 162, 161, 759, 133, 310, 419, 479, 132, 262, 149, 279, 323, 319, 87, 439, 470, 356, 284, 264, 289, 452, 25, 43, 84, 179, 326, 170, 268, 255, 308, 171, 499, 346, 105, 165, 406, 738, 421, 92, 123, 447, 202, 157, 3, 216, 231, 337, 413, 39, 131, 13, 325, 113, 451, 4, 11, 399, 458]. quantum [273, 450, 159, 318, 622, 47, 180, 65]. quantum-classical [450]. Quantum-Mechanical [514, 52, 477]. Quantum-optical [151]. quantum-potential [202]. quantumlogics [377]. quasars [83]. Quash [863]. Quash-Spherical [863]. Quasi [707, 565, 883]. Quasi-Exact [883]. Quasi-set [565]. Quasiprobability [671]. quasistatic [504]. Quaternionic [665]. Qubit [871]. question [743, 430]. questions [265].

R [742, 141]. Radiation [525, 187, 294, 597, 212, 340, 370, 30, 62, 343, 461]. Radiation-induced [525]. Raja [508]. Random [750, 763]. Rastall [94, 93]. Rays [178, 831]. Razor [731, 834]. Reaction [587, 187, 294, 343, 461]. reactivity [721]. Real [787, 590, 764, 504, 271]. Realism [654, 658, 554, 70, 85, 241, 295, 72, 240, 392, 40, 119, 71, 296, 8]. Realistic [750, 408, 308]. reality [300, 265, 80, 90, 505, 350]. really [354]. Reason [17]. Reasoning [935]. Rebits [679]. Rebuttal [960]. Recently [546]. reciprocal [22]. Reciprocity [950, 796]. Recoil [511]. reconsidered [68, 334]. Rectangular [635]. Recurrent [887]. Redhead [8, 188, 121]. Redshift [930, 961]. Reduction [692, 596, 164, 242, 369, 445, 111, 251]. Reference [549, 106, 472]. Reflection [635]. Reflections [944, 778]. Reformulation [512]. Regge [320]. Regge-like [320]. regularization [838]. Reidel [17]. Reimannian [115, 116]. Reinterpretation [338, 507]. Rejection [658]. Related [943, 723, 194, 25, 43]. Relation [865, 510, 122, 45, 24]. Relational [672]. Relations [648, 820, 655, 684, 871, 677, 403, 328, 330, 500, 137, 77, 29, 744, 320]. relative [234]. relative-number [234]. Relatively [604]. Relativistic [510, 574, 310, 749, 648, 402, 793, 782, 268, 842, 872, 899, 637, 688, 554, 307, 174, 293, 500, 445, 298, 179, 244, 401, 355, 251, 124, 55, 146, 181, 441, 232, 317]. Relativistically [830, 549]. Relativity [732, 791, 785, 437, 809, 543, 489, 638, 490, 852, 812, 862, 604, 937, 258, 928, 650, 106, 917, 640, 885, 770, 227, 328, 49, 57, 491, 32, 448, 236, 430, 470, 223, 23, 99, 521, 222, 63, 14, 153, 34, 33, 126, 207, 177, 429]. Relaxation [234, 715]. Religion [678]. remark [382, 208]. Remarks [478, 694, 69, 113, 649]. Remote [794, 92]. Reply [621, 483, 194, 410, 374, 433, 418, 465, 490, 560, 99, 71, 131, 677, 328, 431, 502, 508, 221, 135, 175]. Representation [664, 807, 803, 934, 954, 234, 457, 115, 116]. Representations [735]. Repulsion [863]. require [37, 39]. rescue [254]. Resistive [882]. Resolution [801, 836]. resolutions [564]. resolved [83]. Resolving [221]. resonance [525]. Response [198, 927, 200, 232]. Rest [703, 114, 184]. Result [896, 806, 827, 137, 44]. Results [666, 856, 858, 419]. retardation [52]. retroactive [452]. Retrocausality [158, 175, 200]. retrocausation [384]. Retrodictions [533]. Reversal [937, 633, 578, 478]. Review [493, 48, 8, 47, 9, 180, 87, 65, 678, 289, 18, 19, 17, 36, 35, 385, 332, 386]. reviews [34]. Revisited [922, 896, 821, 562]. Revisiting [445]. revival [6]. Reye [657]. Richard [48, 289]. Riemann [213, 448, 394, 923]. Right [903]. Rigid [865, 607]. Rigid-Body [865]. Rikken [418]. Ring [802]. Robertson [593, 584]. robust [278]. Role [575, 701, 662, 927, 704]. Rosen [651, 361, 544, 423, 21, 28]. Rotating [556, 830, 847, 751, 549, 576, 855, 957, 468, 563]. rotation [244, 141, 7]. Rotational [746, 509]. Rotationally [584]. Ru [713]. Ru-L-Ru [713]. Rule [647, 424]. rules [2, 171].

S [723, 568]. S. [855]. Sachs [710, 708, 709, 693, 791]. safe [360, 387]. Sagnac [839, 758, 806, 352]. Same [590]. Santa [389]. Santilli [611]. Savart [396]. SC=CH [716]. SC=CSCH [716]. scalar [213, 150]. Scale [874, 626, 943, 850, 491, 341, 767]. Scale-Related [943]. Scaling [305]. scandium [720]. Scattering [814, 59, 195, 211, 304]. scenario [225]. Scheme [674]. Schmidt [825]. Schönberg [98]. Schrödinger [351, 6, 309, 600, 752, 457, 769, 841, 439, 76, 934, 916, 607, 627, 534, 613, 768, 883, 202, 50, 766, 618, 646]. Schwarzschild [707, 411, 895, 929, 741]. Schwarzschild-like [741]. Schwinger [145]. Science [437, 101, 570, 553, 570, 17]. scientific [436]. Screened [528]. search [80, 90, 322, 17]. second [189, 280, 142, 18, 476]. Sector [864, 881]. Self [609, 764, 481, 319]. self-consistency [319]. Self-force [764]. Self-Inconsistencies [609]. Self-torque [481]. Semi [635]. Semi-Infinite [635]. Sengupta [129, 194]. separability [478, 185]. Separated [700]. September [437, 18]. Series [803, 613, 637]. set [400, 565]. setting [564]. several [403]. Shapere [17]. shell [160]. Shield [528]. Shift [621, 948, 620, 128, 243]. Sign [801]. Signal [631, 198, 197]. Signals [911, 636, 955]. silico [566]. silicon [715]. Simple [602, 594, 819, 894, 554, 571]. Simplerealizations [346]. simplest [192]. simplified [54]. simulation [111]. Simultaneity [543, 812, 896, 733, 227, 328, 254]. Simultaneous [169, 268]. sine [449]. sine-Gordon [449]. Single [945, 41, 58]. single-particle [58]. singlet [285]. Singularities [614]. Singularity [921, 873, 613]. Sitter [929]. sling [439]. slit [308, 156]. Smearing [517]. Society [437]. Soft [921]. Soliton [542]. Solitons [702]. Solution [330, 607, 372, 397, 412, 443, 718, 220]. Solutions [593, 600, 627, 867, 613, 637, 860, 705, 351, 729, 120, 722, 469, 504, 768, 766]. Solvability [883]. Solve [788, 262]. Solving [656]. Some [69, 830, 798, 663, 188, 702, 738, 113, 778, 316, 149, 430]. source [161]. Space [628, 599, 783, 880, 594, 869, 704, 309, 317, 86, 87, 21, 28, 98, 44, 327, 450]. Spaces [171]. Spacetime [287, 537, 660, 807, 610, 876, 868, 893, 863, 950, 880, 703, 929, 776, 954, 42, 203, 228, 235, 213, 276, 330, 31, 742, 81, 482, 641]. Spacetimes [809]. Spain [414, 435]. Spatial [621, 328, 515, 620, 48, 458]. Spatiality [800]. Spatially [593]. Special [543, 862, 650, 917, 227, 328, 32, 236, 430, 23, 99, 14, 92, 34, 33, 126, 108, 207, 177]. Specific [927]. Specker [657]. Spectra [639, 716]. Spectral [775]. Spectrum [798, 463, 471]. Speed [556, 847, 925, 815, 576, 855, 472, 305]. Spherical [863, 659, 756]. spherically [150]. Spin [905, 695, 795, 229, 575, 629, 880, 777, 886, 938, 285, 310, 210, 169, 336, 401, 425, 760, 148, 256, 671, 793]. Spin-1 [671, 793]. Spin-1/ [671, 793]. spin-Clifford [169]. Spin-Dependent [795]. Spin-Echo [905]. Spindle [883]. Spinning [893]. spinor [523]. Spinors [865]. Spontaneous [790, 242, 369, 566, 41, 64]. Spread [894]. Squares [618, 646]. Squires [4]. St [389]. Stability [676]. stable [320]. Stage [932]. Stairs [221]. standard [390, 207, 442, 884]. Stanley [47]. Stapp [70, 93, 72, 71]. start [63]. State [817, 679, 945, 956, 594, 818, 692, 606, 596, 164, 234, 285, 297, 408, 564, 240, 714, 756, 251, 166, 329]. State-Vector [596, 164]. statement [1]. States [538, 825, 671, 603, 632, 696, 854, 901, 934, 535, 624, 860, 685, 369, 759, 371, 744, 92, 113]. Static [914, 504, 528, 150, 432]. Stationary [934, 243]. Statistical [905, 510, 293, 906, 934, 162, 759, 230, 458, 405]. statistics [118]. Status [897]. Steady [882, 143]. Steerage [598]. steering [5]. Stewart [147]. Stochastic [754, 919, 266, 269, 323, 81, 46, 53, 105, 153, 50]. Stokes [397]. str [254]. Strange [329, 486]. strengths [303]. Stress [577]. String [491, 938, 762, 395]. strings [204]. Strong [886, 835, 314, 441]. strong-field [441]. Strongly [773, 408]. Structure [203, 95, 820, 721, 594, 722, 729, 724, 718, 723, 42, 61, 720, 717, 726, 725, 21, 28, 314, 266, 269]. Structures [674, 675, 845, 727, 106]. Stuart [175, 232]. studies [729, 718]. study [111]. Style [649]. Subquantum [777]. subset [29]. subsystems [403]. successive [270]. suggested [738]. Suggestive [647]. sum [183]. sum-over-histories [183]. Summhammer [96]. Sums [796]. superadditivity [280]. Superarrivals [922]. Superconducting [528, 369]. Superconductivity [550, 580, 288]. superheavy [832]. Superluminal [911, 955, 694, 829]. superoptic [74]. Superselection [171]. supersymmetry [742, 83]. Supported [691]. Surface [882, 860, 883]. Surfaces [782, 659]. surrealism [350]. Symmetric [584, 935, 903, 150]. symmetries [755]. Symmetrized [667]. Symmetry [790, 633, 705, 578, 16, 498, 460, 321, 22]. Symposium [250, 260, 299, 66, 103, 110, 219, 282, 291, 414, 435, 26]. Synchronicity [88]. Synchronisation [177]. Synchronization [940, 917, 207]. Synchronized [689]. Synthesis [725, 832]. System [644, 667]. Systems [661, 700, 604, 851, 856, 858, 582, 871, 234, 478, 167, 41, 714, 76, 506, 73, 488, 134, 196, 364, 407, 159]. Syzygies [761]. Szilard [772].

tachyon [370]. Tachyonic [823, 959, 879]. tachyons [75]. Talbot [796]. Tale [657]. Tambakis [18]. Tangent [537, 660, 203, 228, 235, 213, 276, 287, 330]. Targets [887]. tautomerism [718]. technology [26]. Teleported [624]. Temperature [878, 866, 224]. Tempered [603, 632, 140]. tension [315, 395]. Tensor [577, 448, 377]. Terms [930, 820, 572, 961, 898, 331, 447]. Terrestrial [138, 217]. Test [948, 691, 802, 866, 598, 653, 511, 483, 96, 32, 505, 323, 76, 413, 196, 177, 441]. Tested [750, 277]. testing [316]. Tests [958, 119, 34]. text [17]. Theorem [753, 737, 581, 911, 955, 916, 182, 190, 121, 363, 271, 657, 622]. theorems [407]. Theoretic [900]. theoretical [314]. Theoretically [936]. Theories [574, 862, 430, 99, 833, 459, 456, 33, 177]. Theory [583, 577, 609, 628, 710, 708, 709, 693, 791, 437, 666, 592, 601, 630, 805, 588, 811, 844, 820, 839, 864, 881, 873, 893, 841, 663, 906, 934, 919, 877, 886, 781, 545, 559, 842, 872, 616, 527, 512, 531, 731, 579, 307, 265, 174, 212, 136, 837, 301, 491, 32, 293, 95, 51, 356, 326, 745, 37, 565, 30, 99, 302, 521, 308, 64, 105, 507, 251, 738, 762, 447, 150, 33, 44, 3, 231, 413, 39, 427, 176, 4, 438, 462, 497, 143, 458, 473, 16]. There [840, 341, 432]. Thermal [612, 517, 230, 567, 440]. Thermodynamic [655, 677, 149]. Thermodynamics [611, 699, 915, 476]. Thermofield [663]. These [546]. thesis [254, 186]. Thick [528]. Thomas [751, 244, 7]. Thomson [957]. thought [436]. Three [907, 752, 619, 932, 624, 705, 496, 738]. Three-Body [705]. Three-Dimensional [907, 496]. Three-Particle [624]. Threshold [535]. Tight [649]. till [263]. Time [587, 556, 510, 930, 589, 635, 513, 866, 319, 468, 549, 806, 952, 797, 961, 822, 935, 928, 889, 894, 633, 918, 360, 387, 688, 578, 704, 351, 478, 341, 223, 492, 833, 92, 44, 329, 415, 509, 224, 426, 643]. Time-Dependent [894, 633, 351]. Time-Orthogonal [549, 806]. Time-Reversal [633]. Time-Symmetric [935]. Time-Temperature [866, 224]. Timelike [788]. times [483, 467]. Timing [248]. Tippe [812]. Tm [723]. Top [812]. topological [395]. Topology [908]. torque [481, 126]. Torr [580, 44]. Torsion [395, 287, 330, 272]. Torus [803, 860, 883, 768]. Total [380]. tour [140]. Traces [854]. Tracking [887]. Trajectories [795, 842, 872, 156, 157, 353]. Trajectory [664]. Transactional [913, 25, 43]. Transfer [661, 701, 713]. Transformation [751, 808, 31, 44, 7]. Transformations [538, 737, 857, 924, 604, 884, 940, 898, 885, 778, 496, 745]. transient [191, 243]. Transitions [711, 564, 62]. translation [460]. transmission [318]. Trapped [782, 659]. Trapping [875]. travel [360, 387]. traversable [482]. Treatment [603, 632, 862, 527, 506]. trifluoro [724]. triple [156]. triple-slit [156]. tris [720]. trischelates [719]. Trouton [892]. True [604]. truly [354]. Truth [845]. Tube [868]. Tunneling [944, 467, 483, 325]. Turnaround [952]. Twelfth [226, 233]. Twin [904, 952, 132, 452, 208, 662]. Twists [482]. Two [910, 753, 679, 695, 830, 612, 851, 819, 624, 692, 334, 399, 338, 711, 274, 92, 54]. Two-dimensional [334, 54]. Two-Particle [624]. two-photon [338, 711, 274]. two-time [92]. Type [695, 167, 378, 714, 506, 377, 163, 550].

U [577]. Uffink [655, 137]. Ultradistributions [603, 632]. Uncertainty [510, 610, 641, 137, 866, 655, 744, 684, 677, 631, 60, 491, 77, 122, 29, 45, 24, 224]. Understanding [48]. unhidden [325]. Unification [448, 835, 654, 218, 495]. Unified [811, 844, 820, 839, 873, 652, 545, 527]. Uniformly [597]. unifying [123]. union [570]. unipolar [144]. unique [184, 3, 4]. United [289]. Universal [815, 341]. Universality [272, 762]. Universe [593, 584, 873, 625, 644, 951, 779, 850, 399, 416, 360, 387]. Universes [707, 626, 908]. Universidad [569]. University [434, 9, 678, 289]. unknowable [373]. Unnikrishnan [836]. Unprovability [421]. unsharp [69]. unstable [297]. upon [255]. USA [389]. use [347, 407]. Used [830]. Using [603, 632, 806, 654, 656, 338]. Utrecht [434].

vacuo [340]. Vacuum [628, 639, 708, 592, 601, 943, 615, 781, 779, 931, 765, 168, 161, 474, 357, 345, 396, 381, 455]. vacuum-particle [161]. valence [728, 713]. validity [84]. Value [656, 424]. Values [845]. Variable [750, 274, 459, 292, 163]. Variables [854, 617, 475, 322, 230, 503, 429]. Variant [301]. Variational [25, 43]. variations [447]. Varying [958]. Vavilov [927]. Vector [601, 528, 596, 164, 740, 240, 329]. Vectors [572]. Velocities [809, 783, 55, 366]. Velocity [822, 74, 244, 417, 100, 259, 352]. verifiable [278]. verified [337]. Version [789, 952, 3, 4]. Versus [824, 797, 316, 489, 85, 241, 490, 392]. Via [611, 866, 832, 580, 198, 197, 288]. vibration [719]. Vibrational [716]. View [800, 223, 153]. viewpoint [202]. Vigier [529, 359, 404, 418, 77]. vii [678]. VII0 [626]. viii [8, 626]. vindicated [118]. violated [285, 167]. Violation [520, 77, 29, 903, 462, 390, 45, 413, 349, 541, 68]. Violations [689]. viscous [397]. Vision [636]. Visualization [856].

W [87]. W. [568]. Waals [714, 125]. Walker [593, 584]. Warren [568]. Wave [922, 849, 798, 897, 643, 605, 811, 828, 852, 859, 550, 826, 230, 598, 796, 856, 858, 894, 488, 637, 531, 237, 309, 317, 241, 404, 457, 362, 5, 469, 111, 292, 269, 156, 304]. Wave-function [230]. Wave-Packet [643]. Wave-Particle [796, 237]. Wavefunction [853]. Wavefunctions [667]. Wavelike [824]. Waves [548, 595, 947, 10, 151, 505, 479, 465, 371, 335, 246]. Way [556, 815, 936, 647, 10, 417, 472]. Ways [910]. Weak [849, 636, 891]. Weak-Field [849]. Weber [501, 454, 257, 143, 144, 154, 259, 366].

Weber-like [501]. Weg [701]. weight [243]. Welcher [701]. Well [551, 353]. well-defined [353]. Weyl [162]. Wheeler [249, 650, 507]. Where [112, 457]. Which [624]. width [297]. Wigner [743, 255, 141]. Within [602, 101, 29]. Without [753, 873, 953, 951, 596, 122, 247, 131, 204, 130]. WKB [84]. Workshop [226, 389, 552, 569, 233]. World [918, 763, 3, 199, 4]. worlds [231]. wormholes [482]. Wrong [960].

xiv [9, 289]. xlv [17]. xv [180]. xvi [19].

Yang [954].

Z [77, 630]. Zacatecas [569]. Zanghi [512]. Zare [48]. Zee [954]. Zeilinger [825, 909]. Zeno [327]. Zero [775, 266, 310, 269]. Zero-Point [775, 266, 269]. Zeropoint [636]. zinc [726]. Zou [817, 818].

References

Anonymous:1988:IS

[1] Anonymous. Introductory statement. Foundations of Physics Letters, 1 (1):1–2, March 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Braunstein:1988:QRE

[2] Samuel L. Braunstein and Carlton M. Caves. Quantum rules: an effect can have more than one operation. Foundations of Physics Letters, 1(1): 3–12, March 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Squires:1988:UWE

[3] Euan J. Squires. The unique world of the Everett version of quantum theory. Foundations of Physics Letters, 1(1):13–20, March 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See comments [4].

Sudbery:1988:CUW

[4] A. Sudbery. Comments on “The unique world of the Everett version of quantum theory”, by E. J. Squires. Foundations of Physics Letters, 1(1): 21–23, March 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See [3].

MacGregor:1988:PWS

[5] Malcolm H. MacGregor. A particle-wave steering mechanism. Foundations of Physics Letters, 1(1):25–45, March 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Barut:1988:RSI

[6] A. O. Barut. The revival of Schrödinger’s interpretation of quantum mechanics. Foundations of Physics Letters, 1(1):47–56, March 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ungar:1988:TRP

[7] Abraham A. Ungar. Thomas rotation and the parametrization of the Lorentz transformation group. Foundations of Physics Letters, 1(1):57– 89, March 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Fine:1988:BRI

[8] Arthur Fine. Book review: Incompleteness, nonlocality, and realism. by Michael Redhead. Clarendon Press, Oxford, England, 1987, viii + 191 pp., $45.00 (hard cover). Foundations of Physics Letters, 1(1):91–95, March 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Greenberg:1988:BRI

[9] O. W. Greenberg. Book review: Inward bound by Abraham Pais. Oxford University Press, Oxford, England, 1986, xiv + 666 pp., $24.95 (hardcover). Foundations of Physics Letters, 1(1):97–100, March 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Croca:1988:PWD

[10] J. R. Croca, A. Garuccio, and F. Selleri. On a possible way to detect de Broglie’s waves. Foundations of Physics Letters, 1(2):101–112, June 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Summhammer:1988:IQP

[11] Johann Summhammer. Inference in quantum physics. Foundations of Physics Letters, 1(2):113–137, June 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Garuccio:1988:AIL

[12] A. Garuccio, Ph. Gueret, and F. Selleri. Amplification of incoherent light behind a beamsplitter. Foundations of Physics Letters, 1(2):139–145, June 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Steeb:1988:ELC

[13] W. H. Steeb, A. J. van Tonder, C. M. Villet, and S. J. M. Brits. Energy level crossings in quantum mechanics. Foundations of Physics Letters, 1(2):147–162, June 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sachs:1988:IMC

[14] Mendel Sachs. On the inertial mass concept in special and general relativity. Foundations of Physics Letters, 1(2):163–169, June 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Alsina:1988:MD

[15] C. Alsina and B. Schweizer. Mixtures are not derivable. Foundations of Physics Letters, 1(2):171–174, June 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Zimanyi:1988:SBM

[16] G. T. Zimanyi and K. Vladar. Symmetry breaking and measurement theory. Foundations of Physics Letters, 1(2):175–185, June 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Suppe:1988:BRR

[17] Frederick Suppe. Book review: Reason and the search for knowledge: investigations in the philosophy of science. by Dudley Shapere. D. Reidel Dordrecht, The Netherlands, 1984, xlv + 438 pp., $19.95 (text edition). Foundations of Physics Letters, 1(2):187–191, June 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Marquit:1988:BRD

[18] Erwin Marquit. Book review: Determinism in physics: proceedings of the second international meeting on epistemology, Athens, September 1984. edited by Eftichios Bitsakis and Nikos Tambakis. Gutenberg Publishing Company, Athens, Greece, 1985, 305 pp., $25.00 (paper). Foundations of Physics Letters, 1(2):193–198, June 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Narlikar:1988:BRI

[19] Jayant V. Narlikar. Book review: To infinity and beyond: a cultural history of the infinite. by Eli Maor. Birkhäuser, Boston, 1987, xvi + 275 pp., $49.50 (hard-cover). Foundations of Physics Letters, 1(2):199–201, June 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1988:PA

[20] Anonymous. Preliminary announcement. Foundations of Physics Letters, 1(2):203, June 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Klay:1988:EPRa

[21] Matthias P. Kläy. Einstein–Podolsky–Rosen experiments: the structure of the probability space. I. Foundations of Physics Letters, 1(3):205–244, September 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Shannon:1988:MPR

[22] Darwin L. Shannon. Mach’s principle and the reciprocal symmetry of nature. Foundations of Physics Letters, 1(3):245–275, September 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kennedy:1988:MCL

[23] W. L. Kennedy. The momentum conservation law in special relativity. Foundations of Physics Letters, 1(3):277–286, September 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Stulpe:1988:URJ

[24] W. Stulpe, S. Gudder, and J. Hagler. An uncertainty relation for joint position-momentum measurements. Foundations of Physics Letters, 1(3): 287–292, September 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hokkyo:1988:VFT

[25] Noboru Hokkyo. Variational formulation of transactional and related interpretations of quantum mechanics. Foundations of Physics Letters, 1 (3):293–299, September 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ballentine:1988:PIS

[26] Leslie E. Ballentine. Proceedings of the 2nd international symposium on foundations of quantum mechanics in the light of new technology. Foundations of Physics Letters, 1(3):301–304, September 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1988:IA

[27] Anonymous. Important announcement. Foundations of Physics Letters, 1(4):i, December 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Klay:1988:EPRb

[28] Matthias P. Kläy. Einstein–Podolsky–Rosen experiments: the structure of the probability space. II. Foundations of Physics Letters, 1(4):305–319, December 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Maric:1988:VHU

[29] Zvonko Maric, Karl Popper, and Jean-Pierre Vigier. Violation of Heisenberg’s uncertainty relations on individual particles within subset of γ photons in e^{+}e^{−} → 2γ pair creation. Foundations of Physics Letters, 1(4): 321–332, December 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See comment [77].

Lavenda:1988:BER

[30] B. H. Lavenda. Beyond Einstein’s radiation theory. Foundations of Physics Letters, 1(4):333–341, December 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Chang:1988:DPS

[31] T. Chang and D. G. Torr. Dual properties of spacetime under an alternative Lorentz transformation. Foundations of Physics Letters, 1(4): 343–351, December 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Chang:1988:MLT

[32] T. Chang, D. G. Torr, and D. R. Gagnon. A modified Lorentz theory as a test theory of special relativity. Foundations of Physics Letters, 1 (4):353–372, December 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Spavieri:1988:NET

[33] Giafranco Spavieri. Nonequivalence of ether theories and special relativity. Comment on recent interpretation of Lorentz’ ether theory. Foundations of Physics Letters, 1(4):373–385, December 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See reply [99].

Spavieri:1988:CRT

[34] Gianfranco Spavieri. Critical reviews and tests of the foundations of special relativity. Foundations of Physics Letters, 1(4):387–393, December 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Bub:1988:BR

[35] Jeffrey Bub and David MacCallum. Book review. Foundations of Physics Letters, 1(4):395–399, December 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Bernstein:1988:BR

[36] Herbert J. Bernstein. Book review. Foundations of Physics Letters, 1(4): 401–405, December 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kraus:1989:QTD

[37] K. Kraus. Quantum theory does not require action at a distance. Foundations of Physics Letters, 2(1):1–6, February 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Reents:1989:KK

[38] Georg Reents. Karl Kraus (1938–1988). Foundations of Physics Letters, 2(1):7, February 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Stapp:1989:CQT

[39] Henry P. Stapp. Comments on “quantum theory does not require action at a distance”. Foundations of Physics Letters, 2(1):9–13, February 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Lepore:1989:NIL

[40] V. L. Lepore. New inequalities from local realism. Foundations of Physics Letters, 2(1):15–26, February 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Giovanelli:1989:ASE

[41] Raffaele Giovanelli. Amplified spontaneous emission in dye-laser systems for single photon amplification. Foundations of Physics Letters, 2(1):27– 38, February 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Brandt:1989:MPA

[42] Howard E. Brandt. Maximal proper acceleration and the structure of spacetime. Foundations of Physics Letters, 2(1):39–58, February 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hokkyo:1989:AVF

[43] Noboru Hokkyo. Addendum to “variational formulation of transactional and related interpretations of quantum mechanics”. Foundations of Physics Letters, 2(1):59–60, February 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). URL http://link.springer.

com/content/pdf/10.1007/BF00690078.pdf.

Spavieri:1989:CCT

[44] Gianfranco Spavieri. Comments on Chang and Torr’s “dual properties of space time under an alternative Lorentz transformation”. comparison of ether theory predictions with the result of a recent optical experiment. Foundations of Physics Letters, 2(1):61–70, February 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Scheer:1989:PEV

[45] Jens Scheer, Klaus Götsch, Thomas Koch, Günter Lüning, Michael Schmidt, and Heiko Ziggel. A possible evidence for a violation of Heisenberg’s position-momentum uncertainty relation. Foundations of Physics Letters, 2(1):71–79, February 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Prugovecki:1989:GSQa

[46] Eduard Prugovecki. Geometro-stochastic quantization of gravity. I. Foundations of Physics Letters, 2(1):81–104, February 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Foulis:1989:BRQ

[47] David J. Foulis. Book review: Quantum probability. by Stanley P. Gudder. Foundations of Physics Letters, 2(1):105–106, February 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Farley:1989:BRA

[48] John W. Farley. Book review: Angular momentum: Understanding spatial aspects in chemistry and physics. By Richard N. Zare. Foundations of Physics Letters, 2(1):107–109, February 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1989:ICGa

[49] Anonymous. 12th international conference on general relativity and gravitation. Foundations of Physics Letters, 2(1):111, February 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wallstrom:1989:DSE

[50] Timothy C. Wallstrom. On the derivation of the Schrödinger equation from stochastic mechanics. Foundations of Physics Letters, 2(2):113–126, March 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Eberhard:1989:QFT

[51] Phillippe H. Eberhard and Ronald R. Ross. Quantum field theory cannot provide faster-than-light communication. Foundations of Physics Letters, 2(2):127–149, March 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

DeCaro:1989:QMP

[52] Liberato De Caro. Quantum-mechanical predictions for EPR experiments with retardation plates. Foundations of Physics Letters, 2(2):151–161, March 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Prugovecki:1989:GSQb

[53] Eduard Prugovecki. Geometro-stochastic quantization of gravity. II. Foundations of Physics Letters, 2(2):163–190, March 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wang:1989:MOT

[54] Lipo Wang. Magneto-oscillations in a two-dimensional electron gas: a simplified analytic formulation. Foundations of Physics Letters, 2(2):191– 197, March 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ungar:1989:AAN

[55] Abraham A. Ungar. Axiomatic approach to the nonassociative group of relativistic velocities. Foundations of Physics Letters, 2(2):199–203, March 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1989:Ea

[56] Anonymous. Erratum. Foundations of Physics Letters, 2(2):205, March 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

URL http://link.springer.com/content/pdf/10.1007/BF00696114.
pdf.

Anonymous:1989:ICGb

[57] Anonymous. 12th international conference on general relativity and gravitation. Foundations of Physics Letters, 2(2):207, March 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Westgard:1989:FMS

[58] James B. Westgard. Fluid models of single-particle kinematics. Foundations of Physics Letters, 2(3):209–222, June 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Farina:1989:QCP

[59] J. E. G. Farina. Quantum conditional probability in one-dimensional scattering. Foundations of Physics Letters, 2(3):223–234, June 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

delaTorre:1989:UN

[60] A. C. de la Torre, P. Catuogno, and S. Ferrando. Uncertainty and nonseparability. Foundations of Physics Letters, 2(3):235–244, June 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Carazza:1989:CEF

[61] B. Carazza and G. P. Guidetti. The Casimir effect and the fine structure constant. Foundations of Physics Letters, 2(3):245–250, June 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Lavenda:1989:ABF

[62] B. H. Lavenda and J. Dunning-Davies. Analogy between first-order phase transitions and blackbody radiation. Foundations of Physics Letters, 2(3): 251–259, June 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Mugur-Schachter:1989:QMR

[63] Mioara Mugur-Schächter. Quantum mechanics and relativity: attempt at a new start. Foundations of Physics Letters, 2(3):261–286, June 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Pearle:1989:PIC

[64] Philip Pearle and Jiri Soucek. Path integrals for the continuous spontaneous localization theory. Foundations of Physics Letters, 2(3):287–296, June 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gudder:1989:BRQ

[65] Stanley P. Gudder. Book review: Quantum Probability—Quantum Logic (Lecture notes in physics 321). By Itamar Pitowsky. Foundations of Physics Letters, 2(3):297–298, June 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Viiri:1989:SFMa

[66] Jouni Viiri. Symposium on the foundations of modern physics 1990. Foundations of Physics Letters, 2(3):299, June 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Assis:1989:MP

[67] A. K. T. Assis. On Mach’s principle. Foundations of Physics Letters, 2 (4):301–318, August 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ghirardi:1989:CVE

[68] GianCarlo Ghirardi. CP-violation and Einstein’s locality reconsidered. Foundations of Physics Letters, 2(4):319–329, August 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Busch:1989:SRE

[69] Paul Busch, Marian Grabowski, and Pekka J. Lahti. Some remarks on effects, operations, and unsharp measurements. Foundations of Physics Letters, 2(4):331–345, August 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Clifton:1989:DRS

[70] Robert K. Clifton. Determinism, realism, and Stapp’s 1985 proof of nonlocality. Foundations of Physics Letters, 2(4):347–359, August 1989. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Stapp:1989:RDR

[71] Henry P. Stapp. Reply to “determinism, realism, and Stapp’s 1985 proof of nonlocality”. Foundations of Physics Letters, 2(4):361–364, August 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Fellows:1989:CDR

[72] Frederick Fellows. Comments on “determinism, realism, and Stapp’s 1985 proof of nonlocality”. Foundations of Physics Letters, 2(4):365–369, August 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Marinov:1989:AQC

[73] M. S. Marinov. An approach to quanitzation of constrained systems. Foundations of Physics Letters, 2(4):371–382, August 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Comay:1989:CSG

[74] E. Comay. Comments on superoptic group velocity. Foundations of Physics Letters, 2(4):383–388, August 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Maccarrone:1989:CPM

[75] G. Daniele Maccarrone and Erasmo Recami. Comments on a paper by Marchildon, Antippa, and Everett about tachyons. Foundations of Physics Letters, 2(4):389–394, August 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hokkyo:1989:SEO

[76] Noboru Hokkyo. Schrödinger equation as an operator acting on test functions describing observing systems and the EPR correlation. Foundations of Physics Letters, 2(4):395–399, August 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hilgevoord:1989:CVH

[77] J. Hilgevoord and J. Uffink. Comment on “Violation of Heisenberg’s uncertainty relations ….” by Z. Maric, K. Popper, and J.-P. Vigier. Foundations of Physics Letters, 2(4):401, August 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See [29].

Anonymous:1989:Eb

[78] Anonymous. Erratum. Foundations of Physics Letters, 2(4):403, August 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

URL http://link.springer.com/content/pdf/10.1007/BF00690308.
pdf.

Anonymous:1989:Ec

[79] Anonymous. Erratum. Foundations of Physics Letters, 2(4):405, August 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

URL http://link.springer.com/content/pdf/10.1007/BF00690309.
pdf.

Anonymous:1989:SQRa

[80] Anonymous. In search of quantum reality. Foundations of Physics Letters, 2(4):407, August 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Prugovecki:1989:FAG

[81] E. Prugovecki and S. Warlow. Foundational aspects of geometro-stochastic quantization of Dirac fields in curved spacetime. Foundations of Physics Letters, 2(5):409–423, October 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Datta:1989:BPD

[82] K. Datta. Berry’s phase, density matrices and incompatible observables. Foundations of Physics Letters, 2(5):425–435, October 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Shannon:1989:ABQ

[83] Darwin L. Shannon. Anomalous brightness of quasars resolved by supersymmetry. Foundations of Physics Letters, 2(5):437–469, October 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Holland:1989:VPE

[84] P. R. Holland. The validity of the principle of equivalence in the WKB limit of quantum mechanics. Foundations of Physics Letters, 2(5):471– 485, October 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

DeCaro:1989:QMV

[85] Liberato De Caro. Quantum mechanics versus local realism: compatibility for low detector efficiencies. Foundations of Physics Letters, 2(5):487–497, October 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1989:FSF

[86] A. B. Evans. Four-space formulation of Dirac’s equation. Foundations of Physics Letters, 2(5):499–501, October 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gudder:1989:BRI

[87] Stanley P. Gudder. Book review: An introduction to Hilbert space and quantum logic by David W. Cohen. Foundations of Physics Letters, 2 (5):503–504, October 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Laurikainen:1989:SBB

[88] K. V. Laurikainen. Synchronicity: the bridge between matter and mind by F. David Peat. Foundations of Physics Letters, 2(5):505–507, October 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1989:Ed

[89] Anonymous. Erratum. Foundations of Physics Letters, 2(5):509, October 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

URL http://link.springer.com/content/pdf/10.1007/BF00689819.
pdf.

Anonymous:1989:SQRb

[90] Anonymous. In search of quantum reality. Foundations of Physics Letters, 2(5):511, October 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1989:IDa

[91] Anonymous. Information dynamics. Foundations of Physics Letters, 2(5): 513, October 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Schulman:1989:RTT

[92] L. S. Schulman. Remote two-time boundary conditions and special states in quantum mechanics. Foundations of Physics Letters, 2(6):515–530, December 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Clifton:1989:RMS

[93] Robert K. Clifton. The Rastall model and Stapp’s proof of nonlocality. Foundations of Physics Letters, 2(6):531–547, December 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See comment [94].

Stapp:1989:CRM

[94] Henry P. Stapp. Comment on “The Rastall model and Stapp’s proof of nonlocality”. Foundations of Physics Letters, 2(6):549–551, December 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See [93].

Cooperstock:1989:SEP

[95] F. I. Cooperstock. Structure of elementary particles in a nonlinear field theory. Foundations of Physics Letters, 2(6):553–563, December 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Buonomano:1989:SET

[96] Vincent Buonomano. Summhammer’s experimental test of the nonergodic interpretation of quantum mechanics. Foundations of Physics Letters, 2(6):565–576, December 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

MacGregor:1989:IEA

[97] Malcolm H. MacGregor. On the interpretation of the electron anomalous magnetic moment. Foundations of Physics Letters, 2(6):577–589, December 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sherry:1989:GPS

[98] G. C. Sherry. A generalized phase space for classical fermions arising out of Schönberg’s geometric algebras. Foundations of Physics Letters, 2(6): 591–599, December 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Maciel:1989:RNE

[99] A. K. A. Maciel and J. Tiomno. Reply to “Nonequivalence of ether theories and special relativity. Comment on recent interpretation of Lorentz’ Ether theory”. Foundations of Physics Letters, 2(6):601–605, December 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See [33].

Wesley:1989:BAP

[100] J. P. Wesley. Bradley aberration proposed to measure absolute velocity of closed laboratory. Foundations of Physics Letters, 2(6):607–616, December 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1989:CWS

[101] Anonymous. Consciousness within science. Foundations of Physics Letters, 2(6):617, December 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1989:IDb

[102] Anonymous. Information dynamics. Foundations of Physics Letters, 2(6): 618, December 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Viiri:1989:SFMb

[103] Jouni Viiri. Symposium on the foundations of modern physics 1990. Foundations of Physics Letters, 2(6):619, December 1989. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Garuccio:1990:CBL

[104] A. Garuccio and D. Gutkowski. Comparison between local Kolmogorovian models and Bell’s inequality. Foundations of Physics Letters, 3(1):1–35, February 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Prugovecki:1990:GST

[105] Eduard Prugovecki. On geometro-stochastic theory of measurement and quantum geometry. Foundations of Physics Letters, 3(1):37–58, February 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Rodrigues:1990:FSC

[106] W. A. Rodrigues, Jr., M. E. F. Scanavini, and L. P. de Alcantara. Formal structures, the concepts of covariance, invariance, equivalent reference frames, and the principle relativity. Foundations of Physics Letters, 3(1): 59–79, February 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

deBeauregard:1990:MMG

[107] O. Costa de Beauregard. Massless or massive graviton? Foundations of Physics Letters, 3(1):81–85, February 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Squires:1990:WPS

[108] Euan J. Squires. Why is position special? Foundations of Physics Letters, 3(1):87–93, February 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Rodrigues:1990:CGE

[109] Waldyr A. Rodrigues, Jr. A comment on generalized electromagnetism and Dirac algebra. Foundations of Physics Letters, 3(1):95–99, February 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See reply [135].

Viiri:1990:SFM

[110] Jouni Viiri. Symposium on the foundations of modern physics 1990. Foundations of Physics Letters, 3(1):101, February 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Murayama:1990:SRW

[111] Yoshimasa Murayama. A study of the reduction of wave packets by simulation. Foundations of Physics Letters, 3(2):103–127, April 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cini:1990:WOB

[112] Marcello Cini and Maurizio Serva. Where is an object before you look at it? Foundations of Physics Letters, 3(2):129–151, April 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Stulpe:1990:SRD

[113] Werner Stulpe and Matthias Singer. Some remarks on the determination of quantum states by measurements. Foundations of Physics Letters, 3(2): 153–166, April 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Awobode:1990:DFR

[114] A. M. Awobode. Dynamical fluctuations of the rest mass. Foundations of Physics Letters, 3(2):167–178, April 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

deBeauregard:1990:QRGa

[115] O. Costa de Beauregard. Quantization of Reimannian gravity: Heisenberg representation. Foundations of Physics Letters, 3(2):179–183, April 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

deBeauregard:1990:QRGb

[116] O. Costa de Beauregard. Quantization of Reimannian gravity: interaction representation. Foundations of Physics Letters, 3(2):185–187, April 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gilat:1990:QAN

[117] G. Gilat. On the quantification of asymmetry in nature. Foundations of Physics Letters, 3(2):189–196, April 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Costantini:1990:BSV

[118] D. Costantini and U. Garibaldi. Brillouin statistics vindicated. Foundations of Physics Letters, 3(2):197–202, April 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Lepore:1990:DPO

[119] V. L. Lepore and F. Selleri. Do performed optical tests disprove local realism? Foundations of Physics Letters, 3(3):203–220, June 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Comay:1990:PSM

[120] E. Comay. Physical solutions and mathematical properties of the Lorentz– Dirac equation. Foundations of Physics Letters, 3(3):221–238, June 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Elby:1990:PIH

[121] Andrew Elby. On the physical interpretation of heywood and Redhead’s algebraic impossibility theorem. Foundations of Physics Letters, 3(3):239– 247, June 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Koch:1990:HUR

[122] T. Koch, J. Scheer, H. Ziggel, and M. Schmidt. Heisenberg’s uncertainty relation of position and momentum in experiments without mutual disturbance. Foundations of Physics Letters, 3(3):249–254, June 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sherry:1990:GHF

[123] G. C. Sherry. A generalized Hamiltonian formalism unifying classical and quantum mechanics. Foundations of Physics Letters, 3(3):255–265, June 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sherry:1990:AMC

[124] G. C. Sherry. Algebraic model of a classical non-relativistic electron. Foundations of Physics Letters, 3(3):267–283, June 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Lavenda:1990:PDV

[125] B. H. Lavenda. Probabilistic derivation of van der Waals’ equation. Foundations of Physics Letters, 3(3):285–290, June 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Spavieri:1990:PED

[126] Gianfranco Spavieri. Proposal for experiments to detect the missing torque in special relativity. Foundations of Physics Letters, 3(3):291–302, June 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Barut:1990:QPL

[127] A. O. Barut and Aaron Grant. Quantum particle-like configurations of the electromagnetic field. Foundations of Physics Letters, 3(4):303–309, August 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Boudet:1990:LSF

[128] Roger Boudet. The Lamb shift in finite electrodynamics. Foundations of Physics Letters, 3(4):311–316, August 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Elby:1990:CHS

[129] Andrew Elby. Critique of Home and Sengupta’s derivation of a Bell inequality. Foundations of Physics Letters, 3(4):317–324, August 1990. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See comment [193].

deMuynck:1990:QNC

[130] W. M. de Muynck and W. De Baere. Quantum nonlocality without counterfactual definiteness? Foundations of Physics Letters, 3(4):325–342, August 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Stapp:1990:RQN

[131] Henry P. Stapp. Reply to “quantum nonlocality without counterfactual definiteness?”. Foundations of Physics Letters, 3(4):343–346, August 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dieks:1990:QMT

[132] Dennis Dieks. A quantum mechanical twin paradox. Foundations of Physics Letters, 3(4):347–357, August 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cocolicchio:1990:FIF

[133] Decio Cocolicchio. A ϕ factory to investigate the foundations of quantum mechanics. Foundations of Physics Letters, 3(4):359–365, August 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Steeb:1990:NDS

[134] W.-H. Steeb and N. Euler. Nonlinear dynamical systems, first integrals, Bose operators, and Lie algebras. Foundations of Physics Letters, 3(4): 367–374, August 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Fryberger:1990:RCG

[135] David Fryberger. A reply to “A comment on generalized electromagnetism and Dirac algebra”. Foundations of Physics Letters, 3(4):379–386, August 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See [109].

Ben-Dov:1990:ODE

[136] Yoav Ben-Dov. An “observer decomposition” for Everett’s theory. Foundations of Physics Letters, 3(4):383–387, August 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1990:UPU

[137] A. B. Evans. Uncertainty principle and uncertainty relations: extension of a result of Uffink and Hilgevoord. Foundations of Physics Letters, 3 (4):389–393, August 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wesley:1990:TBA

[138] J. P. Wesley. Terrestrial Bradley aberration cannot be observed. Foundations of Physics Letters, 3(4):395–397, August 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1990:E

[139] Anonymous. Erratum. Foundations of Physics Letters, 3(4):399–401, August 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524

(electronic). URL http://link.springer.com/content/pdf/10.1007/
BF00769713.pdf.

Schmeelk:1990:GTN

[140] John Schmeelk. A guided tour of new tempered distributions. Foundations of Physics Letters, 3(5):403–423, October 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Simon:1990:WRS

[141] R. Simon and N. Mukunda. SO(n,1) Wigner rotation as an SL(2,R) problem. Foundations of Physics Letters, 3(5):425–434, October 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Lavanda:1990:ESL

[142] B. H. Lavanda and J. Dunning-Davies. The essence of the second law is concavity. Foundations of Physics Letters, 3(5):435–441, October 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wesley:1990:WEPa

[143] J. P. Wesley. Weber electrodynamics, Part I. General theory, steady current effects. Foundations of Physics Letters, 3(5):443–469, October 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wesley:1990:WEPb

[144] J. P. Wesley. Weber electrodynamics, Part II: unipolar induction, Zantenna. Foundations of Physics Letters, 3(5):471–490, October 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Berrondo:1990:NSD

[145] M. Berrondo and J. F. Van Huele. Normalized Schwinger–Dyson equations in quantum electrodynamics. Foundations of Physics Letters, 3(5):491– 495, October 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Woodward:1990:NEA

[146] James F. Woodward. A new experimental approach to Mach’s principle and relativistic graviation. Foundations of Physics Letters, 3(5):497–506, October 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See corrigendum [181].

Shimony:1990:JSB

[147] Abner Shimony. John Stewart Bell: 1928–1990. Foundations of Physics Letters, 3(6):507–510, December 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Simpson:1990:PME

[148] Michael Simpson. The Pitowsky model of electron spin is not local. Foundations of Physics Letters, 3(6):511–523, December 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Elitzur:1990:SNT

[149] Avshalom C. Elitzur. On some neglected thermodynamic peculiarities of quantum non-locality. Foundations of Physics Letters, 3(6):525–541, December 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sjodin:1990:SET

[150] Torgny Sjödin. A scalar Euclidean theory of gravitation: motion in a static spherically symmetric field. Foundations of Physics Letters, 3(6): 543–556, December 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Croca:1990:QOP

[151] J. R. Croca, A. Garuccio, V. L. Lepore, and R. N. Moreira. Quantumoptical predictions for an experiment on the de Broglie waves detection. Foundations of Physics Letters, 3(6):557–564, December 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Comay:1990:AIM

[152] E. Comay. Asymptotic inertial motion and Lorentz–Dirac equation. Foundations of Physics Letters, 3(6):565–575, December 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Scheer:1990:VEN

[153] Jens Scheer. A view of EPR non-locality problems based on Aron’s stochastic foundation of relativity. Foundations of Physics Letters, 3(6): 577–580, December 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wesley:1990:WEPc

[154] J. P. Wesley. Weber electrodynamics: part III. mechanics, gravitation. Foundations of Physics Letters, 3(6):581–605, December 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1990:A

[155] Anonymous. Addendum. Foundations of Physics Letters, 3(6):606, December 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524

(electronic). URL http://link.springer.com/content/pdf/10.1007/
BF00666028.pdf.

Schmidt:1991:EWE

[156] M. Schmidt and F. Selleri. Empty-wave effects on particle trajectories in triple-slit experiments. Foundations of Physics Letters, 4(1):1–17, February 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Spiller:1991:CQO

[157] T. P. Spiller, T. D. Clark, R. J. Prance, H. Prance, and D. A. Poulton. Coherent quantum oscillation trajectories. Foundations of Physics Letters, 4(1):19–35, February 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Stuart:1991:RCP

[158] C. I. J. M. Stuart. Retrocausality and conditional probability. Foundations of Physics Letters, 4(1):37–48, February 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See reply [175].

delaTorre:1991:NNQ

[159] A. C. de la Torre, P. Catuogno, and S. Ferrando. Nonseparability and noncommutativity in quantum systems. Foundations of Physics Letters, 4(1):49–59, February 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Land:1991:LFE

[160] M. C. Land and L. P. Horwitz. The Lorentz force and energy-momentum for off-shell electromagnetism. Foundations of Physics Letters, 4(1):61– 71, February 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cetto:1991:EVP

[161] A. M. Cetto and L. de la Peña. Electromagnetic vacuum-particle interaction as the source of quantum phenomena. Foundations of Physics Letters, 4(1):73–80, February 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Castro:1991:NQM

[162] Carlos Castro. Nonlinear quantum mechanics as Weyl geometry of a classical statistical ensemble. Foundations of Physics Letters, 4(1):81–99, February 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Suppes:1991:NBT

[163] Patrick Suppes and Mario Zanotti. New Bell-type inequalities for N > 4 necessary for existence of a hidden variable. Foundations of Physics Letters, 4(1):101–107, February 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Aicardi:1991:DMS

[164] F. Aicardi, A. Borsellino, G. C. Ghirardi, and R. Grassi. Dynamical models for state-vector reduction: do they ensure that measurements have outcomes? Foundations of Physics Letters, 4(2):109–128, April 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Prugovecki:1991:GPI

[165] Eduard Prugovecki. Geometric phases and the informational completeness of quantum frames. Foundations of Physics Letters, 4(2):129–148, April 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Roussel:1991:OPS

[166] Pierre Roussel. Is the one-photon state of an optical field a physical state? Foundations of Physics Letters, 4(2):149–164, April 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Datta:1991:BTI

[167] Amitava Datta and Dipankar Home. Is Bell-type inequality violated for K^{0}−K̄^{0}, B^{0}−B̄^{0}, … systems? Foundations of Physics Letters, 4(2):165– 178, April 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Caser:1991:IFD

[168] S. Caser. Intertial forces in Dirac’s vacuum. Foundations of Physics Letters, 4(2):179–189, April 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Levine:1991:SGM

[169] Robert Y. Levine. Simultaneous SO(2n) generator measurement by spin-Clifford algebra coupling. Foundations of Physics Letters, 4(2):191–195, April 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Laurikainen:1991:CQM

[170] K. V. Laurikainen. Causality and quantum mechanics. Foundations of Physics Letters, 4(2):197–201, April 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Namiki:1991:SRF

[171] Mikio Namiki and Saverio Pascazio. Superselection rules and fluctuations in the many–Hilbert-spaces approach to quantum measurement. Foundations of Physics Letters, 4(3):203–216, June 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Aldinger:1991:NAB

[172] R. R. Aldinger, A. Böhm, and M. Loewe. Non-Abelian Berry phase in a quantum mechanical environment. Foundations of Physics Letters, 4(3): 217–234, June 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wyss:1991:FN

[173] Walter Wyss. Fractional noise. Foundations of Physics Letters, 4(3):235– 246, June 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Arensburg:1991:LLM

[174] Alex Arensburg and L. P. Horwitz. Landau levels as mass excitations in relativistic quantum theory. Foundations of Physics Letters, 4(3):247– 251, June 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

deBeauregard:1991:RCP

[175] O. Costa de Beauregard. Retrocausality and conditional probability: reply to C. I. J. M. Stuart. Foundations of Physics Letters, 4(3):253–263, June 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See [158].

Stuart:1991:PPT

[176] C. I. J. M. Stuart. On a paradox of probability theory. Foundations of Physics Letters, 4(3):265–274, June 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Vetharaniam:1991:SCT

[177] I. Vetharaniam and G. E. Stedman. Synchronisation conventions in test theories of special relativity. Foundations of Physics Letters, 4(3):275– 281, June 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Boya:1991:RPQ

[178] Luis J. Boya and E. C. G. Sudarshan. Rays and phases in quantum mechanics. Foundations of Physics Letters, 4(3):283–287, June 1991. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Horwitz:1991:LCP

[179] L. P. Horwitz and M. Usher. Localizability and causal propagation in relativistic quantum mechanics. Foundations of Physics Letters, 4(3):289– 295, June 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Greenberger:1991:BRQ

[180] Daniel Greenberger. Book review: Quantum Mechanics. by Leslie E. Ballentine, Prentice Hall, Englewood Cliffs, New Jersey 1990, xv + 560 pp., $56.00 (hardcover). Foundations of Physics Letters, 4(3):297–298, June 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Woodward:1991:CNE

[181] James F. Woodward. Corrigendum: A new experimental approach to Mach’s principle and relativistic gravitation. Foundations of Physics Letters, 4(3):299, June 1991. CODEN FPLEET. ISSN 0894-9875 (print),

1572-9524 (electronic). URL http://link.springer.com/content/pdf/
10.1007/BF00665763.pdf. See [146].

Anonymous:1991:ICBa

[182] Anonymous. International conference on Bell’s theorem and the foundations of modern physics. Foundations of Physics Letters, 4(3):301, June 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sinha:1991:SHA

[183] Sukanya Sinha and Rafael D. Sorkin. A sum-over-histories account of an EPR(B) experiment. Foundations of Physics Letters, 4(4):303–335, August 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Oliver:1991:QCU

[184] M. A. Oliver. Quantum calculations in a unique rest frame. Foundations of Physics Letters, 4(4):337–350, August 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Czachor:1991:MNS

[185] Marek Czachor. Mobility and non-separability. Foundations of Physics Letters, 4(4):351–361, August 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

daCosta:1991:CPP

[186] N. C. A. da Costa and F. A. Doria. Classical physics and Penrose’s thesis. Foundations of Physics Letters, 4(4):363–373, August 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ritchie:1991:RR

[187] Burke Ritchie. Radiation reaction. Foundations of Physics Letters, 4 (4):375–384, August 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Jones:1991:SDC

[188] Martin R. Jones. Some difficulties for Clifton, redhead, and Butterfield’s recent proof of nonlocality. Foundations of Physics Letters, 4(4):385– 393, August 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Clifton:1991:SLR

[189] Robert K. Clifton, Michael L. G. Redhead, and Jeremy N. Butterfield. A second look at a recent algebraic proof of nonlocality. Foundations of Physics Letters, 4(4):395–403, August 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1991:ICBb

[190] Anonymous. International conference on Bell’s theorem and the foundations of modern physics. Foundations of Physics Letters, 4(4):405, August 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Woodward:1991:MMT

[191] James F. Woodward. Measurements of a Machian transient mass fluctuation. Foundations of Physics Letters, 4(5):407–423, October 1991. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Mann:1991:SBH

[192] R. B. Mann. The simplest black holes. Foundations of Physics Letters, 4 (5):425–449, October 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Home:1991:CCH

[193] D. Home and S. Sengupta. A comment on “Critique of Home and Sengupta’s derivation of a Bell inequality”. Foundations of Physics Letters, 4(5):451–454, October 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See [129].

Elby:1991:RHH

[194] Andrew Elby. Reply: how is home and Sengupta’s noncontextuality condition related to locality? Foundations of Physics Letters, 4(5):455–457, October 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Fryberger:1991:CSE

[195] David Fryberger. Coulomb scattering of an electron by a monopole. Foundations of Physics Letters, 4(5):459–464, October 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Steeb:1991:HND

[196] W.-H. Steeb, N. Euler, and P. Mulser. On a hierarchy of nonlinear dynamical systems and Painlevé test. Foundations of Physics Letters, 4 (5):465–469, October 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sypel:1991:CQF

[197] R. Sypel. Comment on “Quantum formalism via signal analysis”. Foundations of Physics Letters, 4(5):471–474, October 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See response [198].

Kannenberg:1991:RCQ

[198] L. Kannenberg. Response to “Comment on “Quantum formalism via signal analysis”’. Foundations of Physics Letters, 4(5):475–477, October 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See [197].

Stuart:1991:NFP

[199] C. I. J. M. Stuart. Nonlocality as a feature of the physical world. Foundations of Physics Letters, 4(5):479–492, October 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Stuart:1991:FRR

[200] C. I. J. M. Stuart. Factorization in retrocausality: a response to O. Costa de Beauregard. Foundations of Physics Letters, 4(5):493–497, October 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Graneau:1991:NAI

[201] Peter Graneau. Nonlocal action in the induction motor. Foundations of Physics Letters, 4(5):499–506, October 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Spiller:1991:ADS

[202] T. P. Spiller, P. S. Spencer, T. D. Clark, J. F. Ralph, H. Prace, R. J. Prance, and A. Clippingdale. Adding dissipation to the Schrödinger equation from the quantum-potential viewpoint. Foundations of Physics Letters, 4(6):507–521, December 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Brandt:1991:SST

[203] Howard E. Brandt. Structure of spacetime tangent bundle. Foundations of Physics Letters, 4(6):523–536, December 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wei:1991:MSC

[204] Jiansu Wei and W. E. Baylis. Monopoles without strings: a conflict between the one-photon condition and duality invariance. Foundations of Physics Letters, 4(6):537–556, December 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cannata:1991:CCMa

[205] F. Cannata and L. Ferrari. Canonical conjugate momentum of discrete label operators in quantum mechanics i: formalism. Foundations of Physics Letters, 4(6):557–568, December 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cannata:1991:CCMb

[206] F. Cannata and L. Ferrari. Canonical conjugate momentum of discrete label operators in quantum mechanics II: formalism. Foundations of Physics Letters, 4(6):569–579, December 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Stone:1991:NSC

[207] Anthony P. Stone. Non-standard clock synchronization in special relativity and the hypothetical ether frame. Foundations of Physics Letters, 4(6): 581–591, December 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Vandyck:1991:RTP

[208] M. A. Vandyck. A remark on the twin “paradox”. Foundations of Physics Letters, 4(6):593–600, December 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Treder:1991:QGJ

[209] Hans-Jürgen Treder. Quantum gravity and Jordan’s non-associative algebras. Foundations of Physics Letters, 4(6):601–603, December 1991. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gudder:1992:QPM

[210] S. Gudder. Quantum probability model for spin. Foundations of Physics Letters, 5(1):1–13, February 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gregor:1992:KCE

[211] Malcolm H. Mac Gregor. KeV channeling effects in the Mott scattering of electrons and positrons. Foundations of Physics Letters, 5(1):15–23, February 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Barranco:1992:EER

[212] Antonio V. Barranco and Humberto M. França. Einstein–Ehrenfest’s radiation theory and Compton–Debye’s kinematics. Foundations of Physics Letters, 5(1):25–42, February 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Brandt:1992:RCS

[213] Howard E. Brandt. Riemann curvature scalar of spacetime tangent bundle. Foundations of Physics Letters, 5(1):43–55, February 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Pezzaglia:1992:CAG

[214] William M. Pezzaglia, Jr. Clifford algebra geometric-multispinor particles and multivector-current gauge fields. Foundations of Physics Letters, 5 (1):57–62, February 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Collins:1992:QMC

[215] R. Eugene Collins. Quantum mechanics as a classical diffusion process. Foundations of Physics Letters, 5(1):63–69, February 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Squires:1992:ACB

[216] Euan J. Squires. An apparent conflict between the de Broglie–Bohm model and orthodoxy in quantum cosmology. Foundations of Physics Letters, 5 (1):71–75, February 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wesley:1992:PMT

[217] J. P. Wesley. Proposal to measure terrestrial Bradley aberration. Foundations of Physics Letters, 5(1):77–82, February 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Shannon:1992:UF

[218] Darwin L. Shannon. Unification of the forces. Foundations of Physics Letters, 5(1):83–101, February 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Viola:1992:AIS

[219] G. E. Viola. Announcement: International symposium. Foundations of Physics Letters, 5(1):103, February 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Stairs:1992:AFS

[220] Allen Stairs. On Arthur Fine’s solution to the measurement problem. Foundations of Physics Letters, 5(2):105–124, April 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Fine:1992:RMP

[221] Arthur Fine. Resolving the measurement problem: reply to Allen stairs. Foundations of Physics Letters, 5(2):125–139, April 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Miron:1992:NAL

[222] Radu Miron, Radu Rosca, Mihai Anastasiei, and Klaus Buchner. New aspects of Lagrangian relativity. Foundations of Physics Letters, 5(2):

141–171, April 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hogarth:1992:DGR

[223] Mark L. Hogarth. Does general relativity allow an observer to view an eternity in a finite time? Foundations of Physics Letters, 5(2):173–181, April 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

deSabbata:1992:MTT

[224] Venzo de Sabbata and C. Sivaram. A minimal time and time-temperature uncertainty principle. Foundations of Physics Letters, 5(2):183–189, April 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Lavenda:1992:QCI

[225] B. H. Lavenda and J. Dunning-Davies. Qualms concerning the inflationary scenario. Foundations of Physics Letters, 5(2):191–196, April 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1992:TIMa

[226] Anonymous. Twelfth international MaxEnt workshop. Foundations of Physics Letters, 5(2):197, April 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anderson:1992:DCS

[227] S. J. Ronald Anderson and Geoffrey E. Stedman. Distance and the conventionality of simultaneity in special relativity. Foundations of Physics Letters, 5(3):199–220, June 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Brandt:1992:FST

[228] Howard E. Brandt. Finsler-spacetime tangent bundle. Foundations of Physics Letters, 5(3):221–248, June 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Czachor:1992:CMS

[229] Marek Czachor. On classical models of spin. Foundations of Physics Letters, 5(3):249–264, June 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kobayashi:1992:WFC

[230] Tsunehiro Kobayashi. Wave-function collapse induced by thermal fluctuations of detector variables. statistical collapse of wave functions. Foundations of Physics Letters, 5(3):265–278, June 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Squires:1992:HMW

[231] Euan J. Squires. History and many-worlds quantum theory. Foundations of Physics Letters, 5(3):279–290, June 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

deBeauregard:1992:IRC

[232] O. Costa de Beauregard. Intersubjectivity, relativistic covariance, and conditionals: response to C. I. J. M. Stuart. Foundations of Physics Letters, 5(3):291–294, June 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1992:TIMb

[233] Anonymous. Twelfth International MaxEnt Workshop. Foundations of Physics Letters, 5(3):295, June 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ban:1992:RPS

[234] Masashi Ban. Relaxation of physical systems in relative-number state representation. Foundations of Physics Letters, 5(4):297–313, August 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Brandt:1992:KST

[235] Howard E. Brandt. Kähler spacetime tangent bundle. Foundations of Physics Letters, 5(4):315–336, August 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Friedman:1992:HBS

[236] Yaakov Friedman and Ari Naimark. The homogeneity of the ball in R^{3} and special relativity. Foundations of Physics Letters, 5(4):337–354, August 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cormier-Delanoue:1992:PEW

[237] C. Cormier-Delanoue. Proposed experiments on the wave-particle duality of light. Foundations of Physics Letters, 5(4):355–364, August 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Xu:1992:NHO

[238] Chongming Xu and G. F. R. Ellis. A new higher-order gravitational field equation with L = R + ΩR_{μν}R^{μν}/R + τR_{αβγδ}R^{αβγδ}/R. Foundations of Physics Letters, 5(4):365–371, August 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dotson:1992:NIJ

[239] Allen C. Dotson. Nonlocal interactions: a Jammerian analysis. Foundations of Physics Letters, 5(4):373–382, August 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Finkelstein:1992:CCS

[240] J. Finkelstein. Covariant collapse of the state vector and realism. Foundations of Physics Letters, 5(4):383–392, August 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

DeCaro:1992:QMV

[241] Liberato De Caro and Augusto Garuccio. Quantum mechanics versus local realism: the importance of the wave function normalization. Foundations of Physics Letters, 5(4):393–398, August 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Benatti:1992:BLI

[242] F. Benatti and T. Weber. Bell-like inequality and spontaneous reduction processes. Foundations of Physics Letters, 5(5):399–423, October 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Woodward:1992:SAW

[243] James F. Woodward. A stationary apparent weight shift from a transient Machian mass fluctuation. Foundations of Physics Letters, 5(5):425–442, October 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Mocanu:1992:RVC

[244] Constantin I. Mocanu. On the relativistic velocity composition paradox and the Thomas rotation. Foundations of Physics Letters, 5(5):443–456, October 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Lavenda:1992:DNC

[245] B. H. Lavenda and A. Florio. Derivation of a new class of distributions for the microcanonical ensemble. Foundations of Physics Letters, 5(5):457– 467, October 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Lepore:1992:HDB

[246] Vito Luigi Lepore. Homodyne detection of de Broglie waves. Foundations of Physics Letters, 5(5):469–478, October 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Scharf:1992:CEC

[247] G. Scharf and W. F. Wreszinski. On the Casimir effect without cutoff. Foundations of Physics Letters, 5(5):479–487, October 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

deBeauregard:1992:TEC

[248] O. Costa de Beauregard. Timing in EPR correlations. Foundations of Physics Letters, 5(5):489–491, October 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Macdonald:1992:MWD

[249] Alan L. Macdonald. On the Marzke–Wheeler and Desloge constructions. Foundations of Physics Letters, 5(5):493–494, October 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1992:SFMa

[250] Anonymous. Symposium on the foundations of modern physics 1993. Foundations of Physics Letters, 5(5):495–497, October 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Rimini:1992:FRT

[251] Alberto Rimini and Marco Roncadelli. A framework for a relativistic theory of state reduction. Foundations of Physics Letters, 5(6):499–515, December 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Arntzenius:1992:AIP

[252] Frank Arntzenius. Apparatus independence in proofs of non-locality. Foundations of Physics Letters, 5(6):517–525, December 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

deMuynck:1992:BIJ

[253] W. M. de Muynck, W. De Baere, and H. Martens. The Bell inequalities and the joint measurement of incompatible observables. Foundations of Physics Letters, 5(6):527–534, December 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Coleman:1992:ARC

[254] Robert Coleman and Herbert Korté. On attempts to rescue the conventionality thesis of distant simultaneity in str. Foundations of Physics Letters, 5(6):535–571, December 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Marshall:1992:IQO

[255] Trevor W. Marshall and Emilio Santos. Interpretation of quantum optics based upon positive Wigner functions. Foundations of Physics Letters, 5 (6):573–578, December 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

deSabbata:1992:CS

[256] Venzo de Sabbata and C. Sivaram. Curvature from spin. Foundations of Physics Letters, 5(6):579–584, December 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ragusa:1992:GMW

[257] Silvestre Ragusa. Gravitation with a modified Weber force. Foundations of Physics Letters, 5(6):585–589, December 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Muller:1992:PR

[258] F. Archibald Muller. On the principle of relativity. Foundations of Physics Letters, 5(6):591–595, December 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wesley:1992:WPF

[259] J. P. Wesley. Weber potential from finite velocity of action? Foundations of Physics Letters, 5(6):597–604, December 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1992:SFMb

[260] Anonymous. Symposium on the foundations of modern physics 1993. Foundations of Physics Letters, 5(6):605–607, December 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Durr:1993:DJB

[261] Detlef Dürr, Sheldon Goldstein, and Nino Zanghi. David Joseph Bohm: 1917–1992. Foundations of Physics Letters, 6(1):1–4, February 1993. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Elby:1993:WMI

[262] Andrew Elby. Why “modal” interpretations of quantum mechanics don’t solve the measurement problem. Foundations of Physics Letters, 6(1):5– 19, February 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Bub:1993:MIA

[263] Jeffrey Bub. Measurement: It ain’t over till it’s over. Foundations of Physics Letters, 6(1):21–35, February 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Healey:1993:WEP

[264] Richard Healey. Why error-prone quantum measurements have outcomes. Foundations of Physics Letters, 6(1):37–54, February 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Andaas:1993:QTQ

[265] H. E. Andås and O. K. Gjøtterud. Quantum theory and questions of reality and completeness. Foundations of Physics Letters, 6(1):55–74, February 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Rueda:1993:SEPa

[266] A. Rueda. Stochastic electrodynamics with particle structure part i: Zeropoint induced Brownian behavior. Foundations of Physics Letters, 6(1): 75–108, February 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Benatti:1993:HBL

[267] F. Benatti and T. Weber. Histories, Bell-like inequalities, and QMSL evolution. Foundations of Physics Letters, 6(2):109–118, April 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Levine:1993:RFF

[268] Robert Y. Levine. Relativistic field formulation of simultaneous quantum measurement. Foundations of Physics Letters, 6(2):119–137, April 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Rueda:1993:SEPb

[269] A. Rueda. Stochastic electrodynamics with particle structure. Part II — Towards a zero-point induced wave behavior. Foundations of Physics Letters, 6(2):139–166, April 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Urbanowski:1993:PMM

[270] K. Urbanowski. A proposal for a model of many-successive measuring process. Foundations of Physics Letters, 6(2):167–189, April 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Szabo:1993:RMB

[271] László E. Szabó. On the real meaning of Bell’s theorem. Foundations of Physics Letters, 6(2):191–200, April 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

deSabbata:1993:UCA

[272] Venzo de Sabbata and C. Sivaram. Universality of charge arising from torsion. Foundations of Physics Letters, 6(2):201–206, April 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wood:1993:GFC

[273] W. R. Wood and G. Papini. A geometric formulation of the causal interpretation of quantum mechanics. Foundations of Physics Letters, 6(3): 207–223, June 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Risco-Delgado:1993:VDP

[274] R. Risco-Delgado and F. Selleri. Variable detection probability in twophoton interference. Foundations of Physics Letters, 6(3):225–232, June 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Woodward:1993:MAM

[275] James F. Woodward. Maximal acceleration, Mach’s principle, and the mass of the electron. Foundations of Physics Letters, 6(3):233–244, June 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Brandt:1993:CST

[276] Howard E. Brandt. Complex spacetime tangent bundle. Foundations of Physics Letters, 6(3):245–255, June 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Huang:1993:LCE

[277] Young-Sea Huang. Has the Lorentz-covariant electromagnetic force law been directly tested experimentally? Foundations of Physics Letters, 6(3):

257–274, June 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Woo:1993:CHR

[278] C. H. Woo. Consistent histories, robust histories, and verifiable histories. Foundations of Physics Letters, 6(3):275–281, June 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Elizalde:1993:NQH

[279] E. Elizalde. A note on the quantum Hamilton–Jacobi formalism. Foundations of Physics Letters, 6(3):283–288, June 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dunning-Davies:1993:CSS

[280] J. Dunning-Davies. Concavity, superadditivity and the second law. Foundations of Physics Letters, 6(3):289–295, June 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1993:FFPa

[281] Anonymous. Frontiers of fundamental physics. Foundations of Physics Letters, 6(3):297, June 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1993:ISFa

[282] Anonymous. International symposium on fundamental problems in quantum physics. Foundations of Physics Letters, 6(3):298, June 1993. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Albert:1993:NIM

[283] David Albert and Barry Loewer. Non-ideal measurements. Foundations of Physics Letters, 6(4):297–305, August 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Healey:1993:MQI

[284] Richard Healey. Measurement and quantum indeterminateness. Foundations of Physics Letters, 6(4):307–316, August 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ben-Aryeh:1993:LVS

[285] Y. Ben-Aryeh. Locality is not violated for the singlet spin state. Foundations of Physics Letters, 6(4):317–325, August 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Olsen:1993:CME

[286] Elwood T. Olsen. Classical mechanics and entropy. Foundations of Physics Letters, 6(4):327–337, August 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Brandt:1993:STB

[287] Howard E. Brandt. Spacetime tangent bundle with torsion. Foundations of Physics Letters, 6(4):339–369, August 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Torr:1993:GEC

[288] Douglas G. Torr and Ning Li. Gravitoelectric-electric coupling via superconductivity. Foundations of Physics Letters, 6(4):371–383, August 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See comments [580].

Hemmo:1993:BRP

[289] Meir Hemmo. Book review: The philosophy of quantum mechanics: an interactive interpretation, by Richard Healey. Cambridge University Press, 1989, Cambridge, United Kingdom, xiv + 270 pp., $42.95 (hardcover). Foundations of Physics Letters, 6(4):385–393, August 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1993:FFPb

[290] Anonymous. Frontiers of fundamental physics international conference. Foundations of Physics Letters, 6(4):395, August 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1993:ISFb

[291] Anonymous. International symposium on fundamental problems in quantum physics. Foundations of Physics Letters, 6(4):397, August 1993. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Risco-Delgado:1993:VDA

[292] R. Risco-Delgado. The variable detection approach: a wave particle model. Foundations of Physics Letters, 6(5):399–428, October 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Collins:1993:SBP

[293] R. Eugene Collins. Statistical basis for physical laws; non-relativistic theory. Foundations of Physics Letters, 6(5):429–449, October 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ritchie:1993:RRI

[294] Burke Ritchie. Radiation reaction. II. Foundations of Physics Letters, 6 (5):451–468, October 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dowe:1993:ARC

[295] Phil Dowe. The anti-realism of Costa de Beauregard. Foundations of Physics Letters, 6(5):469–475, October 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

deBeauregard:1993:ARA

[296] O. Costa de Beauregard. “Anti-realism”: answer to Phil Dowe. Foundations of Physics Letters, 6(5):477–479, October 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See [295].

Braun:1993:DWM

[297] M. A. Braun and K. Urbanowski. On the decay width of multiply measured unstable state. Foundations of Physics Letters, 6(5):481–490, October 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Graneau:1993:DBN

[298] Peter Graneau. The difference between Newtonian and relativistic forces. Foundations of Physics Letters, 6(5):491–500, October 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hyvonen:1993:SFM

[299] Tapio Hyvönen. Symposium on the foundations of modern physics 1994. Foundations of Physics Letters, 6(5):501, October 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anandan:1993:PMQ

[300] J. Anandan. Protective measurement and quantum reality. Foundations of Physics Letters, 6(6):503–532, December 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cassinelli:1993:CVM

[301] Gianni Cassinelli and Pekka Lahti. The Copenhagen variant of the modal interpretation and quantum theory of measurement. Foundations of Physics Letters, 6(6):533–544, December 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Mashhoon:1993:QTO

[302] Bahram Mashhoon. Quantum theory and the origin of inertia. Foundations of Physics Letters, 6(6):545–559, December 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

deSabbata:1993:LFS

[303] Venzo de Sabbata and C. Sivaram. On limiting field strengths in gravitation. Foundations of Physics Letters, 6(6):561–570, December 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Tegmark:1993:AWF

[304] Max Tegmark. Apparent wave function collapse caused by scattering. Foundations of Physics Letters, 6(6):571–590, December 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wyss:1993:SGI

[305] Walter Wyss. Scaling gauge invariance and constancy of the speed of light. Foundations of Physics Letters, 6(6):591–596, December 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Adomian:1993:BP

[306] George Adomian. The N-body problem. Foundations of Physics Letters, 6 (6):597–602, December 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Adomian:1993:CNR

[307] George Adomian. A class of nonlinear relativistic partial differential equations in elementary particle theory. Foundations of Physics Letters, 6(6): 603–605, December 1993. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Monaco:1994:DSE

[308] R. E. L. Mónaco and G. G. Cabrera. The double slit experiment, the Aharonov–Bohm effect, and a new quantum potential theory: Towards a realistic interpretation of quantum mechanics. Foundations of Physics Letters, 7(1):1–20, February 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Das:1994:DPSa

[309] A. Das and P. Smoczynski. Discrete phase space. I. Finite difference operators and lattice Schrödinger wave equation. Foundations of Physics Letters, 7(1):21–38, February 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Collins:1994:MBP

[310] R. Eugene Collins. The mathematical basis of physical laws: Relativistic mechanics, quantum mechanics, the Lorentz force, Maxwell’s equations and non-electromagnetic forces for spin zero particles. Foundations of Physics Letters, 7(1):39–58, February 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Woodward:1994:AEE

[311] James F. Woodward. ADM electrons and the equivalence principle. Foundations of Physics Letters, 7(1):59–65, February 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1994:PBL

[312] M. W. Evans. The photomagneton b^{(3)} and longitudinal ghost field b^{(3)} of electromagnetism. Foundations of Physics Letters, 7(1):67–74, February 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Guerra:1994:ADM

[313] Marco Guerra and Antonio Sparzani. About the definition of mass in (Machian) classical mechanics. Foundations of Physics Letters, 7(1):75– 84, February 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Recami:1994:SCC

[314] Erasmo Recami and Vilson Tonin-Zanchin. The strong coupling constant: Its theoretical derivation from a geometric approach to hadron structure. Foundations of Physics Letters, 7(1):85–93, February 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wesley:1994:PEA

[315] J. P. Wesley. Pinch effect and Ampère tension to drive Bering’s pump. Foundations of Physics Letters, 7(1):95–104, February 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Benatti:1994:SRP

[316] Fabio Benatti, GianCarlo Ghirardi, and Renata Grassi. On some recent proposals for testing macrorealism versus quantum mechanics. Foundations of Physics Letters, 7(2):105–126, April 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Das:1994:DPSb

[317] A. Das and P. Smoczynski. Discrete phase space II. Relativistic lattice wave equations and divergence-free Green’s functions. Foundations of Physics Letters, 7(2):127–142, April 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

delaTorre:1994:CTQ

[318] Alberto C. de la Torre. The concept of transmission in quantum mechanics. Foundations of Physics Letters, 7(2):143–159, April 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gonella:1994:TMS

[319] Francesco Gonella. Time machine, self-consistency and the foundations of quantum mechanics. Foundations of Physics Letters, 7(2):161–166, April 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Tonin-Zanchin:1994:RLR

[320] Vilson Tonin-Zanchin, Erasmo Recami, José A. Roversi, and Luis A. Brasca-Annes. “Regge-like” relations for stable (non-evaporating) black holes. Foundations of Physics Letters, 7(2):167–179, April 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Paim:1994:CSB

[321] T. Paim and M. M. Som. A conformai, symmetry broken model for gravity. Foundations of Physics Letters, 7(2):181–190, April 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Folman:1994:SHV

[322] Ron Folman. A search for hidden variables in the domain of high energy physics. Foundations of Physics Letters, 7(2):191–200, April 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ferrero:1994:BDB

[323] M. Ferrero and J. M. Getino. Bounds to the difference between the predictions of stochastic and quantum optics in the optical test of bell inequalities. Foundations of Physics Letters, 7(2):201–208, April 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1994:PEC

[324] M. W. Evans. The photomagneton B̂^{(3)} and electrodynamic conservation laws. Foundations of Physics Letters, 7(2):209–217, April 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Steinberg:1994:HUI

[325] A. M. Steinberg, P. G. Kwiat, and R. Y. Chiao. Hidden and unhidden information in quantum tunneling. Foundations of Physics Letters, 7(3): 223–239, June 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kerner:1994:EAQ

[326] Edward H. Kerner. An essential ambiguity of quantum theory. Foundations of Physics Letters, 7(3):241–258, June 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Westmoreland:1994:ZAC

[327] Michael D. Westmoreland and Benjamin W. Schumacher. Zeno’s arrow and classical phase space logics. Foundations of Physics Letters, 7(3):259– 271, June 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anderson:1994:SMS

[328] Ronald Anderson and Geoffrey E. Stedman. Spatial measures in special relativity do not empirically determine simultaneity relations: a reply to Coleman and Korté. Foundations of Physics Letters, 7(3):273–283, June 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Urbanowski:1994:SPE

[329] K. Urbanowski. Strange properties of early time evolution of the projection of a state vector. Foundations of Physics Letters, 7(3):285–295, June 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Brandt:1994:STR

[330] Howard E. Brandt. Solution to torsion relations in Finsler-spacetime tangent bundle. Foundations of Physics Letters, 7(3):297–301, June 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Esposito:1994:BTM

[331] Giampiero Esposito, Hugo Morales-Técotl, and Giuseppe Pollifrone. Boundary terms for massless fermionic fields. Foundations of Physics Letters, 7(3):303–308, June 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

McCarty:1994:BR

[332] Robert C. McCarty. Book review. Foundations of Physics Letters, 7(3): 309, June 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1994:E

[333] Anonymous. Errata. Foundations of Physics Letters, 7(3):311, June 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). URL

http://link.springer.com/content/pdf/10.1007/BF02275259.pdf.

Schmeelk:1994:TDD

[334] John Schmeelk. Two-dimensional Dirac delta reconsidered. Foundations of Physics Letters, 7(4):315–332, August 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Jeffers:1994:EDE

[335] S. Jeffers and J. Sloan. An experiment to detect “empty” waves. Foundations of Physics Letters, 7(4):333–341, August 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Martens:1994:PME

[336] Hans Martens and Willem M. de Muynck. On the possibility of measuring the electron spin in an inhomogeneous magnetic field. Foundations of Physics Letters, 7(4):343–352, August 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Squires:1994:QNB

[337] Euan J. Squires. Has quantum nonlocality been experimentally verified? Foundations of Physics Letters, 7(4):353–363, August 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hokkyo:1994:RTP

[338] Noboru Hokkyo. Reinterpretation of two-photon interference experiment using advanced field. Foundations of Physics Letters, 7(4):365–371, August 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Schauer:1994:BIB

[339] Daniel L. Schauer. Bell’s inequalities between independent particles. Foundations of Physics Letters, 7(4):373–378, August 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1994:LPE

[340] M. W. Evans. The longitudinal photomagneton of electromagnetic radiation: on the absence of Faraday induction in vacuo. Foundations of Physics Letters, 7(4):379–384, August 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gruber:1994:TUS

[341] Ronald P. Gruber and Stephen M. Matthews. Is there a universal scale of time? Foundations of Physics Letters, 7(4):385–392, August 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Bueno:1994:PLH

[342] M. A. Bueno. Phipps’s Lagrangian and Hamiltonian. Foundations of Physics Letters, 7(4):393–400, August 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Roa-Neri:1994:ARR

[343] J. A. E. Roa-Neri and J. L. Jimenez. An analysis of the radiation reaction in a charged dumbbell. Foundations of Physics Letters, 7(5):403–417, October 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Deck:1994:COD

[344] R. T. Deck and N. Öztürk. Can an observable with discrete eigenvalues be conjugate to a continuum observable? Foundations of Physics Letters, 7(5):419–436, October 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1994:MEP

[345] M. W. Evans. Magnetization of an electron plasma by microwave pulses: Experiment to detect the longitudinal vacuum field B^{(3)}. Foundations of Physics Letters, 7(5):437–446, October 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Pessoa:1994:SGM

[346] Osvaldo Pessoa, Jr. Simplerealizations of generalized measurements in quantum mechanics. Foundations of Physics Letters, 7(5):447–457, October 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ben-Aryeh:1994:UBI

[347] Y. Ben-Aryeh. The use of Bell inequalities in quantum optics. Foundations of Physics Letters, 7(5):459–466, October 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1994:OAB

[348] M. W. Evans. The optical Aharonov–Bohm effect. Foundations of Physics Letters, 7(5):467–474, October 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

deSabbata:1994:GCV

[349] Venzo de Sabbata and C. Sivaram. Gravity and CP violation. Foundations of Physics Letters, 7(5):475–482, October 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Neeman:1994:CSM

[350] Yuval Ne’eman. Cosmological surrealism: More than “eternal reality” is needed. Foundations of Physics Letters, 7(5):483–488, October 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Adomian:1994:AST

[351] G. Adomian and H. J. Efinger. Analytic solutions for time-dependent Schrödinger equations with linear of nonlinear Hamiltonians. Foundations of Physics Letters, 7(5):489–491, October 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wesley:1994:OSD

[352] J. P. Wesley. Oneway Sagnac device to measure absolute velocity. Foundations of Physics Letters, 7(5):493–499, October 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Suppes:1994:DWD

[353] Patrick Suppes and J. Acacio de Barros. Diffraction with well-defined photon trajectories: a foundational analysis. Foundations of Physics Letters, 7(6):501–514, December 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Arunasalam:1994:IPR

[354] V. Arunasalam. Are the invariance principles really truly Lorentzcovariant? Foundations of Physics Letters, 7(6):515–530, December 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ribaric:1994:LRM

[355] Marijan Ribaric and Luka Sustersic. A linear relativistic model of processes with causal faster-than-light effects. Foundations of Physics Letters, 7(6):531–550, December 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hannibal:1994:QFT

[356] Ludger Hannibal. On quantum field theory with invariant parameter. Foundations of Physics Letters, 7(6):551–576, December 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1994:DVL

[357] M. W. Evans. Derivation of the vacuum longitudinal field B^{(3)} from the Dirac equation of the electron in the electromagnetic field. Foundations of Physics Letters, 7(6):577–583, December 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Adomian:1994:OC

[358] G. Adomian. The origin of chaos. Foundations of Physics Letters, 7(6): 585–589, December 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1994:FEV

[359] M. W. Evans. Forms of the Evans–Vigier field B^{(3)}. Foundations of Physics Letters, 7(6):591–600, December 1994. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Woodward:1995:MUSa

[360] James F. Woodward. Making the universe safe for historians: Time travel and the laws of physics. Foundations of Physics Letters, 8(1):1–39, February 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Feldmann:1995:NLE

[361] Michel Feldmann. New loophole for the Einstein–Podolsky-Rosen paradox. Foundations of Physics Letters, 8(1):41–53, February 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Folman:1995:EWD

[362] Ron Folman and Zeev Vager. Empty wave detecting experiments: a comment on auxiliary “hidden” assumptions. Foundations of Physics Letters, 8(1):55–61, February 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1995:NTF

[363] M. W. Evans. Noether’s theorem and the field b^{(3)}. Foundations of Physics Letters, 8(1):63–71, February 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Steeb:1995:BFS

[364] W.-H. Steeb. Bose–Fermi systems and computer algebra. Foundations of Physics Letters, 8(1):73–81, February 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1995:CBP

[365] M. W. Evans. The connection between photon mass and B^{(3)}: the Poincaré group. Foundations of Physics Letters, 8(1):83–92, February 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wesley:1995:PWP

[366] J. P. Wesley. Proposed Weber potential with absolute velocities. Foundations of Physics Letters, 8(1):93–98, February 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Adomian:1995:GKP

[367] G. Adomian. The generalized Kolmogorov–Petrovskii-Piskunov equation. Foundations of Physics Letters, 8(1):99–101, February 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1995:Aa

[368] Anonymous. Announcements. Foundations of Physics Letters, 8(1):103– 104, February 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Buffa:1995:DRE

[369] Mariamargherita Buffa, Oreste Nicrosini, and Alberto Rimini. Dissipation and reduction effects of spontaneous localization on superconducting states. Foundations of Physics Letters, 8(2):105–125, April 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Folman:1995:PTR

[370] Ron Folman and Erasmo Recami. On the phenomenology of tachyon radiation. Foundations of Physics Letters, 8(2):127–134, April 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gregor:1995:MBS

[371] Malcolm H. Mac Gregor. Model basis states for photons and “empty waves”. Foundations of Physics Letters, 8(2):135–160, April 1995. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Adomian:1995:ASL

[372] G. Adomian, R. Rach, and N. T. Shawagfeh. On the analytic solution of the Lane–Emden equation. Foundations of Physics Letters, 8(2):161– 181, April 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Lakhtakia:1995:ELG

[373] Akhlesh Lakhtakia. Is Evans’ longitudinal ghost field B^{(3)} unknowable? Foundations of Physics Letters, 8(2):183–186, April 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See reply [374].

Evans:1995:RLE

[374] M. W. Evans. Reply to A. Lakhtakia: Experimental measurement of B^{(3)}. Foundations of Physics Letters, 8(2):187–193, April 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See [373].

Bao:1995:DQC

[375] Dehai Bao and Z. Y. Zhu. On Dirac’s quantization condition. Foundations of Physics Letters, 8(2):195–201, April 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1995:Ab

[376] Anonymous. Announcement. Foundations of Physics Letters, 8(2):203, April 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Pykacz:1995:BTI

[377] Jaroslaw Pykacz and Emilio Santos. Bell-type inequalities on tensor products of quantumlogics. Foundations of Physics Letters, 8(3):205–212, June 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

DelNoce:1995:AFB

[378] Carlo Del Noce. An algorithm for finding Bell-type inequalities. Foundations of Physics Letters, 8(3):213–229, June 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dickson:1995:FBAa

[379] W. Michael Dickson. Faux-boolean algebras, classical probability, and determinism. Foundations of Physics Letters, 8(3):231–242, June 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gudder:1995:TEE

[380] Stanley Gudder. Total extensions of effect algebras. Foundations of Physics Letters, 8(3):243–252, June 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1995:NAE

[381] M. W. Evans. Non-abelian electrodynamics and the vacuum B^{(3)}. Foundations of Physics Letters, 8(3):253–259, June 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Coddens:1995:RDE

[382] Gerrit Coddens. A remark on the Dirac equation. Foundations of Physics Letters, 8(3):261–277, June 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1995:PMF

[383] M. W. Evans. Photon mass and the field B^{(3)}. Foundations of Physics Letters, 8(3):279–286, June 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

deBeauregard:1995:MR

[384] O. Costa de Beauregard. Macroscopic retrocausation. Foundations of Physics Letters, 8(3):287–291, June 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Griffiths:1995:BR

[385] Robert B. Griffiths. Book review. Foundations of Physics Letters, 8(3): 293–298, June 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wesley:1995:BR

[386] J. Paul Wesley. Book review. Foundations of Physics Letters, 8(3):299– 301, June 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Woodward:1995:MUSb

[387] James F. Woodward. Making the universe safe for historians: Time travel and the laws of physics. Foundations of Physics Letters, 8(3):303, June 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

URL http://link.springer.com/content/pdf/10.1007/BF02187355.
pdf.

Anonymous:1995:CP

[388] Anonymous. Call for papers. Foundations of Physics Letters, 8(3):305, June 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1995:AFI

[389] Anonymous. Announcement: The Fifteenth International Workshop on Maximum Entropy and Bayesian Methods 31 July–4 August 1995 St. John’s College, Santa Fe, New Mexico, USA. Foundations of Physics Letters, 8(3):307, June 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hannibal:1995:CNV

[390] Ludger Hannibal. CP non-violation in neutral K-meson decay: a nonstandard explanation. Foundations of Physics Letters, 8(4):309–326, August 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ruetsche:1995:MEA

[391] Laura Ruetsche. Measurement error and the Albert–Loewer problem. Foundations of Physics Letters, 8(4):327–344, August 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Folman:1995:QMV

[392] Ron Folman and Zeev Vager. Quantum mechanics versus realism. Foundations of Physics Letters, 8(4):345–358, August 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1995:MEP

[393] M. W. Evans. Magnetization of an electron plasma by microwave pulses: Faraday induction. Foundations of Physics Letters, 8(4):359–364, August 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Monaco:1995:RIA

[394] R. L. Monaco, R. E. Lagos, and W. A. Rodrigues, Jr. A Riemann integral approach to Feynman’s path integral. Foundations of Physics Letters, 8 (4):365–373, August 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

deSabbata:1995:TST

[395] Venzo de Sabbata and C. Sivaram. Torsion, string tension, and topological origin of charge and mass. Foundations of Physics Letters, 8(4):375–380, August 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1995:BSA

[396] M. W. Evans. The Biot–Savart–Ampère law and the vacuum field B^{(3)}. Foundations of Physics Letters, 8(4):381–388, August 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Adomian:1995:ASN

[397] G. Adomian. Analytical solution of Navier–Stokes flow of a viscous compressible fluid. Foundations of Physics Letters, 8(4):389–400, August 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dickson:1995:FBAb

[398] W. Michael Dickson. Faux-boolean algebras and classical models. Foundations of Physics Letters, 8(5):401–415, October 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Szabo:1995:QMC

[399] László E. Szabó. Is quantum mechanics compatible with a deterministic universe? two interpretations of quantum probabilities. Foundations of Physics Letters, 8(5):417–436, October 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Coecke:1995:GPE

[400] Bob Coecke. Generalization of the proof on the existence of hidden measurements to experiments with an infinite set of outcomes. Foundations of Physics Letters, 8(5):437–447, October 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Perelstein:1995:NRE

[401] Elcuno A. Perelstein and Maxim E. Perelstein. Notes on relativistic equations of spin motion. Foundations of Physics Letters, 8(5):449–457, October 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1995:OFR

[402] M. W. Evans. On the observability of the field B^{(3)}: Relativistic effects in magneto-optics. Foundations of Physics Letters, 8(5):459–466, October 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Afriat:1995:PRI

[403] Alexander Afriat. Probability relations involving several subsystems. Foundations of Physics Letters, 8(5):467–480, October 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1995:EVF

[404] M. W. Evans. The Evans–Vigier field, B^{(3)}: Derivation of the de Broglie matter–wave equation from the Hamilton–Jacobi equation. Foundations of Physics Letters, 8(5):481–489, October 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Yao:1995:CSA

[405] Carl Yao. Can statistical average be affected by its observer? Foundations of Physics Letters, 8(5):491–495, October 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Reeder:1995:PCH

[406] Nick Reeder. Property composition in Healey’s interpretation of quantum mechanics. Foundations of Physics Letters, 8(6):497–522, December 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Svozil:1995:CUP

[407] K. Svozil. Consistent use of paradoxes in deriving constraints on the dynamics of physical systems and of no-go theorems. Foundations of Physics Letters, 8(6):523–535, December 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Burakovsky:1995:TRE

[408] L. Burakovsky, L. P. Horwitz, and W. C. Schieve. Towards a realistic equation of state of strongly interacting matter. Foundations of Physics Letters, 8(6):537–548, December 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Muller:1995:FH

[409] F. A. Muller. Fixing a hole. Foundations of Physics Letters, 8(6):549–562, December 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1995:RCF

[410] M. W. Evans. Reply to criticisms of the B^{(3)} field. Foundations of Physics Letters, 8(6):563–573, December 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Burghardt:1995:SGO

[411] Rainer Burghardt. Schwarzschild geometry, once more. Foundations of Physics Letters, 8(6):575–582, December 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Adomian:1995:NAS

[412] G. Adomian. A new approach to solution of the Maxwell equations. Foundations of Physics Letters, 8(6):583–587, December 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Squires:1995:ETE

[413] Euan J. Squires. An experiment to test an explanation of a possible violation of quantum theory. Foundations of Physics Letters, 8(6):589– 591, December 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1995:IIS

[414] Anonymous. II International symposium on fundamental problems in quantum physics, Oviedo, Spain 21–26 July 1996. Foundations of Physics Letters, 8(6):593, December 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Woodward:1996:KT

[415] James F. Woodward. Killing time. Foundations of Physics Letters, 9(1):1– 23, February 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Tegmark:1996:DUF

[416] Max Tegmark. Does the universe in fact contain almost no information? Foundations of Physics Letters, 9(1):25–41, February 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Selleri:1996:NOW

[417] F. Selleri. Noninvariant one-way velocity of light and particle collisions. Foundations of Physics Letters, 9(1):43–60, February 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1996:PEV

[418] M. W. Evans. Proof of the Evans–Vigier field from the Dirac equation of the fermion in the classical field: Reply to Rikken. Foundations of Physics Letters, 9(1):61–66, February 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

DeBaere:1996:HAP

[419] W. De Baere. How to avoid Peres’ “incompatible results of quantum measurements”. Foundations of Physics Letters, 9(1):67–78, February 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Carazza:1996:CLM

[420] B. Carazza. Considerations on localization of macroscopic bodies. Foundations of Physics Letters, 9(1):79–90, February 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Schafir:1996:UPE

[421] R. L. Schafir. Unprovability of a “precognition” effect in quantum mechanics. Foundations of Physics Letters, 9(1):91–101, February 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1996:Aa

[422] Anonymous. Announcement. Foundations of Physics Letters, 9(1):103, February 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Israelit:1996:NR

[423] Mark Israelit. Nathan Rosen: 1909–1995. Foundations of Physics Letters, 9(2):105–108, April 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Stocks:1996:VRN

[424] A. C. Stocks and M. L. G. Redhead. A value rule for non-maximal observables. Foundations of Physics Letters, 9(2):109–119, April 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Perelstein:1996:PCD

[425] Elcouno A. Perelstein and Maxim E. Perelstein. On possibility of classical description of spin motion in inhomogeneous electromagnetic fields. Foundations of Physics Letters, 9(2):121–131, April 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

deVries:1996:DAT

[426] D. K. de Vries and W. M. de Muynck. A dynamical approach to time dilation and length contraction. Foundations of Physics Letters, 9(2):133– 144, April 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Stephens:1996:IFT

[427] G. Stephens. On inertial frames in the theory of quantised fields. Foundations of Physics Letters, 9(2):145–150, April 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

deBeauregard:1996:EIP

[428] O. Costa de Beauregard. Electrodynamics of interacting point charges: Excellence of the 1865 Clausius approximation. Foundations of Physics Letters, 9(2):151–156, April 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

vonBorzeszkowski:1996:QGR

[429] H.-H. von Borzeszkowski, V. de Sabbata, C. Sivaram, and H.-J. Treder. On the quantization of general relativity in anholonomic variables. Foundations of Physics Letters, 9(2):157–164, April 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Goy:1996:AQE

[430] François Goy. Aberration and the question of equivalence of some ether theories to special relativity. Foundations of Physics Letters, 9(2):165– 174, April 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1996:OND

[431] M. W. Evans. Optical NMR from the Dirac equation: a reply to Buckingham and Parlett. Foundations of Physics Letters, 9(2):175–181, April 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

vanEnk:1996:TSM

[432] S. J. van Enk. Is there a static magnetic field of the photon? Foundations of Physics Letters, 9(2):183–190, April 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1996:OPM

[433] M. W. Evans. One photon and macroscopic b cyclic equations: Reply to van enk. Foundations of Physics Letters, 9(2):191–204, April 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1996:AMI

[434] Anonymous. Announcement: The Modal Interpretation of Quantum Mechanics 12-14 June 1996, Utrecht University, The Netherlands. Foundations of Physics Letters, 9(2):205, April 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1996:AIS

[435] Anonymous. Announcement: Il international symposium on fundamental problems in quantum physics oviedo, Spain 21-26 July 1996. Foundations of Physics Letters, 9(2):207–208, April 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1996:ADS

[436] Anonymous. Announcement: Descartes and scientific thought. Foundations of Physics Letters, 9(2):209, April 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1996:API

[437] Anonymous. Announcement: Physical interpretations of relativity theory: British Society for the Philosophy of Science 6=-9 September 1996, London. Foundations of Physics Letters, 9(2):211, April 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Suppes:1996:PTC

[438] Patrick Suppes, Adonai S. Sant’Anna, and J. Acacio de Barros. A particle theory of the Casimir effect. Foundations of Physics Letters, 9(3):213– 223, June 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hacyan:1996:QSS

[439] Shahen Hacyan. The quantum sling and the Schrödinger cat. Foundations of Physics Letters, 9(3):225–233, June 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Marciak-Kozlowska:1996:DTE

[440] Janina Marciak-Kozlowska and Miroslaw Kozlowski. Discretization of the thermal excitation in highly excited matter. Foundations of Physics Letters, 9(3):235–245, June 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Woodward:1996:LTM

[441] James F. Woodward. A laboratory test of Mach’s principle and strong-field relativistic gravity. Foundations of Physics Letters, 9(3):247–293, June 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Zecca:1996:NSA

[442] Antonio Zecca. Non-standard aspects of Minkowski causal logic. Foundations of Physics Letters, 9(3):295–300, June 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Adomian:1996:NPS

[443] G. Adomian. A non-perturbative solution of N-body dynamics. Foundations of Physics Letters, 9(3):301–308, June 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1996:Ab

[444] Anonymous. Announcements. Foundations of Physics Letters, 9(3):309– 312, June 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ghirardi:1996:PER

[445] GianCarlo Ghirardi. Properties and events in a relativistic context: Revisiting the dynamical reduction program. Foundations of Physics Letters, 9(4):313–355, August 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Thompson:1996:CBI

[446] Caroline H. Thompson. The chaotic ball: An intuitive analogy for EPR experiments. Foundations of Physics Letters, 9(4):357–382, August 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sherry:1996:GIQ

[447] G. C. Sherry. A geometrical interpretation of the quantum potential in terms of the Hamilton–Jacobi theory in the calculus of variations. Foundations of Physics Letters, 9(4):383–395, August 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1996:UEG

[448] M. W. Evans. Unification of electromagnetism and gravitation: the equations of electrodynamics derived from the Riemann tensor in general relativity. Foundations of Physics Letters, 9(4):397–405, August 1996. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Adomian:1996:DSG

[449] G. Adomian. The dissipative sine-Gordon equation. Foundations of Physics Letters, 9(4):407–410, August 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

dePolavieja:1996:NQC

[450] Gonzalo Garcia de Polavieja. Nonstatistical quantum-classical correspondence in phase space. Foundations of Physics Letters, 9(5):411–424, October 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Suarez:1996:PII

[451] Mauricio Suárez. On the physical impossibility of ideal quantum measurements. Foundations of Physics Letters, 9(5):425–435, October 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Herbut:1996:ROT

[452] F. Herbut. On retroactive occurrence and twin events in quantum mechanics. Foundations of Physics Letters, 9(5):437–446, October 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

McKeon:1996:HDD

[453] D. G. C. McKeon and G. N. Ord. On how the (1 + 1)-dimensional Dirac equation arises in classical physics. Foundations of Physics Letters, 9 (5):447–456, October 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kinzer:1996:WFM

[454] E. T. Kinzer and J. Fukai. Weber’s force and Maxwell’s equations. Foundations of Physics Letters, 9(5):457–461, October 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1996:BFL

[455] M. W. Evans. The b^{(3)} field as a link between gravitation and electromagnetism in the vacuum. Foundations of Physics Letters, 9(5):463–473, October 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Shelupsky:1996:PET

[456] David Shelupsky. The principle of equivalence and theories of gravity. Foundations of Physics Letters, 9(5):475–487, October 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Floyd:1996:WWG

[457] Edward R. Floyd. Where and why the generalized Hamilton–Jacobi representation describes microstates of the Schrödinger wave function. Foundations of Physics Letters, 9(5):489–497, October 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wolf:1996:SMP

[458] Carl Wolf. The statistical mechanics of particles obeying discrete spatial quantum theory. Foundations of Physics Letters, 9(5):499–504, October 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Peruzzi:1996:QMF

[459] Giulio Peruzzi and Alberto Rimini. Quantum measurement in a family of hidden-variable theories. Foundations of Physics Letters, 9(6):505–519, December 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Moussa:1996:EGM

[460] M. H. Y. Moussa and A. O. Caldeira. Extension of the GRWP model to translation symmetry breaking processes. Foundations of Physics Letters, 9(6):521–534, December 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Roa-Neri:1996:DCR

[461] J. A. E. Roa-Neri, L. Nasser, and J. L. Jiménez. Dual charges and radiation reaction. Foundations of Physics Letters, 9(6):535–550, December 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Suppes:1996:VBI

[462] Patrick Suppes, J. Acacio de Barros, and Adonai S. Sant’Anna. Violation of Bell’s inequalities with a local theory of photons. Foundations of Physics Letters, 9(6):551–560, December 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Burakovsky:1996:MSM

[463] L. Burakovsky and L. P. Horwitz. Mass spectrum of a meson nonet is linear. Foundations of Physics Letters, 9(6):561–574, December 1996. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ferber:1996:MLW

[464] R. Ferber. A missing link: What is behind de Broglie’s “periodic phenomenon”? Foundations of Physics Letters, 9(6):575–586, December 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1996:RCC

[465] M. W. Evans and S. Jeffers. Reply to comment by E. Comay on the longitudinal magnetic field of circularly polarized electromagnetic waves. Foundations of Physics Letters, 9(6):587–593, December 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1996:CP

[466] Anonymous. Call for papers. Foundations of Physics Letters, 9(6):595, December 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

delaPena:1997:TTO

[467] L. de la Peña and A. M. Cetto. Tunneling times for opaque barriers. Foundations of Physics Letters, 10(1):1–15, February 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Goy:1997:TRP

[468] F. Goy and F. Selleri. Time on a rotating platform. Foundations of Physics Letters, 10(1):17–29, February 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Munera:1997:NEN

[469] Héctor A. Múnera and Octavio Guzmán. New explicit nonperiodic solutions of the homogeneous wave equation. Foundations of Physics Letters, 10(1):31–41, February 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hadley:1997:LQM

[470] Mark J. Hadley. The logic of quantum mechanics derived from classical general relativity. Foundations of Physics Letters, 10(1):43–60, February 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Burakovsky:1997:MSB

[471] L. Burakovsky and L. P. Horwitz. Mass spectrum of a baryon octet is linear. Foundations of Physics Letters, 10(1):61–71, February 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Selleri:1997:NOW

[472] F. Selleri. Noninvariant one-way speed of light and locally equivalent reference frames. Foundations of Physics Letters, 10(1):73–83, February 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wyss:1997:MGC

[473] Walter Wyss. The meaning of general covariance of Einstein’s theory of gravity. Foundations of Physics Letters, 10(1):85–88, February 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Chubykalo:1997:PFL

[474] A. E. Chubykalo, M. W. Evans, and R. Smirnov-Rueda. The phase-free, longitudinal magnetic component of vacuum electromagnetism. Foundations of Physics Letters, 10(1):93–98, February 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Bar:1997:IHV

[475] D. Bar. Identification of hidden variables through Everett’s formalism. Foundations of Physics Letters, 10(1):99–103, February 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Liboff:1997:MDS

[476] Richard L. Liboff. Maxwell’s demon and the Second Law of Thermodynamics. Foundations of Physics Letters, 10(1):??, February 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

tHooft:1997:QMB

[477] G. ’t Hooft. Quantum-mechanical behaviour in a deterministic model. Foundations of Physics Letters, 10(2):105–111, April 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Busch:1997:RSC

[478] Paul Busch and Pekka J. Lahti. Remarks on separability of compound quantum systems and time reversal. Foundations of Physics Letters, 10 (2):113–117, April 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

DeBaere:1997:DQF

[479] W. De Baere. Does the quantum formalism allow for the observability of empty quantum waves? Foundations of Physics Letters, 10(2):119– 130, April 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Burakovsky:1997:GMO

[480] L. Burakovsky and L. P. Horwitz. Gell-Mann–Okubo mass formula for SU (4) meson hexadecuplet. Foundations of Physics Letters, 10(2):131– 141, April 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Nasser:1997:STD

[481] L. Nasser, J. A. E. Roa-Neri, and J. L. Jiménez. Self-torque on a dual charged dumbbell. Foundations of Physics Letters, 10(2):143–151, April 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Woodward:1997:TFC

[482] James F. Woodward. Twists of fate: Can we make traversable wormholes in spacetime? Foundations of Physics Letters, 10(2):153–181, April 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Bedard:1997:RQT

[483] Katherine Bedard. Reply to “quantum tunneling times: a crucial test for the causal program?”. Foundations of Physics Letters, 10(2):183–187, April 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wesley:1997:ECE

[484] J. P. Wesley. Empirically correct electrodynamics. Foundations of Physics Letters, 10(2):189–204, April 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1997:Aa

[485] Anonymous. Announcement. Foundations of Physics Letters, 10(2):205, April 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ryff:1997:SBE

[486] Luiz Carlos Ryff. The strange behavior of entangled photons. Foundations of Physics Letters, 10(3):207–220, June 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Fain:1997:QMP

[487] Benjamin Fain. Quantum measurement and pure dephasing. Foundations of Physics Letters, 10(3):221–233, June 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sancho:1997:WPP

[488] Pedro Sancho. Wave and particle pictures of correlated systems. Foundations of Physics Letters, 10(3):235–244, June 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Comay:1997:RVL

[489] E. Comay. Relativity versus the longitudinal magnetic field of the photon. Foundations of Physics Letters, 10(3):245–254, June 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See reply [490].

Evans:1997:RCR

[490] M. W. Evans. Reply to Comay’s “Relativity versus the longitudinal magnetic field of the photon”. Foundations of Physics Letters, 10(3):255–271, June 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See [489].

Castro:1997:STS

[491] Carlos Castro. String theory, scale relativity and the generalized uncertainty principle. Foundations of Physics Letters, 10(3):273–293, June 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kozlowski:1997:TAP

[492] Miroslaw Kozlowski and Janina Marciak-Kozlowska. The time arrow in a Planck gas. Foundations of Physics Letters, 10(3):295–299, June 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ahluwalia:1997:BR

[493] D. V. Ahluwalia. Book review. Foundations of Physics Letters, 10(3):301– 304, June 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1997:Ab

[494] Anonymous. Announcement. Foundations of Physics Letters, 10(3):305, June 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Santilli:1997:IGU

[495] R. M. Santilli. Isotopic grand unification with the inclusion of gravity. Foundations of Physics Letters, 10(4):307–327, August 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Goy:1997:DTD

[496] F. Goy. A derivation of three-dimensional inertial transformations. Foundations of Physics Letters, 10(4):329–346, August 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Terenzi:1997:PTB

[497] A. Terenzi and F. Cannata. Perturbation theory in Bohmian mechanics. Foundations of Physics Letters, 10(4):347–356, August 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Majernik:1997:BLS

[498] V. Majerník. Bradyon–luxon symmetry. Foundations of Physics Letters, 10(4):357–370, August 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Nistico:1997:CIQ

[499] Giuseppe Nisticò. ‘counterfactual’ interpretation of the quantum measurement process. Foundations of Physics Letters, 10(4):371–381, August 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dvoeglazov:1997:NRC

[500] Valeri V. Dvoeglazov. A note on the relativistic covariance of the B-cyclic relations. Foundations of Physics Letters, 10(4):383–391, August 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See addendum [648].

Bunchaft:1997:WLI

[501] F. Bunchaft and S. Carneiro. Weber-like interactions and energy conservation. Foundations of Physics Letters, 10(4):393–401, August 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1997:DRC

[502] M. W. Evans. The definition of B^{(3)} = B^{(0)}k: a reply to E. Comay. Foundations of Physics Letters, 10(4):403–407, August 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

SantAnna:1997:IPH

[503] Adonai S. Sant’Anna and Décio Krause. Indistinguishable particles and hidden variables. Foundations of Physics Letters, 10(5):409–426, October 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Omanovic:1997:SQS

[504] G. Omanović. Static and quasistatic solutions of the real Proca field. Foundations of Physics Letters, 10(5):427–440, October 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Croca:1997:ETR

[505] J. R. Croca, M. Ferrero, A. Garuccio, and V. L. Lepore. An experiment to test the reality of de Broglie waves. Foundations of Physics Letters, 10 (5):441–447, October 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Jimenez:1997:CTN

[506] J. L. Jimenez, I. Campos, and G. Del Valle. A canonical treatment of nonautonomous systems of type υ̇ = f(t) + c. Foundations of Physics Letters, 10(5):449–471, October 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ridderbos:1997:WFA

[507] T. M. Ridderbos. The Wheeler–Feynman absorber theory: a reinterpretation? Foundations of Physics Letters, 10(5):473–486, October 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1997:EOB

[508] M. W. Evans. Experimental observation of the b^{(3)} field: a reply to raja et alia. Foundations of Physics Letters, 10(5):487–493, October 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Yahalomi:1997:MTD

[509] E. M. Yahalomi. Measurement of the time dilation in rotational movements. Foundations of Physics Letters, 10(5):495–501, October 1997. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Burakovsky:1997:MPT

[510] L. Burakovsky, L. P. Horwitz, and W. C. Schieve. Mass-proper time uncertainty relation in a manifestly covariant relativistic statistical mechanics. Foundations of Physics Letters, 10(6):503–516, December 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Smith:1997:IEE

[511] Timothy Paul Smith. An intermediate-energy experimental test of Bell’s Inequality with recoil polarimeters. Foundations of Physics Letters, 10 (6):517–531, December 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Salesi:1997:HRQ

[512] Giovanni Salesi and Erasmo Recami. Hydrodynamical reformulation and quantum limit of the Barut–Zanghi theory. Foundations of Physics Letters, 10(6):533–546, December 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gill:1997:PTC

[513] Tepper L. Gill, W. W. Zachary, and James Lindesay. Proper-time classical electrodynamics. Foundations of Physics Letters, 10(6):547–562, December 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

vanVoorthuysen:1997:QMI

[514] E. H. du Marchie van Voorthuysen. Quantum-mechanical indirect measurements. Foundations of Physics Letters, 10(6):563–579, December 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Carazza:1997:SDM

[515] B. Carazza. On the spatial density matrix for the centre of mass of a onedimensional perfect gas. Foundations of Physics Letters, 10(6):581–590, December 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

deLaPena:1997:EPC

[516] L. de La Peña and A. M. Cetto. Estimate of Planck’s constant from an electromagnetic Mach principle. Foundations of Physics Letters, 10(6): 591–598, December 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kozlowski:1997:STI

[517] Miroslaw Kozlowski and Janina Marciak-Kozlowska. The smearing out of the thermal initial conditions created in a Planck era. Foundations of Physics Letters, 10(6):599–603, December 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1997:Ac

[518] Anonymous. Announcement. Foundations of Physics Letters, 10(6):605, December 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A:1022405504680.pdf.

Leaf:1998:FPL

[519] Boris Leaf. Frequency parts of Lorentz-gauge potentials in quantization of the electromagnetic field. Foundations of Physics Letters, 11(1):1–22, February 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cocolicchio:1998:CAC

[520] Decio Cocolicchio and Luciano Telesca. The classical analogue of CP violation. Foundations of Physics Letters, 11(1):23–39, February 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Mehra:1998:OMH

[521] Jagdish Mehra. One month in the history of the discovery of general relativity theory. Foundations of Physics Letters, 11(1):41–60, February 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Oliver:1998:CEP

[522] M. A. Oliver. Classical electrodynamics of a point particle. Foundations of Physics Letters, 11(1):61–82, February 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Datta:1998:EFE

[523] Bidyut Kumar Datta, Renuka Datta, and Venzo de Sabbata. Einstein field equations in spinor formalism: a Clifford-algebra approach. Foundations of Physics Letters, 11(1):83–93, February 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Carneiro:1998:LNH

[524] Saulo Carneiro. The large numbers hypothesis and quantum mechanics. Foundations of Physics Letters, 11(1):95–102, February 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Esposito:1998:RIF

[525] S. Esposito, M. W. Evans, and E. Recami. Radiation-induced fermion resonance. Foundations of Physics Letters, 11(1):103–107, February 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1998:Aa

[526] Anonymous. Announcement. Foundations of Physics Letters, 11(1):109, February 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Redhead:1998:UTE

[527] Michael L. G. Redhead and Fabian Wagner. Unified treatment of EPR and Bell arguments in algebraic quantum field theory. Foundations of Physics Letters, 11(2):111–125, April 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sulcs:1998:SMV

[528] Sue Sulcs and Barry C. Gilbert. Is the static magnetic vector potential screened by a thick superconducting shield. Foundations of Physics Letters, 11(2):127–139, April 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Argyris:1998:MEV

[529] John Argyris, Corneliu Ciubotariu, and Ioannis Andreadis. A metric for an Evans–Vigier field. Foundations of Physics Letters, 11(2):141–163, April 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gersten:1998:IDF

[530] Alexander Gersten. Incomplete delta functions. Foundations of Physics Letters, 11(2):165–178, April 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sjoqvist:1998:BPP

[531] Erik Sjöqvist and Henrik Carlsen. Berry’s phase in pilot–wave theory. Foundations of Physics Letters, 11(2):179–188, April 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cercignani:1998:NDP

[532] Carlo Cercignani. On a nonquantum derivation of Planck’s distribution law. Foundations of Physics Letters, 11(2):189–199, April 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Peruzzi:1998:ICR

[533] Giulio Peruzzi and Alberto Rimini. Incompatible and contradictory retrodictions in the history approach to quantum mechanics. Foundations of Physics Letters, 11(2):201–207, April 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Placek:1998:OSC

[534] Totnasz Placek. Observing a Schrödinger cat. Foundations of Physics Letters, 11(3):209–231, June 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Leifer:1998:ITE

[535] Peter Leifer. Inertia as the “threshold of elasticity” of quantum states. Foundations of Physics Letters, 11(3):233–249, June 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gersten:1998:OBC

[536] Alexander Gersten. Orthogonality and boundary conditions in quantum mechanics. Foundations of Physics Letters, 11(3):251–264, June 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Brandt:1998:QFS

[537] Howard E. Brandt. Quantum fields in the spacetime tangent bundle. Foundations of Physics Letters, 11(3):265–275, June 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Argyris:1998:ILT

[538] John Argyris, Corneliu Ciubotariu, and Ioannis Andreadis. Infinitesimal Lorentz transformations and the states of polarization of free photons. Foundations of Physics Letters, 11(3):277–285, June 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Fujita:1998:QDG

[539] Shigeji Fujita and Salvador Godoy. Quantum derivation of Ginzburg– Landau equation. new formula for penetration depth. Foundations of Physics Letters, 11(3):287–294, June 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Schay:1998:NFE

[540] Geza Schay. A new formulation of the equations of dynamics. Foundations of Physics Letters, 11(3):295–301, June 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cocolicchio:1998:OAC

[541] D. Cocolicchio, L. Telesca, and M. Viggiano. The optical analogue of CP– Violation. Foundations of Physics Letters, 11(4):303–324, August 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kamenov:1998:PS

[542] P. Kamenov and B. Slavov. The photon as a soliton. Foundations of Physics Letters, 11(4):325–342, August 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Budden:1998:GSC

[543] Tim Budden. Geometric simultaneity and the continuity of special relativity. Foundations of Physics Letters, 11(4):343–357, August 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hnizdo:1998:CME

[544] V. Hnizdo. Common misrepresentation of the Einstein–Podolsky-Rosen argument. Foundations of Physics Letters, 11(4):359–369, August 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Monk:1998:UAA

[545] N. A. M. Monk and B. J. Hiley. A unified algebraic approach to quantum theory. Foundations of Physics Letters, 11(4):371–377, August 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Janssen:1998:DRP

[546] Michel Janssen and Robert Schulmann. On the dating of a recently published Einstein manuscript: Could these be the calculations that gave Einstein “heart palpitations”? Foundations of Physics Letters, 11(4): 379–389, August 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Mehra:1998:CGE

[547] Jagdish Mehra. The calculations that gave Einstein “heart palpitations”. Foundations of Physics Letters, 11(4):391–393, August 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Feoli:1998:BMW

[548] A. Feoli and G. Scarpetta. De Broglie matter waves from the linearized Einstein field equations. Foundations of Physics Letters, 11(4):395–403, August 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Klauber:1998:NPR

[549] Robert D. Klauber. New perspectives on the relativistically rotating disk and non–time-orthogonal reference frames. Foundations of Physics Letters, 11(5):405–443, October 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Fujita:1998:PEA

[550] S. Fujita, D. Nguyen, E. S. Nam, and Salvador Godoy. Phonon-exchange attraction. optical phonons and Type II superconductivity, and d-wave Cooper pairs. Foundations of Physics Letters, 11(5):471–482, October 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Callender:1998:NEI

[551] Craig Callender and Robert Weingard. Nonlocality in the expanding infinite well. Foundations of Physics Letters, 11(5):495–498, October 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1998:AIW

[552] Anonymous. Announcement: International Workshop: “Lorentz Group, CPT, and Neutrinos”. Foundations of Physics Letters, 11(5):499, October 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1022587432038.pdf.

Anonymous:1998:AIC

[553] Anonymous. Announcement: The 11th International Congress of Logic, Methodology and Philosophy of Science. Foundations of Physics Letters, 11(5):501, October 1998. CODEN FPLEET. ISSN 0894-9875 (print),

1572-9524 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A:1022539516109.pdf.

Zimba:1998:SRC

[554] Jason Zimba. Simple realism and canonically conjugate observables in non–relativistic quantum mechanics. Foundations of Physics Letters, 11 (6):503–533, December 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Esposito:1998:CPQ

[555] Giampiero Esposito. Complex parameters in quantum mechanics. Foundations of Physics Letters, 11(6):535–547, December 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Bergia:1998:TRP

[556] Silvio Bergia and Mario Guidone. Time on a rotating platform and the one-way speed of light. Foundations of Physics Letters, 11(6):549–559, December 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Srivastava:1998:AIQ

[557] Vipin Srivastava. Arithmetic of the integer quantum Hall effect. Foundations of Physics Letters, 11(6):561–571, December 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Chubykalo:1998:FEF

[558] A. E. Chubykalo, H. A. Múnera, and R. Smirnov-Rueda. Is the free electromagnetic field a consequence of Maxwell’s equations or a postulate? Foundations of Physics Letters, 11(6):573–584, December 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Navarro:1998:TFD

[559] Miguel Navarro. Toward a finite-dimensional formulation of quantum field theory. Foundations of Physics Letters, 11(6):585–593, December 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1998:CCB

[560] M. W. Evans and Lawrence B. Crowell. Charge conjugation and the B^{(3)} field: a reply to Comay. Foundations of Physics Letters, 11(6):595–600, December 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1998:Ab

[561] Anonymous. Announcements. Foundations of Physics Letters, 11(6):601– 602, December 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Bacciagaluppi:1999:NMR

[562] Guido Bacciagaluppi. Nelsonian mechanics revisited. Foundations of Physics Letters, 12(1):1–16, February 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Tartaglia:1999:LRP

[563] A. Tartaglia. Lengths on rotating platforms. Foundations of Physics Letters, 12(1):17–28, February 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Coecke:1999:ORS

[564] Bob Coecke and Isar Stubbe. Operational resolutions and state transitions in a categorical setting. Foundations of Physics Letters, 12(1):29– 49, February 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Krause:1999:QST

[565] D. Krause, A. S. Sant’Anna, and A. G. Volkov. Quasi-set theory for bosons and fermions: Quantum distributions. Foundations of Physics Letters, 12 (1):51–66, February 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Fields:1999:SFI

[566] Dana S. Fields. The spontaneous formation of inhomogeneity in a classical one-particle “gas”: An example in silico. Foundations of Physics Letters, 12(1):67–79, February 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kozlowski:1999:KGT

[567] Miroslaw Kozlowski and Janina Marciak-Kozlowska. Klein–Gordon thermal equation for a Planck gas. Foundations of Physics Letters, 12(1): 93–98, February 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1999:BCW

[568] M. W. Evans. Brief comment on W. S. Warren et al, II: Optical NMR and the B (3) field. Foundations of Physics Letters, 12(1):99–102, February 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1999:IWL

[569] Anonymous. International Workshop: “Lorentz group, CPT, and neutrinos”, Universidad autonoma de Zacatecas, Mexico, June 23–26, 1999. Foundations of Physics Letters, 12(1):103, February 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:1999:IUH

[570] Anonymous. International union of history and philosophy of science division of logic, methodology and philosophy of science: The 11th International Congress of Logic, Methodology and Philosophy of Science. Foundations of Physics Letters, 12(1):104, February 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Oliver:1999:SDE

[571] M. A. Oliver. A simple derivation of the equation of motion for classical electrodynamics. Foundations of Physics Letters, 12(1):??, February 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ivezic:1999:CFE

[572] Tomislav Ivezić. Covariant formulation of electromagnetic 4-momentum in terms of 4-vectors E^{α} and Bα. Foundations of Physics Letters, 12(2): 105–125, April 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Jimenez:1999:MCE

[573] J. L. Jiménez and I. Campos. Models of the classical electron after a century. Foundations of Physics Letters, 12(2):127–146, April 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Coleman:1999:GFR

[574] Robert Alan Coleman and Herbert Korté. Geometry and forces in relativistic and pre-relativistic theories. Foundations of Physics Letters, 12 (2):147–163, April 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Esposito:1999:RSQ

[575] Salvatore Esposito. On the role of spin in quantum mechanics. Foundations of Physics Letters, 12(2):165–177, April 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Rizzi:1999:LGM

[576] G. Rizzi and A. Tartaglia. On local and global measurements of the speed of light on rotating platforms. Foundations of Physics Letters, 12(2):179– 186, April 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anastasovski:1999:IUT

[577] Peter K. Anastasovski, T. E. Bearden, C. Ciubotariu, W. T. Coffey, L. B. Crowell, G. J. Evans, M. W. Evans, R. Flower, S. Jeffers, A. Labounsky, D. Leporini, B. Lehnert, M. Mészáros, J. K. Moscicki, P. R. Molnár, and et al. Inconsistencies of the u(1) theory of electrodynamics: Stress energy momentum tensor. Foundations of Physics Letters, 12(2):187–192, April 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Zeh:1999:NTR

[578] H. D. Zeh. Note on the time reversal symmetry of equations of motion. Foundations of Physics Letters, 12(2):193–196, April 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Zeh:1999:WBQ

[579] H. D. Zeh. Why Bohm’s quantum theory? Foundations of Physics Letters, 12(2):197–200, April 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Harris:1999:CGE

[580] Edward G. Harris. Comments on “Gravitoelectric-Electric Coupling via Superconductivity” by Douglas G. Torr and Ning Li. Foundations of Physics Letters, 12(2):201–208, April 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See [288].

Cereceda:1999:QPC

[581] José L. Cereceda. Quantum perfect correlations and Hardy’s nonlocality theorem. Foundations of Physics Letters, 12(3):211–231, June 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See erratum [622].

Sanchez-Gomez:1999:GID

[582] José L. Sánchez-Gómez and Jesús Unturbe. Gravitationally induced decoherence in macroscopic systems. Foundations of Physics Letters, 12(3): 233–250, June 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anastasovski:1999:GTN

[583] Peter K. Anastasovski, T. E. Bearden, C. Ciubotariu, W. T. Coffey, L. B. Crowell, G. J. Evans, M. W. Evans, R. Flower, S. Jeffers, A. Labounsky, D. Leporini, B. Lehnert, M. Mészáros, J. K. Moscicki, P. R. Molnár, and et al. A general theory of non–Abelian electrodynamics. Foundations of Physics Letters, 12(3):251–265, June 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dariescu:1999:RSM

[584] Ciprian Dariescu and Marina-Aura Dariescu. Rotationally symmetric massless modes in closed Robertson–Walker universe. Foundations of Physics Letters, 12(3):267–279, June 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cardone:1999:EDP

[585] Fabio Cardone, Mauro Francaviglia, and Roberto Mignani. Energydependent phenomenological metrics and five-dimensional Einstein equations. Foundations of Physics Letters, 12(3):281–289, June 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gersten:1999:MEO

[586] A. Gersten. Maxwell’s equations as the one-photon quantum equation. Foundations of Physics Letters, 12(3):291–298, June 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Abolhasani:1999:BPA

[587] M. Abolhasani and M. Golshani. Born’s principle, action–reaction problem, and arrow of time. Foundations of Physics Letters, 12(3):299–306, June 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:1999:AEF

[588] Myron Evans. Advanced electromagnetism: Foundations, theory and applications. Foundations of Physics Letters, 12(3):307–308, June 1999. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Elitzur:1999:BHE

[589] Avshalom C. Elitzur and Shahar Dolev. Black hole evaporation entails an objective passage of time. Foundations of Physics Letters, 12(4):309– 323, August 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sen:1999:WEL

[590] R. N. Sen. Why is the Euclidean line the same as the real line? Foundations of Physics Letters, 12(4):325–345, August 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cardone:1999:EFD

[591] Fabio Cardone, Mauro Francaviglia, and Roberto Mignani. Energy as a fifth dimension. Foundations of Physics Letters, 12(4):347–371, August 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Crowell:1999:SSEa

[592] Lawrence B. Crowell and M. W. Evans. SU(2) × SU(2) electroweak theory I: the B^{3} field on the physical vacuum. Foundations of Physics Letters, 12(4):373–382, August 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dariescu:1999:MFE

[593] Marina-Aura Dariescu and Ciprian Dariescu. A Maxwell field equations class of exact solutions on matter-dominated spatially closed Friedman– Robertson–Walker universe. Foundations of Physics Letters, 12(4):383– 391, August 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Langer:1999:GLS

[594] Helmut Länger. Greechie logics with simple state space structure. Foundations of Physics Letters, 12(4):393–400, August 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Lopez-Bonilla:1999:LPP

[595] J. L. López-Bonilla, G. Ovando, and J. J. Peña. A Lanczos potential for plane gravitational waves. Foundations of Physics Letters, 12(4):401– 405, August 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Weissman:1999:EMD

[596] M. B. Weissman. Emergent measure-dependent probabilities from modified quantum dynamics without state-vector reduction. Foundations of Physics Letters, 12(5):407–426, October 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Shariati:1999:EPR

[597] A. Shariati and M. Khorrami. Equivalence principle and radiation by a uniformly accelerated charge. Foundations of Physics Letters, 12(5):427– 439, October 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kracklauer:1999:PWS

[598] A. F. Kracklauer. Pilot wave steerage: A mechanism and test. Foundations of Physics Letters, 12(5):441–453, October 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Budinich:1999:CCP

[599] P. Budinich. On conformally compacted phase space. Foundations of Physics Letters, 12(5):455–464, October 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dong:1999:ESS

[600] Shi-Hai Dong, Zhong-Qi Ma, and Giampiero Esposito. Exact solutions of the Schrödinger equation with inverse-power potential. Foundations of Physics Letters, 12(5):465–474, October 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Crowell:1999:SSEb

[601] Lawrence B. Crowell and M. W. Evans. SU(2) × SU(2) electroweak theory II: Chiral and vector fields on the physical vacuum. Foundations of Physics Letters, 12(5):475–483, October 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Carazza:1999:DWS

[602] B. Carazza. Decoherence within a simple model for the environment. Foundations of Physics Letters, 12(5):485–495, October 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

DePaoli:1999:TGS

[603] A. L. De Paoli, M. A. Estevez, M. C. Rocca, and H. Vucetich. Treatment of Gamow states, using tempered ultradistributions. Foundations of Physics Letters, 12(5):497–506, October 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See addendum [632].

Ivezic:1999:RMS

[604] Tomislav Ivezić. Relatively moving systems in “true transformations relativity”. Foundations of Physics Letters, 12(6):507–531, December 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Esposito:1999:PWM

[605] S. Esposito. Photon wave mechanics: a de Broglie–Bohm approach. Foundations of Physics Letters, 12(6):533–545, December 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ventura:1999:IAD

[606] Dan Ventura and Tony Martinez. Initializing the amplitude distribution of a quantum state. Foundations of Physics Letters, 12(6):547–559, December 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ma:1999:ESS

[607] Zhong-Qi Ma. Exact solution to the Schrödinger equation for the quantum rigid body. Foundations of Physics Letters, 12(6):561–570, December 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anandan:1999:MDM

[608] J. Anandan and Y. Aharonov. Meaning of the density matrix. Foundations of Physics Letters, 12(6):571–578, December 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anastasovski:1999:SIT

[609] Peter K. Anastasovski, T. E. Bearden, C. Ciubotariu, W. T. Coffey, L. B. Crowell, G. J. Evans, M. W. Evans, R. Flower, S. Jeffers, a Labounsky, B. Lehnert, M. Mészáros, P. R. Molnár, S. Roy, and J.-P. Vigier. Self-inconsistencies of the U(1) theory of electrodynamics: Michelson interferometry. Foundations of Physics Letters, 12(6):579–584, December 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Crowell:1999:GUP

[610] Lawrence B. Crowell. Generalized uncertainty principle for quantum fields in curved spacetime. Foundations of Physics Letters, 12(6):585–591, December 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See corrigenda [641].

Dunning-Davies:1999:TAS

[611] J. Dunning-Davies. Thermodynamics of antimatter via Santilli’s isodualities. Foundations of Physics Letters, 12(6):593–599, December 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Iannuzzi:2000:CET

[612] M. Iannuzzi, V. Merlo, A. Paciaroni, C. Petrillo, and F. Sacchetti. A coincidence experiment of two coherent beams of thermal neutrons. Foundations of Physics Letters, 13(1):1–9, February 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Schulze-Halberg:2000:FNSa

[613] Axel Schulze-Halberg. On finite normal series solutions of the Schrödinger equation with irregular singularity. Foundations of Physics Letters, 13(1): 11–27, February 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Esposito:2000:NFS

[614] Giampiero Esposito. Non-Fuchsian singularities in quantum mechanics. Foundations of Physics Letters, 13(1):29–40, February 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kaloyerou:2000:CIQ

[615] P. N. Kaloyerou. The causal interpretation of quantum fields and the vacuum. Foundations of Physics Letters, 13(1):41–54, February 2000. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Pesic:2000:IFQ

[616] Peter Pesic. Identity and the foundations of quantum theory. Foundations of Physics Letters, 13(1):55–67, February 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

SantAnna:2000:IPH

[617] Adonai S. Sant’Anna, Jeferson de Souza, and Daniel C. de Freitas. Indistinguishable particles and hidden variables: Part II. Foundations of Physics Letters, 13(1):69–77, February 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Yan:2000:HPS

[618] C. C. Yan. Hamilton’s principle and Schrödinger’s equation derived from Gauss’ Principle of Least Squares. Foundations of Physics Letters, 13 (1):79–87, February 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Frieden:2000:CCT

[619] B. Roy Frieden and B. H. Soffer. A critical comparison of three information-based approaches to physics. Foundations of Physics Letters, 13(1):89–96, February 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Peshkin:2000:CQP

[620] Murray Peshkin. Comment on “Quantum Phase Shift Caused by Spatial Confinement” by B. E. Allman et al. Foundations of Physics Letters, 13 (1):97–100, February 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See reply [621].

Allman:2000:RCQ

[621] B. E. Allman, A. Cimmino, and A. G. Klein. Reply to “Comment on Quantum Phase Shift Caused by Spatial Confinement” by M. Peshkin. Foundations of Physics Letters, 13(1):101–103, February 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See [620].

Cereceda:2000:EQP

[622] José L. Cereceda. Erratum to: Quantum Perfect Correlations and Hardy’s Nonlocality Theorem. Foundations of Physics Letters, 13(1):105, February 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A:1017228422417.pdf. See [581].

Oppenheim:2000:WDM

[623] J. Oppenheim, B. Reznik, and W. G. Unruh. When does a measurement or event occur? Foundations of Physics Letters, 13(2):107–118, April 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Marinatto:2000:WKT

[624] Luca Marinatto and Tullio Weber. Which kind of two-particle states can be teleported through a three-particle quantum channel? Foundations of Physics Letters, 13(2):119–132, April 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Feinstein:2000:OEU

[625] Alexander Feinstein and Miguel Angel Pérez Sebastián. The origin of entropy in the universe. Foundations of Physics Letters, 13(2):133–145, April 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dariescu:2000:LSP

[626] Ciprian Dariescu and Marina-Aura Dariescu. Large-scale pathology of universes with VII0 × VIII isometries. Foundations of Physics Letters, 13(2):147–165, April 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ma:2000:ESB

[627] Zhong-Qi Ma. Exact solutions to the N-body Schrödinger equation for the harmonic oscillator. Foundations of Physics Letters, 13(2):167–178, April 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anastasovski:2000:DLF

[628] P. K. Anastasovski, T. E. Bearden, C. Ciubotariu, W. T. Coffey, L. B. Crowell, G. J. Evans, M. W. Evans, R. Flower, S. Jeffers, A. Labounsky, B. Lehnert, M. Mészáros, P. R. Molnár, J. P. Vigier, and S. Roy. Derivation of the Lehnert field equations from gauge theory in vacuum: Space charge and current. Foundations of Physics Letters, 13(2):179–184, April 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gersten:2000:QEM

[629] Alexander Gersten. Quantum equations for massless particles of any spin. Foundations of Physics Letters, 13(2):185–192, April 2000. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Crowell:2000:SSE

[630] L. B. Crowell, P. K. Anastasovski, T. E. Bearden, C. Ciubotariu, W. T. Coffey, G. J. Evans, M. W. Evans, R. Flower, S. Jeffers, A. Labounsky, B. Lehnert, M. Mészáros, P. R. Molnár, J. P. Vigier, and S. Roy. SU(2) × SU(2) electroweak theory in LEP 1 data on Z particle production. Foundations of Physics Letters, 13(2):193–196, April 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Unnikrishnan:2000:PEU

[631] C. S. Unnikrishnan. Popper’s experiment, uncertainty principle, signal locality and momentum conservation. Foundations of Physics Letters, 13(2):197–200, April 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

DePaoli:2000:ATG

[632] A. L. De Paoli, M. A. Estevez, M. C. Rocca, and H. Vucetic. Addendum to “Treatment of Gamow States Using Tempered Ultradistributions”. Foundations of Physics Letters, 13(2):201–204, April 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See [603].

Sisterna:2000:TRS

[633] P. D. Sisterna. Time-reversal symmetry and time-dependent physics. Foundations of Physics Letters, 13(3):205–220, June 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Zeh:2000:PCO

[634] H. D. Zeh. The problem of conscious observation in quantum mechanical description. Foundations of Physics Letters, 13(3):221–233, June 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Floyd:2000:RTG

[635] Edward R. Floyd. Reflection time and the Goos–Hänchen effect for reflection by a semi-infinite rectangular barrier. Foundations of Physics Letters, 13(3):235–251, June 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dechoum:2000:PLD

[636] K. Dechoum, L. de la Peña, and E. Santos. The problem of light detection in the presence of zeropoint fluctuations: a physical model for the vision of weak signals. Foundations of Physics Letters, 13(3):253–264, June 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Schulze-Halberg:2000:FNSb

[637] Axel Schulze-Halberg. On finite normal series solutions of relativistic wave equations. Foundations of Physics Letters, 13(3):265–277, June 2000. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Esposito:2000:AEG

[638] Giampiero Esposito and Cosimo Stornaiolo. On the ADM equations for general relativity. Foundations of Physics Letters, 13(3):279–288, June 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anastasovski:2000:EVE

[639] P. K. Anastasovski, T. E. Bearden, C. Ciubotariu, W. T. Coffey, L. B. Crowell, G. J. Evans, M. W. Evans, R. Flower, A. Labounsky, B. Lehnert, M. Mészáros, P. R. Molnár, J. P. Vigier, and S. Roy. Effect of vacuum energy on the atomic spectra. Foundations of Physics Letters, 13(3):289– 296, June 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Salehi:2000:CIP

[640] H. Salehi, H. R. Sepangi, and F. Darabi. Conformal invariance and particle aspects in general relativity. Foundations of Physics Letters, 13(3):297– 303, June 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Crowell:2000:CGU

[641] Lawrence B. Crowell. Corrigenda: Generalized Uncertainty Principle for Quantum Fields in Curved Spacetime. Foundations of Physics Letters, 13 (3):305, June 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A:1007860323050.pdf. See [610].

Brandt:2000:FQF

[642] Howard E. Brandt. Finslerian quantum fields and microcausality. Foundations of Physics Letters, 13(4):307–328, August 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Engelman:2000:TAW

[643] R. Engelman and A. Yahalom. “Time arrow” in wave-packet evolution. Foundations of Physics Letters, 13(4):329–343, August 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Fink:2000:UQS

[644] Helmut Fink and Hajo Leschke. Is the universe a quantum system? Foundations of Physics Letters, 13(4):345–356, August 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sourbron:2000:NCD

[645] Steven Sourbron. A note on causal duality. Foundations of Physics Letters, 13(4):357–367, August 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Yan:2000:PGS

[646] C. C. Yan. A possible generalization of the Schrödinger equation based on the Gauss Principle of Least Squares. Foundations of Physics Letters, 13(4):369–378, August 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sutherland:2000:SWD

[647] Roderick I. Sutherland. A suggestive way of deriving the quantum probability rule. Foundations of Physics Letters, 13(4):379–386, August 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dvoeglazov:2000:ARC

[648] Valeri V. Dvoeglazov. Addendum to “The Relativistic Covariance of the B-Cyclic Relations”. Foundations of Physics Letters, 13(4):387–393, August 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See [500].

deBeauregard:2000:TMS

[649] O. Costa de Beauregard. Tight Mach-style connection between the Einstein mass–energy and inertia–gravity equivalences. Remarks on a recent experiment by Mikhailov. Foundations of Physics Letters, 13(4):395–399, August 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Pauri:2000:MWC

[650] M. Pauri and M. Vallisneri. Märzke–Wheeler coordinates for accelerated observers in special relativity. Foundations of Physics Letters, 13(5):401– 425, October 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cereceda:2000:QMP

[651] José L. Cereceda. Quantum mechanical probabilities and general probabilistic constraints for Einstein–Podolsky–Rosen–Bohm experiments. Foundations of Physics Letters, 13(5):427–442, October 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gregor:2000:UQH

[652] Malcolm H. Mac Gregor. A unified quantum Hall close-packed composite boson (CPCB) model. Foundations of Physics Letters, 13(5):443–460, October 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Plaga:2000:EPE

[653] R. Plaga. An extension of “Popper’s Experiment” can test interpretations of quantum mechanics. Foundations of Physics Letters, 13(5):461–476, October 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Larsson:2000:PUC

[654] Jan-Åke Larsson. A possible unification of the Copenhagen and the Bohm interpretations using local realism. Foundations of Physics Letters, 13(5): 477–486, October 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Lavenda:2000:CTU

[655] B. H. Lavenda. Comments on “Thermodynamic Uncertainty Relations” by J. Uffink and J. van Lith. Foundations of Physics Letters, 13(5):487– 492, October 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See reply [677].

Wazwaz:2000:NUA

[656] Abdul-Majid Wazwaz. A note on using Adomian decomposition method for solving boundary value problems. Foundations of Physics Letters, 13 (5):493–498, October 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Aravind:2000:HRC

[657] P. K. Aravind. How Reye’s configuration helps in proving the Bell– Kochen–Specker Theorem: a curious geometrical tale. Foundations of Physics Letters, 13(6):499–519, December 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sulcs:2000:RLR

[658] Sue Sulcs, Graham Oppy, and Barry C. Gilbert. The rejection of local realism is premature. Foundations of Physics Letters, 13(6):521–541, December 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Mitra:2000:NOT

[659] Abhas Mitra. Non-occurrence of trapped surfaces and black holes in spherical gravitational collapse. Foundations of Physics Letters, 13(6):543–579, December 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Brandt:2000:CDM

[660] Howard E. Brandt. Causal domain of Minkowski–spacetime tangent bundle. Foundations of Physics Letters, 13(6):581–588, December 2000. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Bedford:2000:MET

[661] Donald Bedford and Peter Krumm. A mechanism for energy transfer in mechanical systems. Foundations of Physics Letters, 13(6):589–594, December 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Nikolic:2000:RAL

[662] Hrvoje Nikolić. The role of acceleration and locality in the Twin Paradox. Foundations of Physics Letters, 13(6):595–601, December 2000. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Jakel:2001:SCE

[663] Christian D. Jäkel. Some comments on entanglement and local thermofield theory. Foundations of Physics Letters, 14(1):1–16, February 2001. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Bouda:2001:PCT

[664] A. Bouda. Probability current and trajectory representation. Foundations of Physics Letters, 14(1):17–35, February 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

DeLeo:2001:QLG

[665] Stefano De Leo. Quaternionic Lorentz group and Dirac equation. Foundations of Physics Letters, 14(1):37–50, February 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Bartocci:2001:LPB

[666] Umberto Bartocci, Fabio Cardone, and Roberto Mignani. Looking for a possible breakdown of local Lorentz invariance for electromagnetic phenomena: Theory and first experimental results. Foundations of Physics Letters, 14(1):51–64, February 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wang:2001:SWQ

[667] Lu-Ya Wang, Xiao-Yan Gu, Bin Duan, and Zhong-Qi Ma. Symmetrized wavefunctions in a quantum N-body system. Foundations of Physics Letters, 14(1):65–76, February 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gsponer:2001:CFG

[668] André Gsponer and Jean-Pierre Hurni. Comment on formulating and generalizing Dirac’s, Proca’s, and Maxwell’s equations with biquaternions or Clifford numbers. Foundations of Physics Letters, 14(1):77–85, February 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anastasovski:2001:EMEa

[669] P. K. Anastasovski, T. E. Bearden, C. Ciubotariu, W. T. Coffey, L. B. Crowell, G. J. Evans, M. W. Evans, R. Flower, A. Labounsky, B. Lehnert, M. Mészáros, P. R. Molnár, J. K. Moscicki, S. Roy, and J. P. Vigier. Explanation of the motionless electromagnetic generator with O(3) electrodynamics. Foundations of Physics Letters, 14(1):87–94, February 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

McCall:2001:IIQ

[670] Storrs McCall. The Ithaca interpretation of quantum mechanics, and objective probabilities. Foundations of Physics Letters, 14(1):95–101, February 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cunha:2001:QPD

[671] M. O. Terra Cunha, V. I. Man’ko, and M. O. Scully. Quasiprobability and probability distributions for Spin-1/ 2 states. Foundations of Physics Letters, 14(2):103–117, April 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Laudisa:2001:EAR

[672] Federico Laudisa. The EPR argument in a relational interpretation of quantum mechanics. Foundations of Physics Letters, 14(2):119–132, April 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Coecke:2001:QDC

[673] Bob Coecke, David J. Moore, and Isar Stubbe. Quantaloids describing causation and propagation of physical properties. Foundations of Physics Letters, 14(2):133–145, April 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Zafiris:2001:QESa

[674] Elias Zafiris. On quantum event structures part i: The categorical scheme. Foundations of Physics Letters, 14(2):147–166, April 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Zafiris:2001:QESb

[675] Elias Zafiris. On quantum event structures. Part II: Interpretational aspects. Foundations of Physics Letters, 14(2):167–177, April 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Clarke:2001:HIS

[676] C. J. S. Clarke. The histories interpretation: Stability instead of consistency? Foundations of Physics Letters, 14(2):179–186, April 2001. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Uffink:2001:TUR

[677] J. Uffink and J. van Lith. Thermodynamic uncertainty relations again: a reply to Lavenda. Foundations of Physics Letters, 14(2):187–193, April 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See [655].

Gudder:2001:BRE

[678] Stanley P. Gudder. Book review: Einstein and Religion. By Max Jammer. Princeton University Press, Princeton, New Jersey, 1999, vii + 279 pp., $22.00 (hardcover). ISBN 0-691-00699-7. Foundations of Physics Letters, 14(2):195–198, April 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Caves:2001:EFA

[679] Carlton M. Caves, Christopher A. Fuchs, and Pranaw Rungta. Entanglement of formation of an arbitrary state of two rebits. Foundations of Physics Letters, 14(3):199–212, June 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Giunti:2001:QMN

[680] Carlo Giunti and Chung W. Kim. Quantum mechanics of neutrino oscillations. Foundations of Physics Letters, 14(3):213–229, June 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hiesmayr:2001:GBI

[681] Beatrix C. Hiesmayr. A generalized Bell inequality and decoherence for the K^{0}K^{o}. Foundations of Physics Letters, 14(3):231–245, June 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

SantAnna:2001:AMM

[682] Adonai S. Sant’Anna and Clóvis A. S. Maia. Axioma for Mach’s mechanics. Foundations of Physics Letters, 14(3):247–262, June 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Caser:2001:EDG

[683] S. Caser. Electrodynamics in Dirac’s gauge: a geometrical equivalence. Foundations of Physics Letters, 14(3):263–274, June 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Short:2001:PEC

[684] A. J. Short. Popper’s experiment and conditional uncertainty relations. Foundations of Physics Letters, 14(3):275–284, June 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Stulpe:2001:QSP

[685] Werner Stulpe and Maciej Swat. Quantum states as probability measures. Foundations of Physics Letters, 14(3):285–293, June 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Houili:2001:GG

[686] H. Houili and T. H. Zerguini. Geons and (4 + 1) gravity. Foundations of Physics Letters, 14(3):295–300, June 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Rubin:2001:LEI

[687] Mark A. Rubin. Locality in the Everett interpretation of Heisenbergpicture quantum mechanics. Foundations of Physics Letters, 14(4):301– 322, August 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Yamaleev:2001:PTR

[688] Robert M. Yamaleev and Ana-Leticia Fernandez Osorio. Proper-time relativistic dynamics on hyperboloid. Foundations of Physics Letters, 14 (4):323–339, August 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Duane:2001:VBI

[689] Gregory S. Duane. Violations of Bell’s inequality in synchronized hyperchaos. Foundations of Physics Letters, 14(4):341–353, August 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Iannuzzi:2001:IEB

[690] M. Iannuzzi and V. Merlo. Interpretation of the experiments on Bell’s inequality. Foundations of Physics Letters, 14(4):355–365, August 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cardone:2001:EDD

[691] Fabio Cardone and Roberto Mignani. Energy-dependent description of gravitation supported by an electromagnetic test of local Lorentz invariance. Foundations of Physics Letters, 14(4):367–376, August 2001. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Mould:2001:CES

[692] Richard Mould. Consciousness and endogenous state reduction: Two experiments. Foundations of Physics Letters, 14(4):377–386, August 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anastasovski:2001:EMEb

[693] P. K. Anastasovski, T. E. Bearden, C. Ciubotariu, W. T. Coffey, L. B. Crowell, G. J. Evans, M. W. Evans, R. Flower, A. Labounsky, B. Lehnert, M. Mészáros, P. R. Molnár, J. K. Moscicki, S. Roy, and J. P. Vigier. Explanation of the motionless electromagnetic generator by Sachs’s theory of electrodynamics. Foundations of Physics Letters, 14(4):387–393, August 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Petroni:2001:ROS

[694] Nicola Cufaro Petroni and Jean-Pierre Vigier. Remarks on observed superluminal light propagation. Foundations of Physics Letters, 14(4):395– 400, August 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cereceda:2001:IAH

[695] José L. Cereceda. Identification of all Hardy-type correlations for two photons or particles with spin 1/2. Foundations of Physics Letters, 14 (5):401–424, October 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dvurecenskij:2001:CSP

[696] Anatolij Dvurecenskij and Thomas Vetterlein. Congruences and states on pseudoeffect algebras. Foundations of Physics Letters, 14(5):425–446, October 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Zareski:2001:EIE

[697] David Zareski. The elastic interpretation of electrodynamics. Foundations of Physics Letters, 14(5):447–469, October 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Biro:2001:CQC

[698] T. S. Biró, S. G. Matinyan, and B. Müller. Chaotic quantization of classical gauge fields. Foundations of Physics Letters, 14(5):471–485, October 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Rembielinski:2001:TPF

[699] Jakub Rembieliński, Kordian Andrzej Smoliński, and Grzegorz Duniec. Thermodynamics and preferred frame. Foundations of Physics Letters, 14 (5):487–500, October 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ischi:2001:GQM

[700] Boris Ischi. Generalized quantum mechanics for separated systems. Foundations of Physics Letters, 14(6):501–518, December 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ludwin:2001:RMT

[701] D. Ludwin and Y. Ben-Aryeh. The role of momentum transfer in Welcher– Weg experiments. Foundations of Physics Letters, 14(6):519–528, December 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kamenov:2001:PSS

[702] P. S. Kamenov and B. D. Slavov. Photon-solitons and some experimental possibilities. Foundations of Physics Letters, 14(6):529–540, December 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Lanciani:2001:RMC

[703] Paolo Lanciani. The rest mass of (charged) leptons in 6-D spacetime. Foundations of Physics Letters, 14(6):541–551, December 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

daCosta:2001:MRT

[704] Newton C. A. da Costa and Adonai S. Sant’Anna. The mathematical role of time and space–time in classical physics. Foundations of Physics Letters, 14(6):553–563, December 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wang:2001:ESA

[705] Lu-Ya Wang and Zhong-Qi Ma. Exact solutions with arbitrary permutation symmetry in a three-body harmonic oscillator. Foundations of Physics Letters, 14(6):565–572, December 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Abbassi:2001:NC

[706] Amir H. Abbassi and Amir M. Abbassi. The Newtonian correspondence. Foundations of Physics Letters, 14(6):573–580, December 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Agnese:2001:SMQ

[707] A. G. Agnese and M. La Camera. Schwarzschild metrics and quasi– universes. Foundations of Physics Letters, 14(6):581–587, December 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anastasovski:2001:DBF

[708] P. K. Anastasovski, T. E. Bearden, C. Ciubotariu, W. T. Coffey, L. B. Crowell, G. J. Evans, M. W. Evans, R. Flower, A. Labounsky, B. Lehnert, M. Mészáros, P. R. Molnár, J. K. Moscicki, S. Roy, and J. P. Vigier. Derivation of the b^{(3)} field and concomitant vacuum energy density from the Sachs theory of electrodynamics. Foundations of Physics Letters, 14 (6):589–593, December 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anastasovski:2001:DST

[709] P. K. Anastasovski, T. E. Bearden, C. Ciubotariu, W. T. Coffey, L. B. Crowell, G. J. Evans, M. W. Evans, R. Flower, A. Labounsky, B. Lehnert, M. Mészáros, P. R. Molnár, S. Roy, and J. P. Vigier. Development of the Sachs theory of electrodynamics. Foundations of Physics Letters, 14(6): 595–600, December 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anastasovski:2001:AGE

[710] P. K. Anastasovski, T. E. Bearden, C. Ciubotariu, W. T. Coffey, L. B. Crowell, G. J. Evans, M. W. Evans, R. Flower, A. Labounsky, B. Lehnert, M. Mészáros, P. R. Molnár, S. Roy, and J. P. Vigier. Anti-gravity effects in the Sachs theory of electrodynamics. Foundations of Physics Letters, 14 (6):601–605, December 2001. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ivanov:2002:IME

[711] V. V. Ivanov, I. V. Boichenko, and A. V. Luzanov. Influence of many-electron effects on the probabilities of one- and two-photon π–π^{∗}- transitions in conjugate organic molecules. Foundations of Physics Letters, 15(1):457–463, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Panin:2002:CMO

[712] A. I. Panin and O. V. Sizova. Construction of molecular orbitals localized on polyatomic fragments. Foundations of Physics Letters, 15(1):464–470, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sizova:2002:IET

[713] O. V. Sizova, V. I. Baranovskii, A. I. Panin, and N. V. Ivanova. Intramolecular electron transfer in mixed-valence complexes [(NH_{3})_{5}Ru-L-Ru(NH_{3})_{5}]^{5+} (L = N_{2}, pyz, pym, 4,4’-bipy, bpa). Foundations of Physics Letters, 15(1):471–479, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Glushkov:2002:CDV

[714] A. V. Glushkov, V. A. Efimov, E. D. Gopchenko, S. V. Ambrosov, V. I. Polishchuk, and I. M. Shpinareva. Calculation of diatomic van der Waals systems inert gas atom–halogen type inert gas ion in the ground state. Foundations of Physics Letters, 15(1):480–483, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ermakov:2002:ORM

[715] A. I. Ermakov and A. E. Merkulov. Orbital relaxation and molecular properties of silicon compounds. Foundations of Physics Letters, 15(1): 484–488, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Frolov:2002:ICV

[716] Yu. L. Frolov, A. V. Knizhnik, and A. G. Malkina. Ab initio calculations and vibrational spectra of CH_{3}SC=CH and CH_{3}SC=CSCH_{3}. Foundations of Physics Letters, 15(1):489–495, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Frolov:2002:CSD

[717] Yu. L. Frolov and A. V. Knizhnik. Conformational structure of divinylacetylene. Foundations of Physics Letters, 15(1):496–501, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Bocharov:2002:QCC

[718] V. N. Bocharov, S. F. Bureiko, A. Koll, and M. Rospenk. Quantum chemical calculations of the molecular structure of diphenylguanidine and experimental studies of its tautomerism in solution. Foundations of Physics Letters, 15(1):502–507, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Morozov:2002:FCV

[719] V. P. Morozov, V. V. Belokopytov, G. D. Zegzhda, and V. N. Moiseenko. Force constants and vibration frequencies and forms of the coordination center of cysteine-containing chromium (III) trischelates. Foundations of Physics Letters, 15(1):508–513, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ezhov:2002:MST

[720] Yu. S. Ezhov, S. A. Komarov, and V. G. Sevast’yanov. Molecular structure of tris(acetylacetonato)scandium Sc(C_{5}H_{7}O_{2})_{3}. Foundations of Physics Letters, 15(1):514–519, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kalninsh:2002:SEA

[721] K. K. Kalninsh, A. D. Kutsenko, and A. S. Archegova. Structure, electronaccepting properties, and reactivity of nitrostilbenes. Foundations of Physics Letters, 15(1):520–531, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Martynets:2002:SLS

[722] V. G. Martynets, P. P. Bezverkhii, and é. V. Matizen. Structure of liquid solutions near the critical point. Foundations of Physics Letters, 15(1): 532–542, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Borisov:2002:CSP

[723] S. V. Borisov, N. V. Podberezskaya, N. V. Pervukhina, and S. A. Magarill. Crystal structure of polymorphous modifications of Tm_{2}S_{3} and related compounds. Foundations of Physics Letters, 15(1):543–548, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Baidina:2002:CSP

[724] I. A. Baidina, G. I. Zharkova, and S. A. Gromilov. Crystal structure of platinum(II) chloro(1,1,1-trifluoro-4-imino-2-pentanononate(hydrate)) [PtC1(ktf)H_{2}O]. Foundations of Physics Letters, 15(1):549–553, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ivanchenko:2002:SCS

[725] A. V. Ivanchenko, S. A. Gromilov, S. M. Zemskova, I. A. Baidina, and L. A. Glinskaya. Synthesis and crystal structure of bis(di-n-butyldithiocarbamato)(1,10-phenanthroline)cadmium(II). Foundations of Physics Letters, 15(1):554–558, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Glinskaya:2002:CMS

[726] L. A. Glinskaya, R. F. Klevtsova, E. I. Berus, S. M. Zemskova, and S. V. Larionov. Crystal and molecular structure of zinc(II) di-n-propyldithiocarbamate complexes with 2,2-bipyridyl and 1,10phenanthroline. Foundations of Physics Letters, 15(1):559–566, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Khlebnikov:2002:AOD

[727] A. I. Khlebnikov. An algorithm for optimal design of chemical structures from molecular fragments. Foundations of Physics Letters, 15(1):567–574, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Orlov:2002:PPI

[728] I. P. Orlov, K. A. Popov, and V. S. Urusov. Program for predicting interatomic distances in crystals by the bond valence method. Foundations of Physics Letters, 15(1):575–579, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Babailov:2002:NMM

[729] S. P. Babailov and Yu. G. Krieger. NMR methods for molecular structure studies of paramagnetic lanthanide complexes in solutions. Applications to crown ether complexes. Foundations of Physics Letters, 15(1):580–593, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Pedash:2002:EDI

[730] Yu. F. Pedash and A. Yu. Semenov. Effect of dispersion interaction on the polarizability and hyperpolarizability of a molecular dimer. Foundations of Physics Letters, 15(1):594–596, February 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Soklakov:2002:ORF

[731] Andrei N. Soklakov. Occam’s razor as a formal basis for a physical theory. Foundations of Physics Letters, 15(2):107–135, April 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ahmedov:2002:ECG

[732] B. J. Ahmedov and M. J. Ermamatov. Electrical conductivity in general relativity. Foundations of Physics Letters, 15(2):137–151, April 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Minguzzi:2002:CS

[733] E. Minguzzi. On the conventionality of simultaneity. Foundations of Physics Letters, 15(2):153–169, April 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dong:2002:DED

[734] Shi-Hai Dong and Zhong-Qi Ma. The (2 + 1) Dirac equation with a delta potential. Foundations of Physics Letters, 15(2):171–178, April 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anastasovski:2002:DEI

[735] P. K. Anastasovski, T. E. Bearden, C. Ciubotariu, W. T. Coffey, L. B. Crowell, G. J. Evans, M. W. Evans, R. Flower, A. Labounsky, B. Lehnert, M. Mészáros, P. R. Molnár, S. Roy, and J. P. Vigier. Derivation of O(3) electrodynamics from the irreducible representations of the Einstein group. Foundations of Physics Letters, 15(2):179–187, April 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Chen:2002:IEM

[736] Jing-Ling Chen and Abraham A. Ungar. Introducing the Einstein metric to quantum computation and quantum information geometry. Foundations of Physics Letters, 15(2):189–197, April 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Asanov:2002:FEL

[737] G. S. Asanov. Finslerian extension of Lorentz transformations and firstorder censorship theorem. Foundations of Physics Letters, 15(2):199–207, April 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Salgado:2002:SCT

[738] D. Salgado and J. L. Sánchez-Gómez. Some comments on three suggested postulates for quantum theory. Foundations of Physics Letters, 15(3):209– 227, June 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Zak:2002:QEN

[739] Michail Zak. Quantum evolution as a nonlinear Markov process. Foundations of Physics Letters, 15(3):229–243, June 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:2002:MGF

[740] M. W. Evans, T. E. Bearden, and A. Labounsky. The most general form of the vector potential in electrodynamics. Foundations of Physics Letters, 15(3):245–261, June 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Singleton:2002:SLM

[741] D. Singleton and A. Yoshida. A Schwarzschild-like model for baryons. Foundations of Physics Letters, 15(3):263–275, June 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dariescu:2002:SSR

[742] M. A. Dariescu and C. Dariescu. SU(N) supersymmetry on S^{3} × r spacetime. Foundations of Physics Letters, 15(3):277–286, June 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Brandt:2002:DWD

[743] Howard E. Brandt. Deconstructing Wigner’s density matrix concerning the mind-body question. Foundations of Physics Letters, 15(3):287–292, June 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Rigolin:2002:URE

[744] G. Rigolin. Uncertainty relations for entangled states. Foundations of Physics Letters, 15(3):293–298, June 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Koltko:2002:GTC

[745] Kurt Koltko. A gauge theory of CPT transformations to first order. Foundations of Physics Letters, 15(3):299–304, June 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ahmedov:2002:RAH

[746] B. J. Ahmedov and M. J. Ermamatov. Rotational analog of the Hall effect: Coriolis contribution to electric current. Foundations of Physics Letters, 15(3):305–309, June 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Matone:2002:EPQ

[747] Marco Matone. Equivalence postulate and quantum origin of gravitation. Foundations of Physics Letters, 15(4):311–328, August 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

John:2002:MBB

[748] Moncy V. John. Modified de Broglie–Bohm approach to quantum mechanics. Foundations of Physics Letters, 15(4):329–343, August 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Criado:2002:RKH

[749] C. Criado and N. Alamo. Relativistic kinematic honeycombs. Foundations of Physics Letters, 15(4):345–358, August 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hnilo:2002:LRM

[750] Alejandro Hnilo, Alejandro Peuriot, and Guillermo Santiago. Local realistic models tested by the EPRB experiment with random variable analyzers. Foundations of Physics Letters, 15(4):359–371, August 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Herrera:2002:TPT

[751] Luis Herrera and Alicia Di Prisco. The Thomas precession and the transformation to rotating frames. Foundations of Physics Letters, 15(4):373– 383, August 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dong:2002:AMA

[752] Shi-Hai Dong. The ansatz method for analyzing Schrödinger’s equation with three anharmonic potentials in D dimensions. Foundations of Physics Letters, 15(4):385–395, August 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Aravind:2002:BTI

[753] P. K. Aravind. Bell’s theorem without inequalities and only two distant observers. Foundations of Physics Letters, 15(4):397–405, August 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cardone:2002:BFC

[754] Fabio Cardone, Roberto Mignani, and Eliano Pessa. Bound on the Qfactor of a G-constant measuring pendulum device from gravitational stochastic background. Foundations of Physics Letters, 15(4):407–413, August 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anandan:2002:CSQ

[755] Jeeva Anandan. Causality, symmetries and quantum mechanics. Foundations of Physics Letters, 15(5):415–438, October 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Mitra:2002:FSS

[756] Abbas Mitra. On the final state of spherical gravitational collapse. Foundations of Physics Letters, 15(5):439–471, October 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sica:2002:CNB

[757] Louis Sica. Correlations for a new Bell’s inequality experiment. Foundations of Physics Letters, 15(5):473–486, October 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gogberashvili:2002:CFS

[758] Merab Gogberashvili. Coriolis force and Sagnac effect. Foundations of Physics Letters, 15(5):487–493, October 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Clarke:2002:ESI

[759] C. J. S. Clarke. Entanglement and statistical independence for mixed quantum states. Foundations of Physics Letters, 15(5):495–500, October 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sidharth:2002:BNE

[760] B. G. Sidharth. A brief note on “extension, spin and non-commutativity”. Foundations of Physics Letters, 15(5):501–506, October 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Colombo:2002:SCL

[761] Fabrizio Colombo, Irene Sabadini, Frank Sommen, and Daniele C. Struppa. Syzygies and conservation laws. Foundations of Physics Letters, 15(6):507–522, December 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Schlesinger:2002:UST

[762] Karl-Georg Schlesinger. On the universality of string theory. Foundations of Physics Letters, 15(6):523–536, December 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Nelson:2002:CCR

[763] R. D. Nelson, D. I. Radin, R. Shoup, and P. A. Bancel. Correlations of continuous random data with major world events. Foundations of Physics Letters, 15(6):537–550, December 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ianconescu:2002:SFC

[764] Reuven Ianconescu and L. P. Horwitz. Self-force of a charge in a real current. Foundations of Physics Letters, 15(6):551–559, December 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anastasovski:2002:ABE

[765] P. K. Anastasovski, T. E. Bearden, C. Ciubotariu, W. T. Coffey, L. B. Crowell, G. J. Evans, M. W. Evans, R. Flower, S. Jeffers, A. Labounsky, B. Lehnert, M. Mészáros, P. R. Molnar, S. Roy, and J. P. Vigier. Aharonov-bohm effect as the basis of electromagnetic energy inherent in the vacuum. Foundations of Physics Letters, 15(6):561–568, December 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Wang:2002:ESD

[766] Lu-Ya Wang, Xiao-Yan Gu, Zhong-Qi Ma, and Shi-Hai Dong. Exact solutions to D-dimensional Schrödinger equation with a pseudoharmonic oscillator. Foundations of Physics Letters, 15(6):569–576, December 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sidharth:2002:PSP

[767] B. G. Sidharth. Planck-scale phenomena. Foundations of Physics Letters, 15(6):577–583, December 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Schulze-Halberg:2002:CFS

[768] Axel Schulze-Halberg. Closed-form solutions of the Schrödinger equation for a particle on the torus. Foundations of Physics Letters, 15(6):585–589, December 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gersten:2002:LFS

[769] Alexander Gersten and Abed Shaqir. Lagrangian formulation of the Schrödinger equation. Foundations of Physics Letters, 15(6):591–594, December 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ungar:2003:APE

[770] Abraham A. Ungar. On the appeal to a pre-established harmony between pure mathematics and relativity physics. Foundations of Physics Letters, 16(1):1–23, February 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Woodward:2003:BPF

[771] James F. Woodward. “Breakthrough” propulsion and the foundations of physics. Foundations of Physics Letters, 16(1):25–40, February 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Devereux:2003:MSE

[772] Michael Devereux. A modified Szilard engine: Measurement, information, and Maxwell’s Demon. Foundations of Physics Letters, 16(1):41–55, February 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dzhunushaliev:2003:NOQ

[773] V. Dzhunushaliev. Nonperturbative operator quantization of strongly nonlinear fields. Foundations of Physics Letters, 16(1):57–70, February 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Selleri:2003:FE

[774] F. Selleri. On the Fizeau experiment. Foundations of Physics Letters, 16 (1):71–82, February 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ibison:2003:EZP

[775] M. Ibison. Electrodynamics in the zero-point field: on the equilibrium spectral energy distribution and the origin of inertial mass. Foundations of Physics Letters, 16(1):83–90, February 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sidharth:2003:CS

[776] B. G. Sidharth. Complexified spacetime. Foundations of Physics Letters, 16(1):91–97, February 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kaniadakis:2003:NQM

[777] G. Kaniadakis. Nonrelativistic quantum mechanics with spin in the framework of a classical subquantum kinetics. Foundations of Physics Letters, 16(2):99–110, April 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Urbantke:2003:LTR

[778] H. K. Urbantke. Lorentz transformations from reflections: Some applications. Foundations of Physics Letters, 16(2):111–117, April 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Overduin:2003:VEE

[779] James Overduin and Hans-Jörg Fahr. Vacuum energy and the economical universe. Foundations of Physics Letters, 16(2):119–125, April 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Mehrafarin:2003:INQ

[780] Mohammad Mehrafarin. The informational nature of quantum mechanics: a novel look at the interference experiment. Foundations of Physics Letters, 16(2):127–133, April 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Modanese:2003:IMV

[781] Giovanni Modanese. Inertial mass and vacuum fluctuations in quantum field theory. Foundations of Physics Letters, 16(2):135–141, April 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Leiter:2003:DPE

[782] Darryl Leiter and Stanley Robertson. Does the principle of equivalence prohibit trapped surfaces from forming in the general relativistic collapse process? Foundations of Physics Letters, 16(2):143–161, April 2003. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cardone:2003:BAD

[783] Fabio Cardone, Alessio Marrani, and Roberto Mignani. Boosts in an arbitrary direction and maximal causal velocities in a deformed Minkowski space. Foundations of Physics Letters, 16(2):163–181, April 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Vigoureux:2003:ECL

[784] J.-M. Vigoureux, P. Vigoureux, and B. Vigoureux. The Einstein constant c in light of Mach’s principle. cosmological applications. Foundations of Physics Letters, 16(2):183–193, April 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anastasovski:2003:IFE

[785] P. K. Anastasovski, T. E. Bearden, C. Ciubotariu, W. T. Coffey, L. B. Crowell, G. J. Evans, M. W. Evans, R. Flower, A. Labounsky, B. Lehnert, M. Mészáros, P. R. Molnár, J. K. Moscicki, S. Roy, and J. P. Vigier. Inverse Faraday effect from the first principles of general relativity. Foundations of Physics Letters, 16(2):195–198, April 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kirkpatrick:2003:QBC

[786] K. A. Kirkpatrick. “Quantal” behavior in classical probability. Foundations of Physics Letters, 16(3):199–224, June 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ianconescu:2003:EMC

[787] Reuven Ianconescu and L. P. Horwitz. Energy mechanism of charges analyzed in real current environment. Foundations of Physics Letters, 16(3):225–244, June 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Brun:2003:CCT

[788] Todd A. Brun. Computers with closed timelike curves can solve hard problems efficiently. Foundations of Physics Letters, 16(3):245–253, June 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Carvalho:2003:NVD

[789] Marcelo Carvalho and Alexandre L. Oliveira. A new version of the Dirac’s Æther and its cosmological applications. Foundations of Physics Letters, 16(3):255–263, June 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dzhunushaliev:2003:NSS

[790] Vladimir Dzhunushaliev. Nonperturbative spontaneous symmetry breaking. Foundations of Physics Letters, 16(3):265–273, June 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anastasovski:2003:DEE

[791] P. K. Anastasovski, T. E. Bearden, C. Ciubotariu, W. T. Coffey, L. B. Crowell, G. J. Evans, M. W. Evans, R. Flower, A. Labounsky, B. Lehnert, M. Mészáros, P. R. Molnár, J. K. Moscicki, S. Roy, and J. P. Vigier. Derivation of O(3) electrodynamics from the Einstein–Sachs theory of general relativity. Foundations of Physics Letters, 16(3):275–280, June 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Crowell:2003:QOD

[792] Lawrence B. Crowell. Quantum optical detection of the holographic principle. Foundations of Physics Letters, 16(3):281–286, June 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hacyan:2003:RQC

[793] Shahen Hacyan. Relativistic quantum correlations between Spin-1/ 2 particles. Foundations of Physics Letters, 16(3):287–291, June 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Afriat:2003:ARP

[794] Alexander Afriat. Altering the remote past. Foundations of Physics Letters, 16(3):293–301, June 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Colijn:2003:SDB

[795] C. Colijn and E. R. Vrscay. Spin-dependent Bohm trajectories for Pauli and Dirac eigenstates of hydrogen. Foundations of Physics Letters, 16 (4):303–323, August 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Matsutani:2003:WPC

[796] Shigeki Matsutani and Yoshihiro Ônishi. Wave-particle complementarity and reciprocity of Gauss sums on Talbot effects. Foundations of Physics Letters, 16(4):325–341, August 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Lynds:2003:TCQ

[797] Peter Lynds. Time and classical and quantum mechanics: Indeterminacy versus discontinuity. Foundations of Physics Letters, 16(4):343–355, August 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dong:2003:QSS

[798] Shi-Hai Dong and Guo-Hua San. Quantum spectrum of some anharmonic central potentials: Wave functions ansatz. Foundations of Physics Letters, 16(4):357–367, August 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:2003:GCF

[799] Myron W. Evans. A generally covariant field equation for gravitation and electromagnetism. Foundations of Physics Letters, 16(4):369–377, August 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Christiaens:2003:NSE

[800] Wim Christiaens. Non-spatiality and EPR-experiments according to the creation–discovery view. Foundations of Physics Letters, 16(4):379–387, August 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Neeman:2003:SMO

[801] Yuval Ne’eman. The sign and micro-origin of complexity as entropy (conjectured resolution of a paradox). Foundations of Physics Letters, 16(4): 389–394, August 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Crowell:2003:RGT

[802] Lawrence B. Crowell. Ring gyro test of the holographic principle. Foundations of Physics Letters, 16(4):395–401, August 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Encinosa:2003:FSR

[803] Mario Encinosa and Babak Etemadi. Fourier series representation of lowlying eigenfunctions for a particle on the torus. Foundations of Physics Letters, 16(4):403–409, August 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Pereverzev:2003:IQM

[804] Andrey Pereverzev. Interpretation of the quantum measurement: Example of a linearly coupled harmonic oscillator. Foundations of Physics Letters, 16(5):411–428, October 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dreisigmeyer:2003:DGP

[805] David W. Dreisigmeyer, Richard Clawson, R. Eykholt, and Peter M. Young. Dynamic and geometric phase formulas in the Hestenes–Dirac theory. Foundations of Physics Letters, 16(5):429–445, October 2003. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Klauber:2003:DGC

[806] Robert D. Klauber. Derivation of the general case Sagnac result using non–time-orthogonal analysis. Foundations of Physics Letters, 16(5):447– 463, October 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Clements:2003:BFC

[807] David J. Clements and Myron Evans. The b^{(3)} field from curved spacetime embedded in the irreducible representation of the Einstein group. Foundations of Physics Letters, 16(5):465–479, October 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Puccini:2003:EUP

[808] G. D. Puccini. Electrodynamics under a possible alternative to the Lorentz transformation. Foundations of Physics Letters, 16(5):481–494, October 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Barashenkov:2003:RGS

[809] V. S. Barashenkov and E. Kapuścik. Relativity groups for spacetimes with invariant velocities. Foundations of Physics Letters, 16(5):495–504, October 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Genovese:2003:DMQ

[810] M. Genovese. On DNA molecules as quantum measurement devices. Foundations of Physics Letters, 16(5):505–511, October 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:2003:GCW

[811] Myron W. Evans. A generally covariant wave equation for grand unified field theory. Foundations of Physics Letters, 16(6):513–547, December 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ghosal:2003:CST

[812] S. K. Ghosal, Biplab Raychaudhuri, Anjan Kumar Chowdhury, and Minakshi Sarker. Conventionality of simultaneity and the Tippe top paradox in relativity. Foundations of Physics Letters, 16(6):549–563, December 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

SantAnna:2003:GHM

[813] Adonai S. Sant’Anna and Christiano Garcia. Gravitation in Hertz mechanics. Foundations of Physics Letters, 16(6):565–578, December 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Alhaidari:2003:DMM

[814] A. D. Alhaidari. Deformation of the J-matrix method of scattering. Foundations of Physics Letters, 16(6):579–591, December 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Minguzzi:2003:UOW

[815] Ettore Minguzzi and Alan Macdonald. Universal one-way light speed from a universal light speed over closed paths. Foundations of Physics Letters, 16(6):593–604, December 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Garola:2003:EQM

[816] Claudio Garola. Embedding quantum mechanics into an objective framework. Foundations of Physics Letters, 16(6):605–612, December 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Boyer:2003:CCZ

[817] Timothy H. Boyer. Comments on Cole and Zou’s calculation of the hydrogen ground state in classical physics. Foundations of Physics Letters, 16 (6):613–617, December 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Milonni:2003:CCZ

[818] Peter W. Milonni. Comment on Cole and Zou’s classical computations of the hydrogen ground state. Foundations of Physics Letters, 16(6):619–621, December 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Lopez-Marino:2004:AAM

[819] M. A. López-Mariño and J. L. Jiménez. Analysis of the Abraham– Minkowski controversy by means of two simple examples. Foundations of Physics Letters, 17(1):1–23, February 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:2004:EGU

[820] Myron W. Evans. The equations of grand unified field theory in terms of the Maurer–Cartan structure relations of differential geometry. Foundations of Physics Letters, 17(1):25–47, February 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Pfister:2004:NFL

[821] Herbert Pfister. Newton’s first law revisited. Foundations of Physics Letters, 17(1):49–64, February 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Manaresi:2004:IAT

[822] R. Manaresi and F. Selleri. The international atomic time and the velocity of light. Foundations of Physics Letters, 17(1):65–79, February 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Caban:2004:LCC

[823] P. Caban, M. Foryś, and J. Rembieliński. Lorentz covariant canonical formalism for free massive, massless and tachyonic particles. Foundations of Physics Letters, 17(1):81–88, February 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Auletta:2004:WCV

[824] Gennaro Auletta and Gino Tarozzi. Wavelike correlations versus path detection: Another form of complementarity. Foundations of Physics Letters, 17(1):89–95, February 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ben-Aryeh:2004:CGH

[825] Y. Ben-Aryeh and A. Mann. The correlations of Greenberger–Horne-Zeilinger states described by Hilbert–Schmidt decomposition. Foundations of Physics Letters, 17(1):97–101, February 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Giunti:2004:CWP

[826] Carlo Giunti. Coherence and wave packets in neutrino oscillations. Foundations of Physics Letters, 17(2):103–124, April 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Klauber:2004:AAB

[827] Robert D. Klauber. Analysis of the anomalous Brillet and Hall experimental result. Foundations of Physics Letters, 17(2):125–147, April 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:2004:DDE

[828] M. W. Evans. Derivation of Dirac’s equation from the Evans wave equation. Foundations of Physics Letters, 17(2):149–166, April 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Shan:2004:QCC

[829] Gao Shan. Quantum collapse, consciousness and superluminal communication. Foundations of Physics Letters, 17(2):167–182, April 2004. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Davidovic:2004:SLI

[830] Dragomir M. Davidović and Dusan Arsenović. Some logical implications of the two Einstein assumptions used in his analysis of geometry on a relativistically rotating disc. Foundations of Physics Letters, 17(2):183– 195, April 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Svetlichny:2004:NLQ

[831] George Svetlichny. Non-linear quantum mechanics and high energy cosmic rays. Foundations of Physics Letters, 17(2):197–202, April 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Adamenko:2004:MSS

[832] S. V. Adamenko and V. I. Vysotskii. Mechanism of synthesis of superheavy nuclei via the process of controlled electron-nuclear collapse. Foundations of Physics Letters, 17(3):203–233, June 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Narnhofer:2004:TEF

[833] Heide Narnhofer. The time evolution of fluctuation algebra in mean field theories. Foundations of Physics Letters, 17(3):235–253, June 2004. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Standish:2004:WOR

[834] Russell K. Standish. Why Occam’s razor. Foundations of Physics Letters, 17(3):255–266, June 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:2004:UGS

[835] M. W. Evans. Unification of gravitational and strong nuclear fields. Foundations of Physics Letters, 17(3):267–276, June 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Rebilas:2004:URE

[836] Krzysztof Rebilas. On the Unnikrishnan resolution of the EPR puzzle. Foundations of Physics Letters, 17(3):277–286, June 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Berger:2004:EBB

[837] Jorge Berger. Extension of the de Broglie–Bohm theory to the Ginzburg– Landau equation. Foundations of Physics Letters, 17(3):287–294, June 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Brandt:2004:IGR

[838] Howard E. Brandt. Intrinsic gravitational regularization of quantum fields. Foundations of Physics Letters, 17(3):295–300, June 2004. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:2004:POS

[839] Myron W. Evans. Physical optics, the Sagnac effect, and the Aharonov– Bohm effect in the Evans unified field theory. Foundations of Physics Letters, 17(4):301–322, August 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Rubin:2004:TNB

[840] Mark A. Rubin. There is no basis ambiguity in Everett quantum mechanics. Foundations of Physics Letters, 17(4):323–341, August 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Grossing:2004:CHF

[841] Gerhard Grössing. From classical Hamiltonian flow to quantum theory: Derivation of the Schrödinger equation. Foundations of Physics Letters, 17(4):343–362, August 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Nikolic:2004:BPT

[842] Hrvoje Nikolić. Bohmian particle trajectories in relativistic bosonic quantum field theory. Foundations of Physics Letters, 17(4):363–380, August 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gregor:2004:QHQ

[843] Malcolm H. Mac Gregor. Quantum Hall quantized cyclotron entrance channels. Foundations of Physics Letters, 17(4):381–391, August 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:2004:DGP

[844] Myron W. Evans. Derivation of the geometrical phase from the Evans phase law of generally covariant unified field theory. Foundations of Physics Letters, 17(4):393–401, August 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Zafiris:2004:QES

[845] Elias Zafiris. On quantum event structures. Part III: Object of truth values. Foundations of Physics Letters, 17(5):403–432, October 2004. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:2004:ELD

[846] Myron W. Evans. The Evans lemma of differential geometry. Foundations of Physics Letters, 17(5):433–455, October 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ghosal:2004:ASL

[847] S. K. Ghosal, Biplab Raychaudhuri, Anjan Kumar Chowdhury, and Minakshi Sarker. On the anistropy of the speed of light on a rotating platform. Foundations of Physics Letters, 17(5):457–477, October 2004. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Szabo:2004:MLC

[848] László E. Szabó. On the meaning of Lorentz covariance. Foundations of Physics Letters, 17(5):479–496, October 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anastasovski:2004:DEW

[849] P. K. Anastasovski, P. Carpenter, C. Ciubotariu, W. T. Coffey, G. J. Evans, M. W. Evans, R. Flower, D. Hamilton, J. B. Hart, T. Kurata, A. Labounsky, B. Lehnert, J. K. Moscicki, and S. Roy. Development of the Evans wave equation in the weak-field limit: the electrogravitic equation. Foundations of Physics Letters, 17(5):497–501, October 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sidharth:2004:PSU

[850] B. G. Sidharth. The Planck scale in the universe. Foundations of Physics Letters, 17(5):503–506, October 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kim:2004:EIB

[851] Ilki Kim and Gerald J. Iafrate. Entanglement in the interaction between two quantum oscillators systems. Foundations of Physics Letters, 17(6): 507–534, November 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:2004:DEW

[852] Myron W. Evans. Derivation of the Evans wave equation from the Lagrangian and action: Origin of the Planck constant in general relativity. Foundations of Physics Letters, 17(6):535–559, November 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Zawisky:2004:BNI

[853] M. Zawisky. Basic neutron interference experiments and the information interpretation of the wavefunction. Foundations of Physics Letters, 17(6): 561–579, November 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hamhalter:2004:TDS

[854] Jan Hamhalter. Traces, dispersions of states and hidden variables. Foundations of Physics Letters, 17(6):581–597, November 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Selleri:2004:COD

[855] F. Selleri. Curing one discontinuity with another: Comment on S. K. Ghosal et al., “On the Anisotropy of the Speed of Light on a Rotating Platform”. Foundations of Physics Letters, 17(6):599–606, November 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Robinett:2004:ARGa

[856] R. W. Robinett and L. C. Bassett. Analytic results for Gaussian wave packets in four model systems: I. visualization of the kinetic energy. Foundations of Physics Letters, 17(7):607–625, December 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Buonaura:2004:MEI

[857] B. Buonaura. Maxwell equations and inertial transformations. Foundations of Physics Letters, 17(7):627–644, December 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Robinett:2004:ARGb

[858] R. W. Robinett and L. C. Bassett. Analytic results for Gaussian wave packets in four model systems: II. autocorrelation functions. Foundations of Physics Letters, 17(7):645–661, December 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:2004:DLB

[859] Myron W. Evans. Derivation of the Lorentz boost from the Evans wave equation. Foundations of Physics Letters, 17(7):663–676, December 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Schulze-Halberg:2004:NEL

[860] Axel Schulze-Halberg. Non-existence of Liouvillian solutions for a free quantum particle on a torus surface,Part 1: Polar states. Foundations of Physics Letters, 17(7):677–690, December 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Khrennikov:2004:EBE

[861] Andrei Khrennikov. EPR–Bohm experiment and interference of probabilities. Foundations of Physics Letters, 17(7):691–700, December 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Iorio:2005:ATC

[862] Lorenzo Iorio. An analytical treatment of the clock paradox in the framework of the special and general theories of relativity. Foundations of Physics Letters, 18(1):1–19, February 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Herrera:2005:GQS

[863] L. Herrera. Geodesics in a Quash-spherical spacetime: a case of gravitational repulsion. Foundations of Physics Letters, 18(1):21–36, February 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:2005:ESEa

[864] M. W. Evans. The electromagnetic sector of the Evans field theory. Foundations of Physics Letters, 18(1):37–51, February 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Boya:2005:RBR

[865] Luis J. Boya and E. C. G. Sudarshan. Relation between rigid-body motion, isotropic cones, and spinors. Foundations of Physics Letters, 18(1):53–63, February 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gillies:2005:ETT

[866] G. T. Gillies and S. W. Allison. Experimental test of a time-temperature formulation of the uncertainty principle via nanoparticle fluorescence. Foundations of Physics Letters, 18(1):65–74, February 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Oyewumi:2005:ASK

[867] K. J. Oyewumi. Analytical solutions of the Kratzer–Fues potential in an arbitrary number of dimensions. Foundations of Physics Letters, 18 (1):75–84, February 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dzhunushaliev:2005:CFT

[868] Vladimir Dzhunushaliev. Colored flux tube in Euclidean spacetime. Foundations of Physics Letters, 18(1):85–93, February 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Stubbins:2005:PDP

[869] Calvin Stubbins. Particle dynamics of a phase space distribution function. Foundations of Physics Letters, 18(1):95–103, February 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:2005:C

[870] Anonymous. Corrigenda. Foundations of Physics Letters, 18(1):105, February 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524

(electronic). URL http://link.springer.com/content/pdf/10.1007/
s10702-005-2484-6.pdf.

Tessier:2005:CRM

[871] Tracey E. Tessier. Complementarity relations for multi–qubit systems. Foundations of Physics Letters, 18(2):107–121, April 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Nikolic:2005:BPT

[872] Hrvoje Nikolić. Bohmian particle trajectories in relativistic fermionic quantum field theory. Foundations of Physics Letters, 18(2):123–138, April 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:2005:NCE

[873] M. W. Evans. New concepts from the Evans unified field theory. Part One: The evolution of curvature, oscillatory universe without singularity, causal quantum mechanics, and general force and field equations. Foundations of Physics Letters, 18(2):139–155, April 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Acosta:2005:MGF

[874] G. D. Acosta and M. Kirchbach. Massive gauge fields and the Planck scale. Foundations of Physics Letters, 18(2):157–170, April 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hoffmann:2005:KFP

[875] Karl Heinz Hoffmann and J. Christian Schön. Kinetic features of preferential trapping on energy landscapes. Foundations of Physics Letters, 18(2):171–182, April 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Crowell:2005:SHH

[876] Lawrence B. Crowell and Robert Betts. Spacetime holography and the Hopf fibration. Foundations of Physics Letters, 18(2):183–193, April 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Leifer:2005:AGT

[877] P. Leifer. An affine gauge theory of elementary particles. Foundations of Physics Letters, 18(2):195–203, April 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Badescu:2005:CET

[878] Viorel Badescu and Peter T. Landsberg. Canonical ensemble with temperature limitation. Foundations of Physics Letters, 18(3):205–226, June 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ostrowski:2005:CET

[879] Marcin Ostrowski. Casimir effect for tachyonic fields. Foundations of Physics Letters, 18(3):227–242, June 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Jin:2005:CSS

[880] Wei Min Jin. Common space of spin and spacetime. Foundations of Physics Letters, 18(3):243–258, June 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:2005:ESEb

[881] M. W. Evans. The electromagnetic sector of the Evans field theory. Foundations of Physics Letters, 18(3):259–273, June 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hernandes:2005:RPC

[882] J. A. Hernandes, E. Capelas de Oliveira, and A. K. T. Assis. Resistive plates carrying a steady current: Electric potential and surface charges close to the battery. Foundations of Physics Letters, 18(3):275–289, June 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Schulze-Halberg:2005:QES

[883] Axel Schulze-Halberg. On quasi-exact solvability of the Schrödinger equation for a free particle on the surface of a spindle torus. Foundations of Physics Letters, 18(3):291–300, June 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ivezic:2005:DBS

[884] Tomislav Ivezić. The difference between the Standard and the Lorentz transformations of the electric and the magnetic fields. Foundations of Physics Letters, 18(4):301–324, August 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Selleri:2005:ITR

[885] F. Selleri. The inertial transformations and the relativity principle. Foundations of Physics Letters, 18(4):325–339, August 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Lurcat:2005:SPT

[886] Francçois Lurçat. Spin physics and the theory of strong interactions. Foundations of Physics Letters, 18(4):341–355, August 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Behera:2005:RQN

[887] Laxmidhar Behera, Indrani Kar, and Avshalom C. Elitzur. A recurrent quantum neural network model to describe eye tracking of moving targets.

Foundations of Physics Letters, 18(4):357–370, August 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Afriat:2005:CLM

[888] Alexander Afriat. Cartesian and Lagrangian momentum. Foundations of Physics Letters, 18(4):371–378, August 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Olum:2005:CCL

[889] Ken D. Olum and Allen Everett. Can a circulating light beam produce a time machine. Foundations of Physics Letters, 18(4):379–385, August 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sidharth:2005:NCG

[890] B. G. Sidharth. A note on the characterization of gravitation. Foundations of Physics Letters, 18(4):387–391, August 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sidhart:2005:NWI

[891] B. G. Sidhart. Neutrinos and the weak interactions. Foundations of Physics Letters, 18(4):393–399, August 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ivezic:2005:AGF

[892] Tomislav Ivezić. Axiomatic geometric formulation of electromagnetism with only one axiom: the field equation for the bivector field F with an explanation of the Trouton–Noble experiment. Foundations of Physics Letters, 18(5):401–429, October 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:2005:SCS

[893] M. W. Evans. The spinning and curving of spacetime: the electromagnetic and gravitational fields in the Evans field theory. Foundations of Physics Letters, 18(5):431–454, October 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Robinett:2005:SEP

[894] R. W. Robinett, M. A. Doncheski, and L. C. Bassett. Simple examples of position–momentum correlated Gaussian free-particle wave packets in one dimension with the general form of the time-dependent spread in position. Foundations of Physics Letters, 18(5):455–475, October 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Petrov:2005:SBH

[895] A. N. Petrov. The Schwarzschild black hole as a point particle. Foundations of Physics Letters, 18(5):477–489, October 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Hogarth:2005:CSM

[896] Mark Hogarth. Conventionality of simultaneity: Malament’s result revisited. Foundations of Physics Letters, 18(5):491–497, October 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dowker:2005:SWF

[897] Fay Dowker and Isabelle Herbauts. The status of the wave function in dynamical collapse models. Foundations of Physics Letters, 18(6):499–518, November 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Schulze-Halberg:2005:FPT

[898] Axel Schulze-Halberg. Form-preserving transformations for Hamiltonians with linear terms in the momentum. Foundations of Physics Letters, 18 (6):535–548, November 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Nikolic:2005:RQM

[899] Hrvoje Nikolić. Relativistic quantum mechanics and the Bohmian interpretation. Foundations of Physics Letters, 18(6):549–561, November 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Grinbaum:2005:ITP

[900] Alexei Grinbaum. Information-theoretic principle entails orthomodularity of a lattice. Foundations of Physics Letters, 18(6):563–572, November 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kapuscik:2005:PMF

[901] Edward Kapuścik and Pawel Szczeszek. The physical mechanism of formation of quantum mechanical Gamow states. Foundations of Physics Letters, 18(6):573–580, November 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

SantAnna:2005:LNI

[902] Adonai S. Sant’Anna. Labels for non-individuals? Foundations of Physics Letters, 18(6):??, November 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Rodriguez:2005:NPC

[903] Yeinzon Rodríguez and Carlos Quimbay. Are the new physics contributions from the left–right symmetric model important for the indirect CP violation in the neutral b mesons? Foundations of Physics Letters, 18 (7):581–602, December 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ghosal:2005:PET

[904] S. K. Ghosal, Saroj Nepal, and Debarchana Das. The principle of equivalence and the twin paradox. Foundations of Physics Letters, 18(7):603– 619, December 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ainsworth:2005:SEE

[905] Peter Ainsworth. The spin-echo experiment and statistical mechanics. Foundations of Physics Letters, 18(7):621–635, December 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Khrennikov:2005:GQM

[906] Andrei Khrennikov. Generalizations of quantum mechanics induced by classical statistical field theory. Foundations of Physics Letters, 18(7): 637–650, December 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Alhaidari:2005:TDD

[907] A. D. Alhaidari. The three-dimensional Dirac oscillator in the presence of Aharonov–Bohm and magnetic monopole potentials. Foundations of Physics Letters, 18(7):651–664, December 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

McCabe:2005:TBU

[908] Gordon McCabe. The topology of branching universes. Foundations of Physics Letters, 18(7):665–676, December 2005. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Shafiee:2006:IBZ

[909] A. Shafiee, F. Safinejad, and F. Naqsh. Information and the Brukner– Zeilinger interpretation of quantum mechanics: a critical investigation. Foundations of Physics Letters, 19(1):1–20, February 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Adamenko:2006:NPN

[910] S. V. Adamenko and V. I. Vysotskii. Neutronization and protonization of nuclei: Two possible ways of the evolution of astrophysical objects and the laboratory electron–nucleus collapse. Foundations of Physics Letters, 19(1):21–36, February 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Chubykalo:2006:HTG

[911] A. Chubykalo, A. Espinoza, R. Alvarado Flores, and A. Gutierrez Rodriguez. Helmholtz theorem and the V -gauge in the problem of superluminal and instantaneous signals in classical electrodynamics. Foundations of Physics Letters, 19(1):37–49, February 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See comment [955].

Tiwari:2006:NEC

[912] S. C. Tiwari. The nature of electronic charge. Foundations of Physics Letters, 19(1):51–62, February 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cramer:2006:TAI

[913] John G. Cramer. A transactional analysis of interaction-free measurements. Foundations of Physics Letters, 19(1):63–73, February 2006. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Jones:2006:CPS

[914] Preston Jones and Lucas F. Wanex. The clock paradox in a static homogeneous gravitational field^{1}. Foundations of Physics Letters, 19(1):75–85, February 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sidharth:2006:BHT

[915] B. G. Sidharth. Black-hole thermodynamics and electromagnetism. Foundations of Physics Letters, 19(1):87–94, February 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kirkpatrick:2006:SHT

[916] K. A. Kirkpatrick. The Schrödinger–HJW theorem. Foundations of Physics Letters, 19(1):95–102, February 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Rowland:2006:NOS

[917] David R. Rowland. Noninertial observers in special relativity and clock synchronization debates. Foundations of Physics Letters, 19(2):103–126, April 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Skulimowski:2006:QWQ

[918] Marcin Skulimowski. Quantum world with quantum time. Foundations of Physics Letters, 19(2):127–141, April 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Lan:2006:SMB

[919] Boon Leong Lan and Ying Oon Tan. A stochastic mechanics based on Bohm’s theory and its connection with quantum mechanics. Foundations of Physics Letters, 19(2):143–155, April 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dzhunushaliev:2006:NAQ

[920] Vladimir Dzhunushaliev. A non-associative quantum mechanics. Foundations of Physics Letters, 19(2):157–167, April 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dzhunushaliev:2006:SSF

[921] Vladimir Dzhunushaliev and Ratbay Myrzakulov. Soft singularity and the fundamental length. Foundations of Physics Letters, 19(2):169–177, April 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ali:2006:QSB

[922] Md. Manirul Ali, A. S. Majumdar, and Dipankar Home. Quantum superarrivals: Bohr’s wave–particle duality revisited. Foundations of Physics Letters, 19(2):179–186, April 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Su:2006:CBR

[923] Ching-Chuan Su. Connections between the Riemann and the Lorentz force laws. Foundations of Physics Letters, 19(2):187–194, April 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Dvoeglazov:2006:NLT

[924] Valeri V. Dvoeglazov and J. L. Quintanar González. A note on the Lorentz transformations for the photon. Foundations of Physics Letters, 19(2):

195–200, April 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Longhi:2006:LHS

[925] G. Longhi. Locality hypothesis and the speed of light. Foundations of Physics Letters, 19(3):201–223, June 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Arminjon:2006:DEH

[926] Mayeul Arminjon. Dirac equation from the Hamiltonian and the case with a gravitational field. Foundations of Physics Letters, 19(3):225–247, June 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Villavicencio:2006:RSR

[927] M. Villavicencio, J. A. E. Roa-Neri, and J. L. Jiménez. The role of a specific response of media in the Vavilov–Cherenkov effect. Foundations of Physics Letters, 19(3):249–258, June 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Nikolic:2006:CPC

[928] Hrvoje Nikolić. Causal paradoxes: a conflict between relativity and the arrow of time. Foundations of Physics Letters, 19(3):259–267, June 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Salti:2006:ESS

[929] Mustafa Salti and Oktay Aydogdu. Energy in the Schwarzschild–de Sitter spacetime. Foundations of Physics Letters, 19(3):269–276, June 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Carmeli:2006:CTT

[930] Moshe Carmeli, John G. Hartnett, and Firmin J. Oliveira. The cosmic time in terms of the redshift. Foundations of Physics Letters, 19(3):277– 283, June 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See comment [961].

Wesson:2006:VI

[931] Paul S. Wesson. Vacuum instability. Foundations of Physics Letters, 19(3):285–291, June 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kak:2006:TSQ

[932] Subhash Kak. A three–stage quantum cryptography protocol. Foundations of Physics Letters, 19(3):293–296, June 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Anonymous:2006:C

[933] Anonymous. Corrigenda. Foundations of Physics Letters, 19(3):297, June 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Khrennikov:2006:PCS

[934] Andrei Khrennikov. Prequantum classical statistical field theory: Complex representation, Hamilton-Schrödinger equation, and interpretation of stationary states. Foundations of Physics Letters, 19(4):299–319, August 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Miller:2006:CRT

[935] D. J. Miller. Counterfactual reasoning in time-symmetric quantum mechanics. Foundations of Physics Letters, 19(4):321–335, August 2006. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Rebilas:2006:WDM

[936] Krzysztof Rebilas. A way to discover Maxwell’s equations theoretically. Foundations of Physics Letters, 19(4):337–351, August 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Minguzzi:2006:RPD

[937] E. Minguzzi. Relativity principles in 1 + 1 dimensions and differential aging reversal. Foundations of Physics Letters, 19(4):353–365, August 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Salesi:2006:DSB

[938] Giovanni Salesi. Deriving spin of the bosonic string^{1}. Foundations of Physics Letters, 19(4):367–376, August 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Carazza:2006:DMB

[939] B. Carazza and R. Tedeschi. On the decoherence of macroscopic bodies. Foundations of Physics Letters, 19(4):379–388, August 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kosinski:2006:NCS

[940] Piotr Kosinski. Note on clock synchronization and Edwards transformations. Foundations of Physics Letters, 19(4):389–397, August 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sidharth:2006:NMP

[941] B. G. Sidharth. A note on massive photons. Foundations of Physics Letters, 19(4):399–402, August 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sen:2006:CCL

[942] D. Sen and S. Sengupta. A critique of the classical limit problem of quantum mechanics. Foundations of Physics Letters, 19(5):403–421, October 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Fahr:2006:CCS

[943] Hans J. Fahr. Cosmological consequences of scale-related comoving masses for cosmic pressure, mass, and vacuum energy density. Foundations of Physics Letters, 19(5):423–440, October 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cardone:2006:MIR

[944] F. Cardone, S. P. Maidanyuk, R. Mignani, and V. S. Olkhovsky. Multiple internal reflections during particle and photon tunneling. Foundations of Physics Letters, 19(5):441–457, October 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Drezet:2006:QNS

[945] A. Drezet. Quantum nonlocality of single photon state? Foundations of Physics Letters, 19(5):459–470, October 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Ord:2006:DDC

[946] G. N. Ord, J. A. Gualtieri, and R. B. Mann. A discrete, deterministic construction of the phase in Feynman paths. Foundations of Physics Letters, 19(5):471–480, October 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

SantAnna:2006:IFM

[947] Adonai S. Sant’Anna and Otávio Bueno. Interaction-free measurement with matter waves. Foundations of Physics Letters, 19(5):481–490, October 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Boyer:2006:PET

[948] Timothy H. Boyer. Proposed experimental test for the paradoxical forces associated with the Aharonov–Bohm phase shift. Foundations of Physics Letters, 19(5):491–498, October 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sidharth:2006:BNN

[949] B. G. Sidharth. A brief note on a new cosmic “coincidence” and new physics. Foundations of Physics Letters, 19(5):499–500, October 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Jarvis:2006:BRG

[950] P. D. Jarvis and S. O. Morgan. Born reciprocity and the granularity of spacetime. Foundations of Physics Letters, 19(6):501–517, November 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Oliveira:2006:CCF

[951] Firmin J. Oliveira and John G. Hartnett. Carmeli’s cosmology fits data for an accelerating and decelerating universe without dark matter or dark energy. Foundations of Physics Letters, 19(6):519–535, November 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kohler:2006:FMT

[952] Matthew Kohler. Finding the missing time in the instantaneous turnaround version of the twin paradox. Foundations of Physics Letters, 19(6):537–551, November 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Nikolic:2006:CMD

[953] Hrvoje Nikolić. Classical mechanics without determinism. Foundations of Physics Letters, 19(6):553–566, November 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Tanaka:2006:CRY

[954] Sho Tanaka. Contracted representation of Yang’s spacetime algebra and Buniy–Hsu–Zee’s discrete spacetime. Foundations of Physics Letters, 19 (6):567–578, November 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Heras:2006:CHT

[955] José A. Heras. Comment on “Helmholtz Theorem and the V -Gauge in the Problem of Superluminal and Instantaneous Signals in Classical Electrodynamics,” by A. Chubykalo et al. Foundations of Physics Letters, 19 (6):579–590, November 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See [911].

Gorka:2006:LQM

[956] Przemysław Górka. Logarithmic quantum mechanics: Existence of the ground state. Foundations of Physics Letters, 19(6):591–601, November 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Xulu:2006:BTE

[957] S. S. Xulu. Bergmann-Thomson energy of a charged rotating black hole. Foundations of Physics Letters, 19(6):603–609, November 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Sidharth:2006:TV

[958] B. G. Sidharth. Tests for varying G. Foundations of Physics Letters, 19 (6):611–617, November 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Caban:2006:ODD

[959] P. Caban, J. Rembieliński, K. A. Smoliński, and Z. Walczak. Oscillations do not distinguish between massive and tachyonic neutrinos. Foundations of Physics Letters, 19(6):619–623, November 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Kracklauer:2006:WWR

[960] A. F. Kracklauer. What’s wrong with this rebuttal? Foundations of Physics Letters, 19(6):625–629, November 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Macdonald:2006:CCT

[961] Alan Macdonald. Comment on “The Cosmic Time in Terms of the Redshift,” by Carmeli et al. Foundations of Physics Letters, 19(6):631–632, November 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic). See [930].